A Bibliography of Publications in Theoretical Computer Science: 1995–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.23

Title word cross-reference

(3/2) log 3 [723]. (α) [1110]. (A ∨ B) ⇒ C [1133]. (q) [1379]. ∗ [36]. + [666, 1191]. 0 [888, 104, 760, 161, 335, 430, 1164]. 0, 1, , n − 1 [548]. 1 [880, 133, 162, 760, 1164]. 101 [1254]. 180 [1037]. 2 [1066, 1283, 171, 266, 1179, 583, 1000, 1369]. 2.5 [167]. 2λ [528]. 3 [51, 729, 1348, 822, 782]. 4 [549]. = [17]. ^{0} [1225]. A^{∗} [1180]. α [1373, 597]. B_{n} [548]. β [1136]. β = λ [282]. cβ [829]. D [1330, 1358]. Δ [604]. E [1344, 98, 1368]. E_{2} [98]. η [1136]. F [770, 106]. F_{<} [37]. F_{q^{n}} [1383]. G [278]. h [879]. H(X) [1285]. K [1123, 58, 1143, 1080, 376, 438, 10, 1180, 879, 1039, 1082]. κ [666, 1191]. L_{1} [815]. λ [628, 228, 1193, 1393, 1394, 700, 517, 1347, 89, 4, 1198, 464, 276, 365, 366]. λ^{{}} [628]. λβ [829]. log_{2} n + O(1) [862]. m [558, 163]. μ [216, 177, 1048, 1237, 407, 1365]. N [1208, 616, 971]. n + c^{ste} [1208]. n + c^{th} [1208]. ≠ [159]. O(n) [971]. O(n^{5}/2) [720]. Ω [756, 217, 745, 849, 940, 847, 592, 90]. P [294, 1424, 342, 329, 163]. P_{4} [693, 136]. π [1004, 485, 480, 969]. P ω [883, 283]. r [204]. R^{q} [967]. S [969]. σ [270, 969]. σ^{+} [270]. ε [10]. W [1] [62]. W [2] [1147]. X [364]. x_{0}u_{1}^{i}x_{1}u_{2}^{i}x_{2}u_{3}^{i}x_{3} = y_{0}v_{1}^{i}y_{1}v_{2}^{i}y_{2}v_{3}^{i}y_{3}(i = 0, 1, 2,) [1374]. Z/nZ [1380]. Z^{3} [421, 782]. Z_{m} [1330, 1085]. Z_{N} [1382].

- [163]. -abstraction [628]. -adjacent [376]. -approximation [1179]. -approximations [10]. -Arbiter [879]. -arboretum [1143].

-ary [58]. -automata [969]. -autonomous [36]. -bisimilar [1347]. -body [971]. -bounded [880]. -calculi [517]. -Calculus [480, 228, 216, 177, 1193, 1393, 1004, 485, 4, 528, 1198, 464, 407, 365, 366, 1237, 1365]. -categories [756]. -center [1123]. -chain [1082]. -competitive [723]. -complete [163]. -connected [549]. -consistent [204]. -denotational [666, 1191]. -dependent [106]. -dimensional [1330]. -écritures [1208]. -free [616]. -game [270]. -generator [1066]. -Horn [770]. -infix [822]. -label [10]. -languages [217, 745, 849, 847, 90, 604]. -logics [1048]. -long [1136]. -Machine [829]. -machines [1394]. -matrices [969]. -median [1123]. -networks [278]. -normal [1136]. -notation [276]. -opt [1180]. -out [879]. -paperfolding [329]. -partition [1039]. -polynomials [969]. -reducible [693, 136]. -reduction [829]. -regular [940, 592]. -safe [162]. -SAT [1348]. -selective [342, 1424, 294]. -solvable [133]. -space [597]. -Spaces [1358]. -sparse [693]. -stage [729]. -structures [266]. -tape [1369]. -terms [282, 700, 89]. -trees [438]. -unification [1344, 1368]. -words [1373]. -writings [1208].

0 [1398].

1-optimal [1061]. 150 [213, 211, 210]. 160 [374]. 170 [530]. 171-180 [701].

2 [366]. 218 [1428].

3rd [352].

58 [344]. 59 [440]. 5NF [779]. 5th [791].

60 [542]. 61 [736]. 62 [737].

’94 [345]. ’95 [886]. ’96 [964, 1253].

A. [1149, 72]. abducibles [1106].

Abduction [827]. abductive [446, 1106]. abelian [1141, 1263]. Abstract [986, 573, 81, 636, 977, 898, 199, 268, 1047, 577, 1335, 1342, 645, 1234, 1408, 1345, 397, 1016, 362, 1233]. abstraction [1392, 628, 473, 223, 874, 45, 488]. AC [1225, 85]. AC-compatible [85]. acceptance [185, 578, 892]. access [958, 950]. according [584]. accumulations [507]. Accuracy [396]. achieve [1307]. Achilles [1166]. action [1407, 1322]. actions [873, 884, 47]. active [977]. Actor [1320]. acyclic [1058]. adaptation [1351]. adaptive [184, 1309]. Addition [862, 449, 396, 955]. Additive [584, 270, 315, 1362, 724]. additives [999]. adequacy [632]. adjacent [376]. Adjoining [1093, 1177]. adjunction [638]. admissible [1377]. Advances [1252, 764]. advice [1126, 271]. Affine [963, 129, 295]. affines [295]. agent [826, 874, 480]. agent-passing [480]. aggregating [7]. aggregation [882]. aggregative [462]. ahead [297, 105]. Ajtai [1381]. aléatoire [349, 346]. Algebra [791, 644, 828, 199, 311, 942, 405, 203, 647, 1181, 801, 280, 317, 524, 641, 646]. Algebraic [310, 758, 1238, 1011, 209, 1012, 1384, 1019, 1015, 63, 307, 520, 113, 838, 870, 442, 31, 514, 315, 358, 21, 1258, 749, 1292, 746, 747, 794, 1024, 363, 39, 470, 535, 462, 1200, 812, 1184, 1016, 796, 259, 362, 281, 1360]. algebraically [935]. algebras [406, 758, 1238, 643, 14, 1005, 820, 677, 801, 45, 1299, 1184, 1300, 800]. Algèbres [801]. ALGOL [488, 977, 1392]. ALGOL-like [488]. algorithm [1379, 347, 167, 138, 586, 293, 336, 659, 356, 1042, 1160, 658, 511, 434, 353, 1314, 414, 1179, 334, 439, 136, 118, 853, 823, 1311, 508, 1385, 109, 1140, 866, 723, 971, 1210, 800, 1418]. algorithme [353]. Algorithmes [377]. Algorithmic [310, 717, 18, 466, 1351, 256, 1267, 940, 1148]. algorithmically [856]. Algorithms [1420, 317, 1256, 457, 532, 733, 1061, 1145, 169, 606, 1062, 694, 408, 1146, 718, 1415, 706, 887, 859, 1417, 401, 1180, 893, 974, 413, 1277, 1416, 400, 966, 1296, 1041, 182, 1022, 57, 821, 1014, 55, 123, 1058, 1211, 1039, 168, 377]. Alignment [97, 733, 138, 158]. alignments [1180]. Alimohamed [142]. all-optical [1324]. allocating [1161]. Allocation [539, 403, 1052]. allowing [1341, 1280]. Almost [974, 1327, 425]. along [572]. Alphabet [1091, 1175, 1149, 135, 437]. Alphabet-independent [1091]. alphabetic [695]. alternating [172, 96, 91, 61, 685]. Alternation [691, 938, 685]. alternative [97]. always [369]. AM [17]. ambiguity [711]. amidst [51, 54]. among [1398, 1157]. analog [1034]. Analysis [1421, 118, 1420, 1041, 968, 445, 18, 19, 862, 169, 1332, 1288, 1289, 894, 467, 1259, 808, 1229, 116, 1338, 1198, 1348, 1294, 350, 119, 427, 1297, 477, 1022]. Analytic [1290, 1117, 1173]. Analytical [1031]. ancestor [1114]. animals [350]. annealing [1209]. Annotated [536]. Anomalous [424]. Anonymous [126]. answer [1006]. Answering [537]. anti [1363]. anti-founded [1363]. Antimirov [285, 286]. any [1377]. aperiodic [1158, 420]. application [319, 435, 1208, 645]. Applications [809, 295, 789, 347, 692, 607, 375, 1247, 947, 74, 12, 455, 811, 1169, 1031, 256, 549, 174, 477, 1250, 1305, 1023, 809]. Applying [1248]. approach [631, 241, 900, 758, 1238, 894, 1055, 904, 577, 665, 40, 1167, 157, 21, 1003, 912, 535, 940, 1105, 684, 926, 1322, 1334]. approximability [166, 1155, 1372, 709]. Approximate [205, 196, 1059, 813, 1292, 1180, 712, 1041]. approximately [591]. Approximating [609, 1092, 1157, 1175]. Approximation [733, 1176, 692, 442, 333, 1146, 404, 1179, 324, 661, 840, 1202, 1426, 110].

Approximations [1122, 10]. Arbiter [879]. arbitrary [404]. arboretum [1143]. area [346]. Argument [151]. arguments [539, 298]. arithmetic [1144, 1276, 301, 569, 395, 1164, 159, 261, 590]. arithmetical [1166, 1341, 565]. arithmetics [743]. arithmétique [301]. Array [959, 91, 73]. arrays [972, 973]. articles [1205, 212]. artificially [1245]. ary [58]. Aspects [595, 232, 1255, 1361, 1137, 1138]. assertional [825]. assertions [562]. Assigning [731]. assignment [520, 1124, 229]. associated [331, 426]. Associative [801]. associatives [801]. associés [331]. Assumption [474, 42]. Assumption/guarantee [474]. asteroidal [612]. Asymptotic [839, 117, 1277, 437]. asymptotically [963]. asymptotics [112, 115, 1229]. asynchronous [944, 1314, 946, 510, 582]. atomic [201, 948, 198]. atomicity [197]. Attack [1380]. Attribute [1251]. augmented [51]. authenticated [1312]. Author [781]. Autocorrelation [861]. automata [165, 819, 1186, 1246, 270, 299, 420, 580, 122, 1220, 383, 651, 1037, 1243, 894, 31, 1283, 171, 584, 848, 1216, 512, 1248, 811, 1337, 711, 459, 96, 1219, 300, 1119, 1330, 691, 1244, 180, 61, 510, 1369, 1371, 555, 1077, 478, 969, 1085, 1021, 463, 598, 1245, 306, 1282, 1242, 617, 1023, 1017]. automated [410]. automates [848]. Automatic [562, 807, 545, 412, 479]. Automaticity [819, 688]. automatiques [545]. automaton [287, 1094, 1388, 1210]. Automorphism [450]. autonomous [36]. Avant [1274]. Average [169, 862, 1272, 1027, 437]. averaging [716]. avoid [915]. Avoiding [566, 80]. axiom [1018]. axiomatics [1297]. Axiomatisation [1107]. Axiomatising [834]. axiomatization [238, 1189, 524].

B [72, 996]. B-graphs [996]. background [1251]. balancing [1314]. balistique [809]. ballistics [809]. Banach [1291, 984, 1401]. bananas [509]. banded [1042]. based [1149, 1231, 894, 1380, 753, 569, 230, 5, 873, 874, 1120, 939, 923, 1387, 85, 479, 193, 1389, 50, 1124, 198, 1303]. bases [1377, 1383, 537, 722, 1039]. Basic [405, 1095, 866]. basis [445, 353, 362]. BDDs [70]. be [1225, 458]. behavior [1243, 117, 638, 824]. behaviors [830, 794, 617]. behaviour [420, 1153]. Behavioural [441, 836, 1232, 473, 1239]. behaviours [238, 220, 664]. beholder [674, 673]. Beimel [1029]. belief [527]. Belnap [534]. Bernoulli [1229]. Berry [293]. Bertrand [704]. best [138]. beta [13, 920]. beta-reduction [920]. better [298]. between [880, 333, 46, 323, 1371, 426, 390]. beyond [219]. bidimensional [1156]. bidimensionnelle [1156]. bifix [1266, 137]. bifurcation [808, 713]. billard [432]. billiard [432]. Binary [710, 296, 156, 124, 309, 371, 292, 548, 681, 1318, 1270, 135, 810, 713]. binding [871, 1184]. bioremediation [1248]. bipartite [355, 613, 168]. bipartitioning [861]. bird [1415]. Bireflectivity [1406]. bisimilar [1347]. bisimilarities [1004]. Bisimilarity [1405, 334, 188]. Bisimulation [1365, 1334, 238, 676, 250, 937, 529]. Bisimulation-invariant [1365]. Bisimulations [1026, 944, 144, 30]. bit [1383]. bit-serial [1383]. bits [423, 557]. black [680]. black-box [680]. blessings [981]. Block [712, 1375]. Blum [896, 967]. Blum-Shub-Smale [967]. body [971]. Böhm [1200]. Böhm-out [1200]. Boltzmann [1247]. Book [1113]. Boolean [735, 692, 2, 124, 608, 788, 1354, 1096, 373, 613, 8, 1253, 70, 419, 844, 1130, 1391]. Boolean-valued [8]. border [251]. Borel [1301]. borné [848]. borrow [572]. Bottom [147, 297, 4]. Bottom-up [147, 297, 4]. bound [1094, 484, 108, 318, 557, 1164]. boundaries [546, 559]. Bounded [727, 483, 878, 626, 880, 550, 411, 1143, 1065, 1008, 99, 597, 380, 848, 711, 963, 193, 685]. bounded-complete [597]. Bounding [1117]. bounds [319, 1144, 692, 1148, 1192, 133, 1032, 309, 543, 321, 291, 298, 449, 812]. box [680, 247, 1181]. BPP [655, 671]. BPP-nets [671]. branch [360]. branches [961]. branching [557, 1049]. bridging [1416]. brief [27]. broadcast [1029, 160]. broadcasting [1052, 1026]. Broadening [995]. Bruijn [1343]. BSP [1061]. BSS [390]. Bubbles [865]. Buchberger [356]. Buchi [510]. bucket [119]. budget [1059]. build [1327]. building [1030]. bulk [950]. bulk-synchronous [950]. Bulletin [344, 440, 542, 736, 737]. Bus [806].

Caenorhabditis [1249]. cakes [1253]. calcul [809, 377, 801]. calculational [665]. calculations [1355]. calculi [361, 941, 517, 514, 1196, 1407, 67, 480, 275, 1201, 1202, 1426]. Calculus [480, 228, 944, 216, 518, 177, 88, 236, 1193, 978, 1393, 1195, 1004, 938, 343, 1333, 1009, 809, 751, 871, 246, 1188, 247, 485, 1005, 4, 1026, 528, 1198, 464, 601, 1237, 1410, 567, 991, 874, 1030, 407, 1365, 939, 365, 366]. Call [1410, 1423, 1108, 1338]. Call-by-name [1410, 1423]. call-by-need [1410]. call-by-value [1108, 1338, 1410]. can [1225, 458, 1173]. canonical [883]. capabilities [20]. capacity [727]. Caractérisation [1208]. cardinality [193]. cardinality-bounded [193]. cares [654]. Cartesian [468]. Case [5, 1323, 169, 1037, 67, 1348, 427, 688]. Case-based [5]. cases [1180]. cash [1389]. catch [600]. categorial [1011, 434]. categorical [143, 35, 926, 367]. categories [36, 634, 935, 1396, 756]. category [152, 268, 629, 468]. catenation [697]. causality [251, 1322]. Cayley [1371, 1078]. CCC’s [274]. CCS [43, 642]. cell [100]. cellular [165, 819, 1246, 270, 420, 122, 651, 1037, 1243, 1283, 171, 584, 1248, 811, 459, 91, 1219, 1330, 691, 1244, 180, 497, 510, 1371, 1085, 1245, 73, 306, 1282, 1242]. center [1123]. centralized [1230]. certain [694]. certificates [1064]. CGOOD [367]. Chain [383, 1103, 610, 1082]. change [1316]. channel [1332, 1124]. channels [925]. chaos [584]. chaotic [1037, 1243, 584]. character [408]. characteristic [1098]. characteristics [63]. Characterization [266, 849, 880, 143, 307, 392, 1147, 556, 421, 992, 824, 1297, 272, 1085, 1281, 259, 1208]. Characterizations [1088, 90, 1382, 415, 607]. characterized [1409]. checker [1197]. Checking [921, 1333, 752, 37, 940, 924]. Chervonenkis [9]. chip [854, 547]. chip-firing [547]. Choffrut [1224]. choice [1007]. Chor [1029]. chord [1068]. circles [1283]. circuit [850]. circuits [1144, 17, 859, 315, 815, 298, 583, 395, 8, 1164]. Circumscribing [885]. circumscription [730]. class [238, 1186, 127, 1381, 934, 120, 925, 419]. classes [546, 1126, 278, 438, 1057, 294, 1424, 454, 338, 323, 386, 814, 55, 415, 1154]. classical [978, 1393, 994, 600]. classification [1330, 525]. Classifying [1178]. clauses [284]. CLEAN [923]. climbing [1166]. clock [1186, 1008]. clocks [1319]. Clos [729]. Closed [343, 1407, 1173, 1287, 102, 468]. Closed-form [1173]. Closing [1053]. closure [740, 849, 462, 940, 513, 224, 1390]. CLP [465, 569]. cluster [1124]. CML [1232, 280]. CNF [1207]. Co [261, 159]. Co-NP [261, 159]. coalgebraic [1334]. coarsest [1008]. code [409, 1226, 559, 962, 807]. codes [1065, 851, 1266, 435, 292, 734, 416, 822, 722, 137]. codicity [165]. codings [186]. coefficient [861]. coefficients [1269, 377]. Cohen [340]. Coherence [1396]. Coherent [984, 1401].

Coin [436, 723]. Coin-sliding [436]. coincide [155]. collage [129]. collection [192]. coloured [1376, 618, 1023]. colouring [1063, 708]. Combination [86, 868, 1156]. combinational [108]. combinations [11]. Combinatoire [1156]. combinatoires [432]. combinatorial [1275, 1146, 114, 1255, 401, 1024, 273, 53, 432]. combinatorics [743, 1261]. combinators [504]. Combinatory [1009, 40]. Combining [514, 534]. combustion [808]. committed [939]. common [282, 858, 1129, 437, 110]. communicating [828, 1227, 588, 589, 237, 641]. Communication [1151, 1152, 319, 201, 1306, 1312, 484, 1002, 775, 953, 1214, 653]. commutation [645]. commutations [1364]. commutative [1144, 987, 555]. Commutativity [1067]. commuting [553]. Compact [725, 152, 880, 1285, 1287, 730, 883, 672]. companion [682]. Comparing [464]. comparison [408, 484, 732, 684]. Comparisons [1050, 336]. compatibility [894]. compatible [85]. competitive [968, 1162, 723]. Compilation [146]. compiler [59, 1012, 1250]. complementation [177, 510]. Complete [128, 238, 1189, 550, 150, 785, 156, 522, 171, 597, 935, 1007, 1110, 1318, 153, 1330, 1366, 1035, 163]. Completeness [368, 639, 216, 979, 127, 1370, 1227, 62, 209, 983, 1400, 98, 431, 476, 28, 993, 231, 460]. Completion [214, 137, 149, 876, 734, 83, 519, 272, 446, 362]. completions [1396]. complets [1035]. complex [192, 1248, 1170, 626]. complexité [1208]. Complexity [310, 1148, 162, 1271, 512, 816, 321, 9, 319, 1151, 166, 1316, 1092, 1126, 1192, 133, 1150, 158, 320, 885, 1118, 268, 831, 170, 1208, 484, 850, 1218, 827, 1104, 266, 380, 1152, 94, 1133, 315, 108, 438, 1053, 675, 742, 533, 316, 325, 989, 1402, 623, 449, 379, 100, 323, 190, 688, 339, 386, 814, 1027, 8, 812, 415, 1303, 1082, 341, 74]. component [825]. composed [409]. Composition [617]. compositional [445, 199, 904, 1237, 48]. Compositionality [1187]. Compositions [889, 297]. compressed [649]. COMPUGRAPH [886, 909]. COMPUGRAPH/ SEMAGRAPH [886]. Comput [1425, 735, 1428, 1424, 1112, 761, 1426]. Computability [1287, 896, 315, 1295, 1301, 1305, 1167, 1292, 1363, 390, 1304]. Computable [1165, 1291, 41, 1356, 392, 1288, 1289, 95, 132, 604, 1330, 394, 1297, 927]. Computation [886, 1245, 1300, 411, 20, 321, 1338, 1408, 834, 255, 391, 397, 109, 1299]. Computational [994, 859, 325, 728, 500, 1227, 312, 877, 742, 583, 339, 225]. computations [88, 580, 1247, 694, 99, 429, 1002, 987, 91, 948, 921, 892, 855, 397, 1298, 218, 489]. compute [901, 960]. Computer [791, 595, 242, 801, 1185]. computers [1034, 855]. Computing [195, 417, 507, 66, 279, 835, 582, 496, 1016, 1160, 905, 439, 298, 377, 1164, 9, 875, 971, 967]. concatenable [524]. concept [546, 930]. concerning [324, 1115, 593]. conclusion [447]. Concrete [1299]. concurrency [207, 260, 32, 1105, 525, 845]. Concurrent [905, 918, 830, 1044, 1425, 580, 753, 754, 67, 664, 1397, 567, 48, 41, 943]. concurrent-object [943]. condition [1050]. conditional [234, 444, 912, 917]. conditions [1308, 1050]. configurations [1332]. Conflict [1124, 252]. Conflict-Free [1124]. Confluence [754, 516, 269, 228, 444, 1025, 1102, 864, 866, 239, 603]. confluent [567]. conformisms [922]. confusion [251]. Congruence [1254, 1004, 1008, 1188, 1255, 43]. congruences [836, 1135, 745, 90]. conjecture [1036, 327, 1069]. conjugation [936]. conjunctive [517]. conjunctive-disjunctive [517]. connected [1357, 549]. connectivity [1064, 511]. conquer [400]. consensus [1306, 158, 775]. conservative [643]. considerations [741]. consistency [570, 209, 912, 563]. Consistent [448, 95, 135, 204]. Constant [1272, 19, 477, 1121]. Constant-space [1272]. constants [887]. Constrained [564, 83]. Constraint [26, 78, 1331, 868, 1101, 754, 569, 533, 563, 567, 340, 1106, 218]. constraints [1372, 565, 475, 373, 965, 767, 882, 759]. construct [323]. constructed [296]. Constructible [113]. Constructing [389, 634, 793, 265, 695]. construction [93, 303, 1012, 102, 103, 1077, 1250, 1202, 1426]. constructions [287, 375, 114]. Constructive [1288, 360, 912, 1294, 304]. constructivism [1185]. constructor [444]. constructor-sharing [444]. Contact [251]. containing [976]. contaminated [1248]. content [337]. Contents [344, 440, 542, 736, 737]. Context [917, 109, 198, 1227, 698, 888, 334, 384, 454, 974, 537, 450, 1050, 741]. Context-free [109, 1227, 698, 888, 334, 450, 1050, 741]. Context-sensitive [917, 537]. Context-specific [198]. contextual [1081, 742]. Continued [887, 1385, 1069]. Continuity [635, 1302]. continuous [20, 633, 984, 1401, 677, 1059, 25, 391]. continuous-state [25]. continuous-time [391]. contraction [1052]. contractum [241]. Contributions [200, 1113, 738]. control [38, 826, 958, 1337, 957, 174, 1411, 845, 1242]. Controllability [316]. controlled [664]. convergence [433, 1314, 1417, 1416, 756]. conversion [14]. conversions [1201]. convert [1210]. convex [288, 586, 51, 963, 812, 168, 1279]. convex-bipartite [168]. Cook [431]. Cooperation [698]. coordinated [858]. coordinating [874]. coordination [870, 873, 1307]. copies [763]. copying [1350]. copyright [1387]. corner [857]. corrected [779]. correction [1042]. Correctness [695, 358, 1012, 49, 508, 479, 302, 1014, 1360]. correlation [1382]. correlation-immune [1382]. Correspondence [378]. Corrigendum [1425, 1428, 1112, 761, 1426]. cost [309, 433, 895, 957, 1296]. costed [300]. costs [731]. counter [495]. counterfeit [723]. Counting [1068, 686, 1043, 194, 1308, 415, 981]. Coupled [1249, 384]. coupled-context-free [384]. Covariant [774]. Coven [420]. Cover [613, 1372, 1375, 1082]. Cover-preserving [613]. coverage [257]. covering [1273]. Covers [803, 89, 439]. CPS [1423]. CRCW [291, 1090]. criteria [83]. criterion [1360]. Crochemore [336]. Crochemore-Perrin [336]. crossbars [954]. crowding [1421]. Cryptographic [1387]. cryptosystem [1388]. cryptosystems [1380]. CSP [27, 29, 1182, 642]. cubes [1376]. cubic [1380]. Cumulative [1206, 213]. Curry [773]. curve [658, 132]. curves [1379, 703, 1380, 102]. customers [278]. cut [518, 718, 993, 1403]. cut-elimination [993, 1403]. cut-free [518]. cycle [102]. cycles [1078]. cyclic [1042, 916, 990]. cylindrical [425].

D [167, 51, 1283, 171, 782]. D0L [425]. dans [1070, 683, 432, 346]. DAQV [959]. data [703, 825, 577, 1335, 1259, 313, 1310, 897, 573, 222, 452, 120, 70, 960, 615, 367, 193, 198, 509, 1233]. data-movement-intensive [897]. data-parallel [825]. data-types [1233]. database [982, 779]. databases [372, 541, 839, 1104, 116, 533, 840, 843, 198, 538]. dataflow [477]. DATALOG [885, 841]. datatypes [35]. dead [962]. deadlines [1217, 121]. deadlock [1340, 1218, 222]. deadlock-freedom [1340]. Decidability [1215, 478, 1239, 937, 896, 744, 751, 425, 598]. decidable [934, 883, 1109]. Deciding [43, 260, 529, 334]. decision [86, 124, 186, 320, 681, 1348, 1362, 461, 515, 1140, 812, 1130, 1368, 63]. decisional [679]. Deco [348]. decoding [1341]. decomposable [1276]. decomposition [353, 487, 141, 800, 353]. decompositions [1357]. decorations [1016]. deduction [234, 622, 600]. deductions [1183]. deductive [1104]. default [521, 1322]. Defaults [637]. definability [143, 258]. definable [504, 892]. defined [354, 523, 627, 432, 696, 924]. définies [432]. defining [796]. definition [584, 575, 782, 80, 779, 1013]. definitions [232]. degenerate [794]. degree [389, 1092, 388, 318]. Degrees [633, 163]. delai [848]. delay [1094, 1065, 951, 848, 624, 1003, 952]. delays [653]. deletion [740]. Delta [926]. Delta-grammars [926]. demonic [467, 203]. denotational [666, 1191, 984, 983, 1401, 1400, 1181, 995]. dense [389, 828, 1043]. density [1372, 1270]. Dépendance [331]. dependence [580, 331]. dependences [838]. dependencies [839, 196, 1107, 564]. dependent [106]. depoissonization [1031]. depth [1144, 757, 583, 8, 1121]. depth- [583]. depth-first [757]. derandomization [1323]. derivation [899, 860, 44]. derivations [1187, 918]. derivatives [287, 19]. derived [1367]. Descriptional [688]. descriptions [669, 453, 927]. Design [1345, 703, 328, 1053, 57, 1389, 779]. designs [1122]. DESIR [797]. DESIR-II [797]. Detecting [794]. detection [789, 65, 1218, 79]. detector [1306]. determination [1171]. determining [975]. determinism [79]. deterministe [848]. Deterministic [1216, 548, 105, 889, 848, 690, 297, 255, 766, 492, 161, 1369, 478, 933]. determinizable [1186]. Devaney [584].

Developing [603]. development [1249]. developments [346, 603]. dévéloppements [346]. devising [788]. DFA [649]. diagnosis [1060]. diagonal [1346]. diagram [586]. diagrams [124, 1068, 681]. dictionary [328]. DIEGO [930]. differences [769, 115]. differentiable [1298]. Differential [1146, 795, 793, 314, 794, 317, 798, 796]. differential-algebraic [796]. differentially [113]. diffusion [1151, 1152]. digital [167, 658, 265, 117, 1387, 120, 224]. digraphs [1062, 1055, 1124]. Dijkstra [148]. dimension [1318, 661, 9, 8]. dimension-exchange [1318]. dimensional [819, 270, 1087, 584, 96, 1091, 1330, 1244, 428, 180, 777, 1365, 1041, 558, 1282, 611]. dimensions [1173, 9]. Diophantine [566]. direct [272, 860]. directed [1328, 350, 925, 1058]. directional [56]. discipline [1339]. discovery [6]. Discrete [1283, 1337, 52, 1379, 499, 782]. Discrete-time [1337]. discretization [313]. discretized [102]. discrimination [130]. discussion [1249]. disjoint [444]. disjunctive [372, 517, 1104, 1071]. disk [579]. disorder [483]. dispersal [1145]. distance [880, 689, 55]. distances [1123]. Distributed [1397, 38, 133, 942, 1063, 298, 948, 697, 775, 1307, 1235, 959]. distribution [1029, 178, 433, 1318]. distributions [716, 114, 719]. distributive [455, 1140, 390]. disunification [86]. Divergence [37, 359, 22]. Divide [400]. Divide-and-conquer [400]. division [541]. DLOGTIME [184]. DNA [763]. do [291]. documents [1019]. does [1303]. Domain [221, 843, 1285, 283, 883, 1291, 1167, 543, 526, 360, 637]. Domain-independent [843]. domain-theoretic [1167]. domaine [346]. domains [876, 635, 1234, 677, 963, 1345, 672, 468, 281, 233]. dominating [1147]. Domineering [305]. Domino [355, 662]. dominoes [353]. dominos [353]. don’t [654]. double [819, 241, 758].

Double-pushout [758]. doubly [168].

down [1089, 889, 297, 105, 1017].

downwards [507]. DPO [898]. Drawing [611, 549, 1252]. drawings [1054]. driven [70]. Dual [1383, 1118, 355, 290]. Duality [216, 219, 1048]. d’un [331, 301]. d’une [377, 1156]. duplication [653]. durational [47]. durations [24]. Dynamic [457, 116, 79, 759, 838, 456, 1055, 577, 871, 339, 1016, 26, 1233, 933]. dynamic-data [577]. Dynamical [420, 104, 1034, 19, 500, 20, 1243, 25, 426, 699]. dynamics [1249, 587, 501, 1418].

E-pattern [790]. E-unification [471, 40]. each [1222]. eager [476]. early [155]. ears [487]. easy [1370]. EATCS [344, 440, 542, 736, 737]. Echanges [585]. eco [1153]. eco-grammar [1153]. écritures [1208]. edge [1190, 1063, 549, 708]. edge-colouring [708]. edges [1160, 1056]. edit [97, 689, 712]. Editorial [1404, 1414, 945, 210, 82, 832, 242, 1010]. education [907]. effect [483]. Effective [1285, 268, 1304, 1293, 589]. effectively [890]. Effectiveness [1302]. effects [776]. efficiency [947]. Efficient [949, 951, 1057, 401, 952, 1386, 687, 55, 1039, 168, 1061, 138, 958, 157, 487, 1389]. efficiently [1330]. Ehrenfeucht [581]. election [647]. ELECTRE [146]. electronic [1389, 242]. elegans [1249]. element [807]. elementary [785, 122, 651, 1037, 321, 998]. Elements [526, 883]. Eliminating [962]. Elimination [1221, 476, 993, 1403]. elusive [1354]. embedded [1343]. Embedding [203, 876, 1412]. embeddings [717, 613]. EMPA [1425, 1044]. empty [641, 422]. emulations [289]. Encapsulated [910]. eNCE [787, 1352]. encircle [1222]. Encoding [991, 1341, 611]. encryption [1029]. end [594]. endoscopy [1125]. engendrés [683]. engine [505]. engineering [907]. Enhanced [1180]. Enlargements [1005]. Ensuring [912]. Entailment [873]. Entailment-based [873]. Entropy [1427]. enumerable [1081, 1088, 164, 967]. enumeration [1209, 687]. Enumerative [1325]. environment [923]. environmental [1248]. environments [835, 1001]. equal [252]. Equality [1286, 1033]. equation [793]. Equational [14, 935, 1000, 365, 366, 1149, 1189, 78, 445, 234, 663]. equations [566, 1364, 795, 645, 314, 1374, 794, 486, 125, 629, 1241]. Equivalence [674, 1089, 307, 1220, 1008, 700, 690, 1320, 1398, 202, 453, 790, 478, 430, 1239, 937, 593, 671, 453]. equivalences [836, 254, 751, 676, 260, 39, 529]. equivalent [127]. ERCW [953]. EREW [429]. Ergodic [1085, 1131]. Erratum [735, 1424, 1427]. erroneous [763]. error [1060]. errors [732]. ESM [914]. essence [1331]. Essential [676]. estimates [396]. estimation [236]. étude [1208]. Euclidean [1287]. Euler [63]. evaluation [659, 961, 147, 533, 100, 939, 395, 538, 1172]. even [256, 766]. Event [1186, 667, 248]. Event-clock [1186]. events [948, 32]. Eventually [1310]. Eventually-serializable [1310]. evidence [539]. evolutionary [1415]. evolved [1245]. Exact [1220, 449, 11, 1265, 1428]. example [541, 1223, 4]. Examples [1066, 544, 766]. exchange [1318]. Exchanges [585]. exclusion [398, 247, 879]. executable [670]. execution [156, 506]. exist [291]. existence [125]. existentials [878]. expansions [1385]. Experiments [622]. Explicit [1190, 485, 784, 534]. explicitly [962]. Exploiting [413, 399]. Exploration [1209, 1251]. exponent [69]. exponential [482, 338]. exponential-time [338]. expression [901, 917, 422, 1210]. expressions [977, 1392, 1149, 92, 287, 659, 649, 43, 807, 1116]. Expressive [1104, 885, 194, 844, 193]. expressiveness [941, 277]. expressivity [1231]. extended [92, 1099, 393, 693, 1096, 834, 1116, 340, 527, 575]. Extending [1197]. extension [1061, 643, 1036]. extensional [514, 222]. extensions [1194, 609, 985, 493]. external [843]. extracting [315]. Extremal [174]. eye [1415, 674, 673].

facteurs [451, 545]. factoring [151].

factorization [799, 75, 846]. factorizations [435, 553, 1267, 734]. factors [451, 1098, 1163, 545, 1271, 1262]. facts [6]. fail [181]. failure [1306, 534]. failures [1182]. Fairness [826, 1003, 636]. faithful [225]. families [1218, 787]. family [277, 502]. fans [803]. Fast [16, 289, 1211, 1172, 167, 1095, 511, 427, 54]. fault [1151, 1152, 775, 960]. fault-tolerance [775]. fault-tolerant [1151, 1152, 960]. faulty [429]. Feasible [1378, 166, 1127]. feature [150]. features [1175]. feedback [60, 1384]. few [1327, 413]. fiber [1328]. Fibonacci [1428, 1265, 556]. field [799, 971]. fields [324, 846, 397, 326]. FIFO [1332]. FIFO-channel [1332]. figures [353, 453]. fill [612]. fill-in [612]. filling [167, 703]. Filter [517]. find [1095, 262, 1057]. Finding [558, 554, 166, 138, 1059, 135, 326]. Fine [847, 1268, 1264]. Finite [299, 22, 578, 696, 860, 631, 287, 1020, 1158, 113, 799, 648, 146, 257, 1271, 1377, 624, 1003, 277, 43, 96, 260, 324, 807, 629, 846, 824, 1369, 1350, 1077, 1078, 463, 326, 1281, 1388, 437, 593, 1132, 1261, 1049, 115]. finite-element [807]. finite-state [43, 824]. finitely [1149, 153, 1075, 1023, 1049]. finiteness [1075]. Firing [459, 547, 498]. First [371, 757, 145, 368, 838, 1341, 619, 1232, 794, 369, 1345, 776, 1403, 925, 164, 1142]. first- [1403]. First-order [371, 145, 368, 1341, 619, 1232, 369, 1345, 1142]. Fixed [274, 62, 28, 824, 1228, 206, 152, 663, 1346, 388, 33, 661, 1256]. Fixed-parameter [62]. fixed-point [152]. fixpoint [901, 659, 1003, 370]. fixpoints [514, 626]. Flatwords [378]. Flexible [541, 313]. floating [1276, 955]. floating-point [1276, 955]. flow [222, 181, 1017]. folding [80]. fonctions [131]. Fondement [353]. food [857]. force [971]. Ford [181]. Ford-Fulkerson [181]. forests [75, 1016, 1228]. Foreword [748, 1051, 792, 715, 869, 837, 863, 191, 490, 630, 472, 596, 932, 226, 762]. fork [820]. form [886, 964, 1136, 1173, 939]. Formal [825, 506, 647, 1235, 1269, 809, 672]. formalisation [466, 508]. formalization [1367]. Formalizing [148]. formel [809, 801]. forms [240]. formula [1133]. formulae [1207, 69]. formulas [770]. formulation [462]. Forum [997, 447]. foundation [149, 1288, 753]. Foundations [882, 777, 875]. founded [333, 469, 147, 1363]. four [8]. Fourier [1420]. fourth [178]. fractal [12, 699]. fractals [129]. fraction [887, 1385]. fractions [1069]. fragment [992, 1366]. fragmentary [170]. Fraïssé [581]. frames [219]. framework [410, 948, 771, 959]. Free [234, 406, 1124, 518, 868, 1356, 799, 1313, 1067, 1227, 698, 1007, 888, 1374, 334, 384, 1073, 612, 1270, 677, 454, 974, 1362, 450, 1050, 741, 109, 616]. free-choice [1007]. freedom [1340]. Frege [1184]. French [451, 809, 349, 1208, 1070, 13, 331, 545, 353, 301, 848, 683, 1072, 432, 377, 801, 453, 295, 1035, 131, 585, 1156, 346]. Frequences [451]. frequencies [1161, 451]. Frequency [411]. fringe [101]. front [433]. frontier [259]. Fugitive [552]. Fugitive-search [552]. Fulkerson [181]. Full [1392, 223, 488, 773, 979, 1333, 257, 1021]. Fully [456, 199, 977, 1408, 775]. function [408, 232, 321, 11]. Functional [1194, 145, 632, 902, 839, 887, 1005, 1405, 1408, 196, 1107, 79, 623]. Functional-logic [1194]. functionals [256]. functions [692, 1284, 1356, 392, 124, 1224, 686, 1346, 114, 608, 1382, 1354, 749, 1168, 681, 66, 179, 1059, 325, 394, 131, 8, 1298, 70, 1202, 1426, 843, 415, 1304, 154, 1391, 1172]. functor [206]. Fundamental [543, 59]. Fusion [1225].

GA [1421]. Gabbay [33]. Galmiche [1112]. Game [1338, 977, 1222, 270, 107, 1093, 547, 732]. Game-theoretic [1338]. Games [407, 176, 581, 552, 854, 988, 975, 1402, 590, 1023, 341]. Gap [814, 1053, 1416]. Gap-languages [814]. gaps [368]. gas [808]. GASCOM [345]. gates [321, 815, 583]. GCD [397]. General [1145, 1079, 1249, 632, 643, 1052, 1100, 544, 303, 1395, 248, 879, 323, 164, 370]. generalised [969]. generalization [1355, 1264, 4, 1001, 1184]. Generalized [1029, 876, 1037, 388, 1267, 999, 330, 1216, 890, 638, 526, 821, 9, 1017]. generalizing [102]. generated [819, 652, 683]. Generating [129, 1258, 805, 1352, 349, 114]. generation [347, 348, 1275, 562, 1276, 1043, 807, 1277, 346, 346, 349]. generator [1066]. Genericity [482, 1127]. Genetic [1420, 1416, 1418, 1417, 1022]. genome [1179]. gentle [575]. Gentzen [518, 1201]. genus [1379]. Geometric [337, 310, 298, 703, 313]. geometry [312, 12]. GF [1379]. given [323]. global [1008, 570, 313, 563, 921, 561, 1302]. global-clock-bounded [1008]. Glushkov [1210]. Go [436]. goal [56]. golden [1074]. Graded [540, 1123]. grammar [1230, 1011, 1153, 1227, 907, 458, 918, 1215, 926, 1159]. grammars [1114, 904, 905, 405, 698, 129, 888, 434, 1038, 621, 1352, 1088, 351, 924, 1018, 430, 926, 1177]. grammatical [292, 764]. Graph [886, 521, 915, 1250, 70, 931, 861, 175, 896, 717, 618, 903, 1160, 904, 905, 405, 907, 908, 910, 511, 1038, 710, 94, 1057, 895, 912, 913, 549, 787, 918, 919, 1137, 269, 604, 921, 1178, 71, 992, 923, 53, 924, 339, 1138, 926, 687, 929, 930, 1252, 367]. graph-grammar [926]. graph-oriented [367]. graph-theoretic [992]. graphical [353]. graphique [353]. Graphs [102, 1222, 389, 652, 900, 607, 1143, 551, 1343, 355, 106, 220, 364, 906, 1357, 552, 1043, 505, 910, 706, 402, 693, 1395, 1057, 911, 1213, 1199, 136, 549, 916, 612, 1363, 579, 1278, 450, 996, 1371, 1078, 928, 123, 1058, 1039, 1375, 168, 1023, 341, 708]. greatest [206]. greedy [169, 1311]. grid [958]. grids [156]. Grossberg [340]. ground [757, 469, 1106]. group [680, 1377, 416, 513, 1078]. groupoids [1158]. groups [874, 450, 1141]. growth [122, 1223]. Guthrie [254].

Hajós [435, 734]. Hall [513]. Hamiltonian [102, 1078]. Hamming [722]. Hammock [487]. Hammock-on-ears [487]. hard [895, 1209]. harder [716]. hardness [1323, 1161, 1157]. hardness-randomness [1323]. Harmonic [112]. hashing [130, 726]. having [19]. head [1427, 1242]. Heaps [423, 1317]. heapsort [1317, 414]. heartbeat [1306]. height [682]. Helios [561]. helps [424]. Herbrand [1368]. Hermite [377]. Hermitian [377]. heterogeneous [947]. heuristic [1419]. heuristics [427]. hidden [236, 1077]. hierarchical [910, 956, 231, 302, 971, 1357]. Hierarchically [579]. hierarchies [592]. Hierarchy [1080, 361, 1117, 1166, 938, 633, 850, 1002, 1096, 787, 1221, 189, 338, 589, 1120, 847]. high [1092, 949, 313, 775, 993, 783]. high-level [949, 783]. high-order [313, 993]. Higher [864, 761, 625, 520, 40, 749, 473, 1408, 1136, 208, 1403, 1365, 925]. higher-dimensional [1365]. Higher-order [864, 761, 625, 520, 40, 749, 473, 1408, 1136, 208, 1403]. highest [389]. Higman [1032]. histories [134]. history [27]. holes [353]. holonomic [1172]. homogeneous [795]. Homogeneously [551]. homomorphic [638]. horizontal [354, 288]. horizontally [1279]. Horn [445, 608, 770]. Howard [773]. HRNCE [1352]. hull [597]. hulls [812]. Hybrid [197, 18, 500, 20, 21, 1337, 23, 1161, 26]. hyper [1166]. hyper-arithmetical [1166]. hypercube [58, 661, 400, 891]. hypercubes [330]. hyperedge [351]. hyperelliptic [1379]. Hypergraph [922, 1352, 920]. hyperuniverses [303]. hypothesis [540].

I/O [617]. IBDDs [1080]. ideal [519, 362]. idealized [977, 1392]. ideals [1135, 803]. identical [588, 529]. identification [1316, 818]. If [17, 431]. II [357, 1238, 1062, 62, 1327, 797, 688, 515]. II. [35]. image [262, 57, 1049]. images [295, 1387]. immune [1382]. imperative [1258]. Implementation [798, 201, 1046, 67, 401, 561, 956]. implementations [77, 1313]. implemented [1225]. Implementing [920, 838]. implicants [1207]. implications [442]. Implicit [125, 143]. imply [44]. Improved [1372, 659, 1042, 1147, 856, 1033, 823]. Improving [1329, 559]. inaccurate [385]. incomplete [1107]. increasing [1346]. Incremental [78, 771]. indefinite [533]. independants [1072]. Independent [1350, 140, 1072, 486, 1091, 661, 173, 843]. Index [213, 211, 374, 530, 1413, 1422, 1289, 210, 386, 781, 701, 1206, 1204, 483]. Indexed [212, 802, 1205, 418, 1223]. indexing [1175]. indices [139]. indifference [708]. indistinguishability [42, 859]. individual [1217, 1131]. Induction [1169, 1356, 522, 665, 639]. inductive [1284, 120]. Inductively [25, 523, 627]. inequalities [1192, 643]. inequations [566, 174]. inference [1284, 518, 196, 764, 768]. Inferring [763, 387, 25]. infinis [683]. Infinitary [601, 916, 195]. Infinite [1193, 240, 1023, 1333, 258, 1271, 683, 621, 578, 1030, 824, 924, 1132, 218]. infinite-valued [1030]. Infiniteness [1110]. infix [822]. influences [331]. Information [1017, 1151, 337, 1145, 267, 934, 385, 1152, 1129, 1303]. information-based [1303]. inhomogeneous [731]. initial [177, 233]. input [172, 31, 1083]. input/output [31]. Insertion [740, 1088]. instance [380, 895]. instantiation [279]. institution [1233]. Institutions [576]. instructions [1225]. instrumentation [957]. Integer [966, 1157, 1141]. integers [1]. integrals [115]. integration [1194, 221]. Intensional [232, 464]. intensive [897, 960]. interacting [894]. Interaction [357, 833]. Interactive [875, 1029]. interconnection [156, 1122, 1124]. interference [1412, 1411]. interior [132]. interleaving [1236, 379, 1105, 275]. intermediate [562, 1223]. internal [941, 1081, 480]. International [352, 489]. interpretation [1047, 81]. interpretations [1342, 1399, 1017]. interprocedural [477]. interprocess [201]. Intersection [520, 229]. Intersections [139]. Interval [134, 461, 515, 93, 1084, 569, 1301]. intervalles [585]. intervals [614, 585]. intrinsic [1297]. Introduction [909, 481, 311, 312, 575]. Intuitionism [981]. Intuitionism-counting [981]. Intuitionistic [600, 1196, 360, 1201, 1183, 1360]. invariance [1165]. invariant [1365]. invariants [562, 245, 1377]. Inventories [1036]. inverse [1067, 1214]. inversion [171]. inversions [413]. inward [926]. irrational [1098]. irregular [404]. irreversible [980, 1394, 180]. isomorphic [1068]. isomorphism [438]. isomorphisms [994]. Issue [352, 345]. items [546]. iterated [105]. iteration [1189, 817, 935, 645]. Iterative [527, 343, 453]. iterées [453]. IX [220]. jobs [403, 661]. join [1116]. Joining [1240]. joins [1096]. Joint [886]. Jordan [782]. jump [1286].

Kahn [933]. Karp [431]. Karp-Levin [431]. key [1029, 1388]. keys [839, 296]. kind [81]. Kleene [14]. Kleijn [376]. Knapsack [388]. knowledge [1251, 1307, 537, 721]. known [101]. Kolmogorov [818, 1378, 1027, 844]. Kripke [924].

L [888, 104, 161, 335, 430]. Label [228, 10]. Label-selective [228]. labeled [266]. labeling [549]. Labelled [1083]. labels [88, 1040]. Lambda [504, 45, 143, 88, 978, 1195, 1009, 871, 980, 514, 1005, 601, 1410, 1001, 1202, 1426]. lambda-calculus [88, 978, 871]. lambda-machines [980]. langages [349]. Language [1150, 64, 306, 1316, 145, 406, 825, 632, 130, 146, 227, 405, 1223, 997, 1408, 290, 125, 784, 767, 561, 1034, 381, 488, 1235, 422, 646]. Languages [1158, 77, 217, 293, 185, 1114, 376, 227, 1212, 349, 207, 1081, 12, 418, 669, 425, 660, 384, 742, 277, 740, 914, 5, 1409, 68, 787, 591, 766, 454, 974, 662, 604, 79, 745, 1088, 1320, 849, 335, 790, 940, 513, 1050, 741, 814, 1036, 1015, 721, 847, 382, 1281, 1079, 844, 90, 50, 1097, 259]. Large [396, 1379, 947, 1180]. Large-scale [396]. large-size [947]. largest [468]. Late [155]. latency [955]. latest [797]. Lattice [1247, 851, 1118, 1255, 558, 615]. lattices [1249, 1118, 1381, 1254, 455, 174, 1362, 613]. Latticial [1259]. law [856]. laws [1258, 760]. layer [1253]. layering [74]. layout [718, 1250]. Lazy [476, 805, 222]. leader [647]. learnable [970]. Learning [765, 769, 686, 385, 591, 766, 770, 546, 1148, 424, 544, 410, 7, 5, 64, 179, 290, 767, 11, 771, 813]. leaves [1325]. left [204]. left-linear [204]. lemma [418, 999]. Lempel [117]. Length [1207, 741, 435, 101, 718, 654, 437]. lengths [176, 44]. lenses [509]. letter [1089]. letter-to-letter [1089]. level [949, 783]. levels [1202, 1426]. Levin [431]. lexicographic [619, 44]. Lie [800]. lightwave [1343, 891]. like [651, 1174, 488]. Liminf [756]. Limit [1260]. limited [888]. Limiting [620, 114]. Limits [433]. Linda [870]. Lindenmayerian [746, 747]. line [167, 532, 1324, 606, 1162, 408, 1167, 610, 1169, 60, 179, 394, 1321, 1389]. Linear [706, 982, 890, 811, 454, 1363, 925, 1022, 927, 933, 342, 566, 773, 819, 166, 1155, 978, 1393, 632, 311, 1160, 972, 609, 795, 262, 3, 793, 569, 314, 240, 136, 474, 987, 1397, 766, 989, 1402, 102, 308, 1330, 1410, 1362, 622, 990, 849, 992, 11, 1385, 1403, 998, 1112, 1366, 995, 973, 1085, 204, 796, 1142, 1399, 1049, 964, 364]. Linear-time [706, 136, 474]. Linearity [977, 394]. linearizability [1308]. Linearizable [1315]. linearly [127]. linkage [1178]. links [333]. Liouvillian [795]. List [212, 961, 968, 1205, 74, 1226]. lists [1341, 417, 328, 118]. live [1007]. LLNC [1398]. load [1314]. Local [293, 570, 1070, 789, 1094, 752, 187, 1314, 563, 989, 921, 421, 554]. locales [1070]. Locality [275]. locally [892, 672]. Locating [241]. Location [202, 725]. Löf [523, 627]. log [814]. log-time [814]. logarithmic [1126]. logarithms [1379]. Logic [964, 906, 195, 773, 78, 867, 368, 1101, 148, 663, 1194, 199, 755, 1045, 572, 35, 577, 200, 220, 364, 1357, 994, 521, 40, 678, 505, 827, 1100, 982, 3, 1361, 997, 473, 474, 834, 369, 987, 1397, 24, 989, 1402, 1319, 79, 1345, 1362, 776, 623, 990, 447, 535, 48, 992, 1403, 576, 777, 582, 34, 998, 940, 1112, 1366, 995, 80, 884, 461, 515, 771, 534, 446, 1367, 1399, 714, 370, 527, 1235, 1322]. Logical [595, 467, 258, 757, 572, 25, 927]. Logics [595, 372, 934, 512, 1359, 1048, 189, 46, 750, 461, 515, 463, 598]. LogP [1214]. LogP-machine [1214]. logs [1327]. LOGSPACE [511, 1409, 604]. long [1136]. longest [110]. look [297, 105]. look-ahead [297, 105]. lookahead [968, 1162]. loops [675]. Loveland [1378]. low [337, 680, 815, 1096, 775]. Lower [319, 812, 1144, 1148, 309, 484, 543, 321, 108, 291, 557, 1164]. lower-bound [484]. Luby [178]. Lukasiewicz [1030]. Lyndon [656, 1262].

m [653]. MA [17]. Machine

[829, 1214, 396, 495, 6, 950, 967]. Machines [489, 172, 1290, 650, 257, 980, 1394, 690, 96, 7, 1173, 494, 492, 587, 1427, 493, 491, 1242, 685, 724, 220]. made [1370]. magic [651, 81]. majority [921]. Makanin [853]. malicious [769]. malignness [719]. management [913]. Mandelbrot [713]. Many [778, 553, 529]. Many-to-many [778]. map [666, 1191, 1249, 896]. MAPLE [808]. Mapping [1293, 57]. mappings [817]. maps [572, 836, 1173]. Marking [114]. Markov [320, 1294]. Martin [523, 627]. Martin-Löf [523, 627]. Massively [404, 553, 401]. Master [701, 374, 530]. Matching [1195, 654, 16, 336, 408, 1272, 157, 264, 60, 712, 1041, 996, 1321, 684, 646]. matchings [1043, 778, 140]. Mathematical [1128, 590, 665, 107, 6]. mathematics [1288, 806]. Matos [72]. matrice [377]. matrices [682, 969, 821]. matrix [1095, 460, 1066, 694, 377]. Max [1370, 569]. maxima [1059]. Maximal [1266, 1279, 1065, 106, 375, 822, 135, 173, 1211, 154]. Maximum [720, 166, 437, 181, 709]. May [572, 181]. McNaughton [61]. mean [955, 341]. meaning [233]. means [921, 1088, 90]. measure [482, 268, 333, 132]. measures [1127, 1301]. mechanism [867]. median [1213, 1123]. meet [851]. Mellin [112, 115]. membership [434, 1038, 721, 1154]. memory [399, 452, 956, 50]. men [1222]. Merit [1163]. meshes [404, 966]. message [1313]. message-passing [1313]. Meta [466, 227, 910]. meta-language [227]. Meta-resolution [466]. metaphysics [1128]. method [1323, 127, 1063, 157, 775, 323, 956]. methodology [642]. methods [484, 1348, 470, 317]. Metric [24, 519, 152, 876, 1045, 877, 635, 750, 1302, 1304, 1049]. metrical [893]. metrics [224]. migration [759]. min [569]. mind [1316]. Minimal [101, 1194, 1226, 818, 878, 279, 1270, 135, 585, 687]. minimales [585]. Minimality [629]. minimizing [1155, 852]. minimum [1372, 1108, 718, 706, 417, 895, 160, 612, 452]. Mixed [15, 236, 1064, 1039]. ML [1197, 575]. mobile [1236, 1336, 275]. mobility [941, 480]. modal [216, 938, 1333, 1048, 1361, 46]. modality [599]. Model [1333, 924, 670, 1061, 896, 752, 200, 877, 946, 358, 1214, 1030, 1182, 1202, 1426, 225, 340, 304, 367, 390, 1303, 26]. Model-checking [924]. modeling [561]. Models [840, 525, 143, 517, 1104, 897, 233, 464, 279, 147, 91, 25, 712, 29, 1299, 159, 193, 284, 967, 310]. modes [172]. modified [817, 689]. Modifying [1056]. Modular [107, 1137, 231, 1138, 204, 753, 1100, 49, 1078]. Modularity [235, 865]. module [645, 930]. modules [465, 1141]. modulo [245, 1364, 548, 583, 558]. modulo-invariants [245]. modulus [1302]. Moez [142]. molecular [1178]. moment [178]. Monad [599]. Monadic [1344, 623, 463, 220, 364, 1357, 890, 884]. monoid [1374, 153, 1135]. monoids [1067, 207, 1073, 1135, 1075, 860]. monotone [1354, 11, 1164, 128]. Monotonic [290, 755]. monotonicity [767]. morphismes [683]. morphisms [1373, 831, 683, 936, 657]. Morse [1163]. Most [164, 156, 1160]. mot [1070]. motifs [1156]. motion [51, 313]. mots [349, 545, 683]. move [433]. move-to-front [433]. movement [897, 452]. moves [1093]. Moving [857, 51, 1427]. MPA [1189]. mu [938, 1333]. mu-calculus [938, 1333]. Multi [74, 68, 535, 777, 201, 826, 1421, 512, 874, 1018]. multi-agent [826, 874]. multi-axiom [1018]. Multi-dimensional [777]. Multi-list [74]. multi-niche [1421]. Multi-pattern [68]. multi-pebble [512]. Multi-valued [535]. multi-writer [201]. multichannel [891]. Multidimensional [1084, 969]. Multilayered [1246]. multiparty [855]. Multiple [264, 684, 733, 278, 1180, 1362, 447]. multiple-conclusion [447]. multiplication [1095, 1383]. multiplicative [992, 1360]. multiplicatives [996]. multiplicity [1120]. multiply [44]. multirate [729]. multisearch [1061]. multiset [130, 246, 872]. Multisets [1188]. multivariate [799]. Mutual [398, 247, 879].

N [682]. N-rational [682]. name

[941, 1423, 1410, 1339]. name-passing [941]. named [714]. narrowing [911, 476]. Natural [84, 622, 600]. nature [32]. NC [106, 314, 1211]. NCAS [389]. nd [1000]. nd-order [1000]. Near [1063]. Near-optimal [1063]. nearest [505]. nearly [586, 1058]. need [1410]. negation [867, 757, 200, 534]. Negative [373, 278]. neighbor [505]. nested [167, 769, 197]. net [785, 901, 67, 524]. Nets [244, 2, 735, 1231, 162, 752, 618, 905, 1347, 1218, 1007, 1046, 833, 248, 188, 260, 988, 249, 250, 1105, 996, 671, 26, 1360]. network [1056, 960, 304, 1124]. networks [1306, 156, 1324, 1312, 1343, 947, 982, 278, 729, 499, 160, 502, 563, 1311, 1308, 1161, 103, 954, 582, 891, 1090, 181, 933]. neural [2, 735, 651, 499, 59, 103]. neural-like [651]. Newton [561]. NFA [649]. ngsm [593]. nibble [1063]. niche [1421]. niching [1421]. Nim [107]. NLC [1038]. No [716, 282, 1377, 153, 421, 1386]. nodes [1325, 1329]. noise [769]. Noisy [768]. nombre [331]. Non [1144, 595, 1236, 1278, 1132, 755, 903, 1372, 1068, 469, 553, 240, 987, 454, 1245, 1159, 370, 1106]. non-approximability [1372]. Non-commutative [1144, 987]. non-commuting [553]. non-ground [469, 1106]. Non-interleaving [1236]. non-isomorphic [1068]. non-linear [240]. non-monotonic [755]. non-orthogonal [903]. non-returning [1159]. non-safe [370]. Non-Standard [595]. Non-stochastic [1132]. Non-uniform [1278, 1245]. nonblocking [729]. noncommutative [979, 992]. noncopying [235]. Nondeterminism [218, 631, 1044, 1425, 359, 711]. Nondeterministic [1020, 650, 91, 467, 96, 1408, 61]. nondistinct [296]. Nonholonomic [316]. nonholonomy [318]. nonlinear [108]. nonlinearity [1391]. nonmonotonic [372, 536]. Nonnegative [682]. nonoverlapping [138]. nonprimitive [266]. nonreachability [245]. nonrecursive [132]. nonregular [1332]. Nonreturning [458]. nonsubsequences [110]. nonsupersequences [110]. Nonsymmetric [36]. Nontrivial [1354]. nonuniform [1296]. nonzero [1155]. Normal [655, 240, 1136]. normalization [1190, 1347, 3, 993, 1403, 1112]. normed [334, 937]. norms [815]. notation [276, 1001]. Note [1355, 677, 416, 744, 292, 53, 926]. Notes [242, 642]. notions [405, 431]. NP [17, 205, 785, 716, 460, 127, 1370, 171, 1007, 66, 431, 1209, 1035, 128, 159, 261, 159]. NP-complete [785, 171, 1007, 1035]. NP-completeness [460, 127, 1370]. NP-complets [1035]. NP-hard [1209]. Number [344, 440, 542, 736, 737, 284, 1207, 1171, 609, 331, 1265, 1428, 1303]. numbering [13]. numberings [818]. numbers [1098, 1335, 393, 1292, 1134, 396, 1157, 1170, 655, 225]. numeration [704, 1072, 13, 331, 1072]. numerical [1257, 620].

O [617]. OBDDs [1080]. object [227, 134, 943, 759, 367, 193, 50, 198]. object-based [193, 50, 198]. object-oriented [227, 50]. Objects [1412, 167, 234, 1275, 1195, 192, 1102, 421, 1315, 626]. observability [574]. observable [2, 735]. observation [219]. Observational [42, 307, 43]. observed [320]. obstacles [51, 54]. obtain [1056]. of- [879]. off [1389]. off-line [1389]. offs [1323]. On-line [532, 1324, 606, 610, 60, 179, 1321, 1162, 408, 394]. once [69]. one [319, 819, 1087, 371, 584, 707, 856, 1173, 1244, 180, 73, 306, 1142, 1242]. one-dimensional [819, 1087, 584, 1244, 180]. one-head [1242]. one-step [1142]. one-way [319, 73, 306]. Online [1298]. only [1225, 886, 964, 1280, 452]. Open [836, 1006, 513, 1297]. operation [851, 1234, 1116]. Operational [1181, 1399, 1028, 144, 700, 359]. operations [1020, 274, 1096, 322, 182, 845]. operator [541, 1342, 33, 369]. operators [887, 574]. opfibrations [898]. opponent [1222]. opt [1180]. optical [1324, 953, 954, 1124]. Optimal [692, 156, 1343, 1328, 414, 271, 1318, 96, 732, 810, 1027, 821, 123, 1375, 357, 457, 1061, 586, 658, 980, 1394, 1063, 1180, 439, 695, 1091, 974, 975, 963, 498, 554, 1210, 1062]. optimally [1307]. optimisation [169]. Optimization [1402, 205, 333, 1146, 147, 807, 561, 340, 1125]. optimized [118]. oracles [323, 768]. ORDAL [1253]. Order [670, 749, 145, 368, 520, 93, 148, 1341, 408, 619, 220, 364, 1357, 40, 371, 313, 795, 1232, 793, 1377, 473, 794, 1408, 1136, 369, 987, 1345, 776, 864, 208, 993, 1403, 1000, 1366, 761, 625, 80, 884, 925, 164, 488, 1142, 616, 780]. Order-sorted [670, 749, 780]. orderable [551]. ordered [1054, 681]. ordering [187, 1136, 1363, 308, 85, 714]. orderings [619, 364, 1136, 44]. Orders [1199, 606, 580, 610, 1025, 1073, 613]. Ordinal [1316, 1032, 449]. ordinals [299]. ordinary [314, 798]. oriented [227, 367, 50]. orthogonal [903, 770]. other [1275, 20, 1057, 487, 1050, 1368]. outfix [822]. output [1083, 1120]. outward [926]. overgeneralization [64]. overhead [957, 775]. overloading [628].

P [1093, 139, 261]. P-positions [1093]. P-selectivity [139]. PAC [546, 1148, 544]. PAC-learning [544]. packing [720, 427]. page [707]. Paging [707, 1162]. Painlevé [798]. pairs [1083]. pairwise [1178]. Palindromes [1349, 65]. palindromic [817]. paper [72]. paperfolding [329]. papers [886, 964]. parabolas [1283]. paradigm [789]. Parallel [65, 900, 93, 667, 375, 694, 718, 1087, 397, 182, 395, 783, 616, 1242, 347, 1092, 156, 1061, 586, 949, 825, 16, 1062, 1160, 1227, 658, 1052, 506, 402, 328, 404, 1002, 553, 401, 675, 265, 66, 1091, 487, 962, 1217, 974, 413, 1385, 1350, 996, 812, 173, 123, 950, 327, 168, 724, 1210]. parallel/distributed [724]. parallelisation [353]. parallelism [633, 1350]. parallelization [841]. Parameter [1356, 62]. Parameterized [850, 158, 522, 264]. parameters [552, 116, 1221]. parametric [706, 404, 202]. Parametrization [102]. Parikh [1050]. Park [639]. parse [1016]. parsers [454, 1018]. Parsing [1014, 384, 117]. Part [1062, 461, 274, 175, 1327, 515, 365, 366]. Partial [287, 1268, 1102, 234, 1143, 758, 1238, 1364, 1104, 438, 1169, 279, 866, 1300, 714]. partially [1312, 320, 962]. participant [1294]. partition [175, 463, 1039]. partitionable [1306]. partitioning [428]. partitions [610]. passing [652, 941, 1313, 480]. passive [1392]. past [189]. path [375, 619, 101, 1136, 123, 15, 1058, 1211, 44, 1375]. Paths [88, 1062, 487, 54, 109]. Pattern [122, 263, 408, 5, 68, 790, 1041]. patterns [765, 1099, 264, 1091]. Pavage [353]. pavages [295]. pavants [432]. paved [432]. payoff [341]. PC [1230, 458, 1159]. PCF [393, 1169, 1109, 1239]. PDL [379]. pebble [512]. distributed [724]. guarantee [474]. non-stochastic [236]. output [31]. SEMAGRAPH [886]. Transition [249]. without [296]. write [1315]. Penelope [895]. peptide [684]. Perfect [726, 996, 140]. Performance [960, 1220]. period [176, 1070]. périodes [1070, 1070]. Periodic [1, 103]. Periodicities [1086]. Periodicity [1074]. periods [1264, 1070]. permanent [1164]. permanents [1209]. Permutability [1196]. permutation [347, 518, 1377]. Permutations [652, 348, 1083]. permutative [1201]. perpetual [1198]. Perrier [1112]. Perrin [336]. person [1222]. perspective [32]. Petri [1231, 752, 618, 905, 1347, 1007, 1046, 248, 188, 249, 250, 1105, 524]. PH [338]. Phase [993, 1403]. phenomena [808]. PHRASE [1018]. physics [1247]. pi [944, 751, 246, 1188]. pi-calculus [944, 246, 1188]. picture [662, 1281]. pictures [167, 353]. piecewise [19, 660]. piecewise-constant [19]. pile [1087]. Pisot [331]. Place [249]. Place/Transition [249]. placement [457]. plan [301]. Planar [607, 355, 106, 1218, 706, 1057, 123]. planarity [1353]. plane [353, 549, 975, 54, 318]. planes [353]. planning [51, 313, 54]. player [1222]. point [206, 152, 274, 356, 1276, 33, 824, 955, 1256]. pointfree [360]. points [663, 1346, 28, 796]. polygon [586]. Polygonal [1170, 54]. polygons [53, 55]. polyhedral [51]. polyhedric [963]. polyhedron [51]. POLYLOGTIME [184]. Polymorphic [1013, 1197]. polymorphism [143].

Polynomial [1398, 513, 17, 1117, 333, 1347, 434, 1038, 1377, 1110, 334, 95, 132, 1221, 377, 1363, 803, 100, 846, 397, 866, 1121]. polynomial-size [17]. polynomial-space [1110]. polynomial-time [1117, 132]. polynomials [799, 690, 553, 324, 969]. polynomiaux [377]. Polyomino [165]. Polyominoes [352, 354, 348, 288, 1279]. popular [156]. posets [611]. positional [590]. positions [1093]. positive [278, 766]. possibility [1111]. Post [378]. pour [1035]. Power [1038, 728, 1269, 876, 185, 500, 885, 1104, 187, 1234, 815, 194, 746, 369, 805, 583, 1016, 844, 193, 685]. powerset [1342]. PP [680]. Practical [721, 108]. practice [357, 929, 1252, 780]. PRAM [429, 897, 946, 289, 1090]. PRAMs [291, 953]. precedence [403]. Precedences [77]. Precise [477]. Predicate [208, 148, 371, 1367]. prediction [1177]. Preface [352, 571, 702, 605, 183, 215, 345, 1203, 503, 531, 253, 76, 560, 640, 111, 243, 739, 772]. Preference [1257]. prefix [16, 409]. prefix-matching [16]. presence [769, 385, 961]. presentation [153]. presentations [744, 332, 386]. presented [153, 1075]. preservation [760]. preserve [890, 719]. preserving [613, 1115, 1202, 1426]. primal [1118, 621]. prime [1207, 281]. primitive [881, 417, 853, 381, 326]. primitives [870]. primitivity [425]. principle [589]. Principles [192]. priorities [1044, 1425, 828, 942, 1028]. prioritized [29]. priority [727, 1083, 182, 1139]. Probabilistic [705, 350, 29, 119, 767, 351, 237, 319, 690, 91, 841, 1119, 141, 537, 617, 538, 1334]. probabilities [1044, 1425, 296]. probability [1127, 1301]. probe [100]. problem [968, 1089, 1092, 861, 785, 1376, 460, 1118, 106, 1042, 1108, 254, 171, 505, 1006, 434, 1038, 1071, 559, 459, 160, 153, 921, 498, 790, 713, 52, 304, 796, 1039, 723, 1125, 1375, 1082, 378]. problem-solving [304]. problèmes [1035, 346]. problems [550, 680, 205, 532, 86, 716, 169, 175, 186, 130, 1066, 1055, 127, 388, 1146, 718, 706, 897, 710, 816, 1248, 425, 188, 1024, 549, 324, 487, 725, 1123, 679, 428, 1215, 1209, 614, 1035, 1297, 339, 927, 812, 87, 173, 123, 128, 340, 593, 1368, 346, 709]. procedure [4, 1106, 181]. procedures [149, 1362, 461, 515, 302]. Proceedings [1253]. Process [644, 914, 249, 939, 646, 828, 902, 870, 942, 634, 643, 751, 31, 647, 516, 913, 230, 986, 874, 280, 641]. process-calculus-based [874]. Processes [673, 1340, 38, 1044, 1425, 320, 1333, 1236, 588, 1336, 47, 334, 918, 1237, 1102, 41, 589, 275, 524, 925, 237, 937, 641, 529, 1049]. processing [130, 262, 1015, 57]. processor [160]. processors [402, 653]. product [667]. productions [926]. Products [257, 860]. Program [598, 835, 479]. programmed [913]. Programming [1045, 77, 78, 949, 902, 755, 192, 35, 754, 1405, 997, 1409, 1397, 919, 842, 567, 535, 576, 777, 923, 1015, 1016, 365, 366, 527, 1235, 218]. programs [445, 867, 1101, 825, 199, 753, 200, 467, 521, 827, 1100, 1258, 620, 962, 623, 208, 696, 943, 80, 771, 557, 960, 446, 1300, 370]. progress [753]. projections [354, 288]. projective [786]. Prolog [757, 508]. Proof [574, 441, 825, 1224, 700, 3, 820, 660, 528, 1110, 988, 1237, 989, 1402, 1112, 996, 1106, 1241, 1360]. proof-nets [1360]. Proofs [928, 1196, 209, 61, 993, 1403, 231, 721, 44]. propagation [1331, 1070, 477]. proper [361, 15]. proper-path-width [15]. Properties [145, 1119, 441, 1004, 903, 753, 1187, 839, 14, 935, 1038, 89, 444, 104, 432, 1267, 1111, 921, 1277, 849, 786, 940, 1304, 1154]. property [217, 1025, 273, 1115, 1256]. Propos [1274]. propositional [361, 518, 644, 730]. Propriétés [432]. protection [1387]. protein [1176]. protocol [958, 647]. protocols [1309]. provably [1356]. Proving [442, 245, 89, 1111, 479, 127, 1103, 4, 1141]. proving-by-example [4]. pseudo [311]. pseudo-linear [311]. PSPACE [1366]. PSPACE-complete [1366]. PTIME [1409, 1365]. public [1388]. public-key [1388]. puissances [331]. pullback [900]. pumping [594]. pursuit [94, 975]. push [1017]. push-down [1017]. Pushdown [711, 1219, 1119, 937]. pushout [241, 1238, 918, 758]. PWD [104]. PWD0L [425]. PX [985].

Qu [508]. Qu-Prolog [508]. quadratic [817]. quantum [690]. Quasi [342, 161, 281, 868, 1025, 1073, 295, 796, 295]. Quasi-affine [295]. Quasi-deterministic [161]. quasi-free [868]. Quasi-linear [342, 796]. Quasi-prime [281]. quasilinear [190]. quasilinear-time [190]. Quasiperiodicity [1273, 1326]. Queries [565, 765, 1117, 411, 541, 686, 134, 66, 11, 537, 55, 843]. Query [959, 538, 147, 533, 462, 844]. Querying [372]. question [1149, 75]. questions [744]. queue [648, 946, 182]. queue-read [946]. queue-write [946]. queueing [121]. queues [727, 1083, 731]. quickly [507]. quiescent [1306]. quorum [775]. quorums [958]. quotient [1260, 1047]. quotients [1078].

R [996]. Rabin [177, 61]. radiosity [956]. radix [1170]. Ramsey [543]. Random [1275, 403, 121, 1419, 348, 1207, 1381, 839, 349, 1276, 1043, 856, 1134, 1278, 1295, 950, 1131, 346, 171]. randomization [312]. Randomized [330, 893, 1130, 511, 255, 399]. randomness [550, 1323, 1127, 1128]. ranking [716]. Rapid [1314, 254]. rate [267]. rates [433, 1417]. ratio [1074, 713]. Rational [475, 1269, 217, 682, 1325, 830, 349, 679, 131, 1157, 90, 1097]. rationnelles [131]. rationnels [349]. Rauzy [426, 699]. Reachability [19, 1007, 187, 1332, 618, 671].

Reaching [56]. reactive [146, 46, 1319]. read [172, 946, 69, 1315, 452]. read-modes [172]. read-once [69]. read-only [452]. read/write [1315]. reading [557]. real [1284, 392, 1335, 1167, 393, 1292, 1168, 1169, 396, 1034, 515, 225, 73, 306, 714, 1303, 1169]. real-time [1034, 515, 714]. real-valued [1284, 392]. realizability [89]. realization [505, 469, 638]. realizing [323]. reals [1168, 412, 391]. rearrangements [1179]. reasoning [755, 753, 1367]. received [886, 964]. Recents [346]. receptiveness [1339]. reciprocal [1194, 295]. réciproques [295]. recognition [93, 136, 974, 1034, 109, 616]. recognizability [704, 890]. Recognizable [207, 662, 1281, 306, 137]. recognized [1158]. recognizer [1282]. recognizers [1034]. Recognizing [1213]. Reconciling [635]. reconfigurable [966, 1090]. reconfiguration [972, 973]. Reconstructing [288]. recovery [845]. rectangular [1337]. recurrence [1260, 705]. recurrent [502]. Recursion [391, 415, 1197, 881, 417, 1169, 223, 479]. recursion-based [479]. Recursive [392, 103, 1284, 150, 901, 1032, 710, 157, 1168, 1136, 853, 119, 44, 509]. Recursively [967, 1081, 1088, 164]. reduced [1234]. Reducibility [294, 1424]. reducible [693, 136]. Reducing [955]. reduction [1144, 443, 1009, 230, 1199, 915, 920, 377, 835, 829, 151, 931, 377]. reduction-based [230]. reductions [342, 184, 357, 254, 917, 1378]. Redundancy [83, 399]. Reference [1205, 212]. refined [1202, 1426]. refinement [901, 904, 1046, 48]. Refining [872]. reflexive [283, 883]. register [650, 494, 1052]. registers [201, 1384]. Regular [549, 422, 238, 1149, 92, 287, 185, 355, 649, 297, 153, 591, 940, 847, 672, 592, 1210]. regularity [1115]. regulated [888]. rehashing [289]. related [552, 188, 397, 812, 1368]. Relating [1135, 1121]. Relation [363, 1220, 1118, 203, 358]. Relation-algebraic [363]. Relational [71, 1020, 541, 570, 116, 569, 779, 538]. relations [166, 1155, 392, 705, 881, 14, 548, 196, 638, 784, 1107, 679, 1157, 390]. relationship [813]. Relationships [426, 323]. relative [970, 1234]. Relatively [1168]. Reliable [1312, 429, 1306, 1160]. relying [721]. remained [1077]. remark [72]. Remarks [888, 998, 990, 1130]. Removing [543]. repetition [1270, 327]. repetition-free [1270]. replacement [351, 783]. replica [958]. replication [246, 1188]. Representability [672, 820]. representable [631]. Representation [580, 883, 5, 161, 1369, 45, 1027, 615]. representations [93, 1285, 730, 1257, 1255, 664, 1170]. represented [1141]. Representing [523, 627, 1183, 124, 1081, 681]. reproduction [497]. Reset [555, 1347]. residue [1171]. Resolution [894, 1030, 1103, 466, 1035, 780, 370]. Resolution-based [894]. Résolutions [1035]. Resource [550, 380]. resource-bounded [380]. resources [457, 828]. respect [1106]. restricted [1059]. Restriction [1391, 685]. Restrictions [175, 664]. result [632]. resulting [817]. results [1230, 162, 1372, 512, 61, 1115]. retrospective [931]. returning [458, 1159]. reversal [685]. reversal-bounded [685]. reversals [1179]. reversed [590]. Reversible [99, 980, 1394, 180, 459, 591, 497, 495]. Revision [842, 527]. revisited [705, 695, 1411]. Rewrite [1136, 867, 444, 890, 864, 602, 239]. rewriter [356]. Rewriting [886, 92, 898, 520, 900, 522, 903, 514, 833, 911, 912, 913, 528, 43, 240, 916, 1352, 1137, 235, 920, 1135, 1113, 922, 923, 1350, 231, 1138, 1139, 204, 929, 1142, 44, 603, 1028]. rewritings [71]. Rhine [791]. Rice [115]. rich [845]. Riemann [1293]. right [90]. rigid [471, 1344, 1368]. ring [1377]. rings [126, 1382, 1314, 1135]. risk [1060]. RNA [1177]. RNC [1211]. Road [1418]. roadmap [51]. robust [358]. role [50]. Ronald [1113]. root [315]. roots [326]. rotation [816]. rotations [1150, 1280]. rotator [1124]. rough [678]. round [133]. Routing [788, 954, 1324, 951, 1328, 330, 1084, 1053, 952, 778, 1321]. Royal [1418]. Rozenberg [376]. RPO [85]. RSA [1380, 1389]. RSA-based [1389]. RSA-type [1380]. Rudimentary [881]. rule [651, 1037, 433, 536, 370]. rules [600]. RWCA [791].

s [1404, 1414]. safe

[162, 1007, 260, 879, 370]. safety [753].

Safra [61]. SAGBI [1377]. Salomaa [1149]. Sample [1148]. sand [1087]. sans [353]. SAT [1348]. satisfaction [78, 473]. Satisfiability [614, 532, 284, 127]. Saving [336]. SBTA [669]. SBTA-languages [669]. Scalable [1390]. scale [396]. scales [724]. scan [167]. scan-line [167]. schedules [1343]. Scheduling [402, 1217, 1311, 1214, 724, 398, 961, 661, 963, 858, 653]. schemata [598, 1014]. schematized [621]. scheme [958, 117, 661, 1033, 879]. schemes [456, 267, 1029, 1084, 1389, 788]. Sci [1425, 735, 1428, 1424, 1112, 761, 1426]. science [1185, 595, 242]. scope [1183]. score [138]. Scott [283, 883]. search [1054, 309, 552, 433, 710, 265, 117, 1091, 296, 118, 989, 810, 120, 554, 1419]. searching [732, 821, 15]. second [148, 220, 364, 1357, 795, 987, 1366, 884, 488]. second-order [148, 220, 364, 1357, 987, 1366, 884]. secondary [1317]. secret [456, 267]. Secure [855]. See [509]. segmentation [658]. SEGRAGRA [886]. Selection [452]. selective [342, 228, 294, 1424]. selectivity [139]. Self [497, 728, 789, 129]. self-affine [129]. Self-reproduction [497]. self-stabilization [789]. self-stabilizing [728]. selling [1386]. SEMAGRAPH [929]. semantic [445, 1403, 26]. Semantical [1198]. semantically [845]. Semantics [624, 1139, 977, 757, 177, 668, 898, 666, 1191, 144, 1423, 902, 199, 200, 1236, 827, 246, 359, 1100, 984, 1401, 469, 230, 248, 620, 155, 1408, 805, 1181, 835, 363, 1320, 249, 535, 636, 993, 995, 1105, 275, 370, 1028, 214]. Semi [224, 1055, 890, 866, 860, 87]. semi-direct [860]. semi-dynamic [1055]. Semi-metrics [224]. semi-monadic [890]. semi-Thue [866, 87]. semidecision [149]. semigroup [548, 696]. semigroups [1066, 744, 486]. semilattice [203]. Semirings [908]. semisimple [800]. sense [729]. sensitive [917, 537]. separable [786]. Séparateurs [683]. Separating [431, 338, 421]. Separation [589]. separations [388]. separators [687, 683]. sequence [733, 158, 763, 684, 1156]. sequences [819, 1325, 1150, 1260, 817, 1098, 170, 1208, 1163, 1271, 747, 432, 1180, 856, 1378, 1178, 135, 585, 1131, 1132]. sequent [361, 1196, 1362, 991]. sequential [443, 852, 1333, 413]. sequentiality [1139]. serial [1383]. serializable [1310]. series [1269, 682, 113, 451, 1160, 402, 746, 996, 1016, 1097]. series-parallel [996]. server [731]. services [1310]. set [206, 169, 106, 1071, 277, 322, 475, 654, 1033, 193, 1124]. set-system [322]. Sethi [293]. sets [337, 342, 1332, 1289, 375, 1147, 678, 523, 627, 970, 294, 1424, 621, 604, 1363, 629, 1, 513, 386, 173, 1281, 592, 140, 1211, 1301, 1304, 284]. setting [202]. seven [629]. several [113, 314, 1180, 160, 1130]. shapes [408]. shared [828, 399, 956]. shared-memory [956]. Sharing [38, 977, 145, 456, 267, 444, 1395]. sheaves [89]. Shellsort [823]. shift [408, 690, 1384]. shift-equivalence [690]. shifted [1385]. shifts [1037, 332]. short [375, 1056, 820, 976]. shorter [1241]. Shortest [1062, 95, 858, 1058, 1118, 487, 1157, 558, 109, 55, 110]. shrinkage [69]. shrinking [418]. Shub [967, 896]. Shuffle [965, 406]. SI [666, 1191]. Sign [1171]. Signals [1244, 644]. SIMD [506, 57]. similarities [1178]. similarity [460]. Simon [660]. Simple [602, 862, 825, 53, 924, 501]. simplification [568]. simplified [57]. simply [528]. simulate [1173]. simulated [458, 1209]. Simulating [987, 61, 1090, 1159]. simulation [236, 99, 1248, 96, 180]. Simulations [1371, 399, 971]. Simultaneous [1368, 471, 1344, 807]. single [179, 731, 918, 1238]. Single-pushout [1238]. singleton [1149]. singular [1380, 1262, 796]. Singularity [1229]. Situated [568]. six [1354]. Size [681, 1144, 1148, 17, 124, 947, 543, 486, 163, 1154]. sizes [970]. skip [328, 118]. SL-systems [365, 366]. slack [566]. SLD [1187]. SLD-derivations [1187]. SLDNF [370]. Slender [335]. sliding [436]. Smale [896, 967]. Small [494, 492, 491, 722, 1062, 431, 419]. smaller [543]. smallest [1060, 1377, 181]. smooth [404, 179]. sober [219]. Sobolev [1305]. Sofic [332]. Software [806, 907, 913]. soils [1248]. solution [205, 1376, 487]. solutions [1095, 1207, 795, 174, 125, 498, 1241]. Solvable [680, 133, 153, 555]. solve [130, 173]. solvers [868]. Solving [1364, 569, 784, 340, 566, 1042, 314, 304]. Some [1230, 1192, 1004, 903, 854, 1255, 679, 990, 1115, 87, 176, 692, 186, 1146, 709, 116, 401, 188, 324, 350, 1277, 397, 1297, 557, 9, 699, 593]. sorted [670, 749, 750, 821, 780]. Sorting [265, 413, 452]. Soundness [1106]. source [1160]. sources [655]. Space [703, 319, 880, 138, 1249, 1285, 1287, 99, 1272, 597, 157, 1110, 412, 497, 244]. space-bounded [99]. Space-filling [703]. Spaces [1358, 152, 880, 219, 876, 651, 877, 1291, 635, 984, 1401, 10, 1361, 475, 672, 224, 1302, 1304, 1305, 1049]. Spanning [709, 706, 1329, 1278, 1228]. sparse [693, 1057, 324, 308]. special [55, 352, 345]. speciation [1421]. specific [198]. Specification [577, 919, 1340, 236, 1231, 831, 506, 1240, 470, 447, 1019]. specifications [77, 670, 234, 42, 648, 522, 749, 574, 233, 474, 940]. specified [296, 579]. Specifying [1322, 1039]. spectra [371, 1199]. spectrum [1209]. speed [811, 399]. speed-up [811]. Speedup [947]. Splicing [1174, 496]. squad [459, 498]. square [799, 427]. square-free [799]. squares [1263, 1265, 1428, 556, 427]. Stability [1182]. stabilization [789]. Stabilizing [1309, 728]. Stable [1021, 200, 469, 154, 468]. stack [650]. stage [607, 729]. Standard [595, 657, 936, 1367, 1197]. standardization [999]. star [682]. star-height [682]. starting [858]. state [977, 236, 257, 753, 43, 25, 103, 824]. state-based [753]. statecharts [1237]. statements [1240]. Static [302, 457, 1233]. Statistical [1418]. steady [103]. Steiner [1228]. step [838, 930, 1142]. steps [862]. Sticker [1076]. stochastic [236, 1220, 1378, 1132]. stochastic/non [236]. stochastic/non-stochastic [236]. Stochastically [300]. storage [1317, 399]. strange [1037]. strategies [581, 262, 1198, 975, 835, 732]. strategy [443, 57]. strict [938]. Strictly [1054]. Strictly-upward [1054]. string [1189, 65, 16, 336, 1042, 1272, 157, 439, 556, 1273, 60, 712, 1113]. string-matching [1272, 157]. string-rewriting [1113]. string-to-string [1042]. strings [265, 654]. Strong [460, 35, 1190, 1139, 476]. Structural [144, 430, 1188]. Structure [822, 252, 1340, 1126, 887, 455, 248, 1024, 53, 1078, 70, 615, 1177, 1304, 1305, 743]. structured [574]. structures [703, 868, 1020, 1011, 667, 572, 1276, 521, 1259, 266, 1257, 986, 924, 1174, 120, 1184, 509]. structuring [904]. studies [67]. study [1249, 1208, 543, 59]. Studying [39]. Sturmian [656, 451, 1150, 1260, 1264, 1373, 1349, 273, 1262, 1156, 657, 743]. sturmienne [1156]. sturmiennes [451]. sub [1037]. sub-shifts [1037]. subclasses [974]. subcubes [661]. subexponential [1379]. subgraph [389, 1092, 438, 1082]. subgraphs [1056, 554]. Subject [1204, 211, 210]. sublinear [1272, 974]. submodular [608]. suboptimal [54]. subquadratic [1213]. subsequence [437]. subsequences [448]. subsequential [1224]. subset [1122, 1361, 488]. subsets [1285, 1287, 699, 967]. substitution [1190]. Substitutions [13, 1263, 485, 426, 593, 13]. subsystems [166]. subtraction [176]. subtype [1192]. subtyping [761, 625]. subwords [976]. Succinctness [669, 424]. suffice [1327]. suffices [1323]. suffix [852, 409, 1040]. suite [1156]. suites [451, 1208, 432, 585]. summary [886, 964]. summation [548]. summing [966]. sums [112, 1229, 419]. super [501]. Superlinear [176, 889]. Superposition [1141]. superpositions [1279]. supersequences [448]. superstrings [375, 95, 858]. supervisory [174]. surfaces [404, 782]. Switchbox [1053]. switching [731]. Symbolic [1332, 30, 618, 170]. symbols [1175]. Symmetric [499, 1354, 322]. Synchronisation [848, 848]. Synchronization [1212, 1116, 1094, 1065, 459, 498, 198]. synchronized [478]. synchronizers [811]. synchronizing [332]. synchronous [1340, 1311, 950]. syndéticite [1072]. Syntactic [1411, 745, 965]. syntactical [700]. syntax [228, 1320, 1013]. synthesis [785]. System [1060, 899, 825, 1231, 314, 1374, 985, 913, 322, 664, 1352, 1237, 629, 603, 1108]. Systèmes [1072, 13, 331]. systems [728, 18, 1155, 19, 357, 1230, 785, 520, 236, 169, 1332, 500, 20, 1171, 704, 826, 146, 1313, 903, 1153, 1227, 458, 13, 331, 833, 888, 569, 444, 21, 890, 104, 1026, 574, 240, 893, 914, 573, 486, 917, 23, 46, 316, 784, 25, 1102, 1319, 39, 141, 308, 864, 991, 1113, 1398, 602, 1215, 263, 48, 49, 1307, 536, 161, 824, 1076, 1350, 231, 1139, 600, 798, 121, 204, 87, 866, 426, 261, 783, 930, 252, 796, 1159, 239, 44, 529, 26, 1334, 229, 1072, 365, 366].

T [1108]. table [342]. tables [1134]. tabular [1017]. tape [1369]. task [893, 724]. task-trees [724]. tasks [133, 1217, 653]. taxonomical [982]. Team [1153]. teams [7]. technique [487, 1200]. techniques [86, 1135, 1387]. Temporal [1361, 670, 368, 1045, 577, 33, 474, 369, 533, 563, 24, 189, 1319, 48, 750, 940, 884, 714]. tension [895]. term [356, 903, 35, 833, 911, 240, 1137, 235, 923, 231, 1138, 1139, 204, 929, 44]. term-rewriting [833]. Terminal [1049]. terminate [181]. Termination [645, 1201, 44, 84, 444, 602, 204, 87]. terms [773, 504, 282, 700, 89, 621, 1136, 1001, 11, 1391]. test [798, 1353]. testable [660, 382, 259]. Testing [972, 690, 973, 155, 1033, 589, 813, 866]. their [319, 546, 703, 354, 220, 1257, 512, 256, 1267, 1178, 864, 161, 461, 515, 743, 1320]. Theorem [1103, 152, 177, 1268, 979, 1224, 1264, 1032, 643, 660, 300, 141, 782, 992, 1141, 1368, 1293]. theorems [1080, 61, 45, 164, 1131]. Theoret [1425, 735, 1428, 1424, 1112, 761, 1426]. theoretic [1167, 1338, 992, 415]. Theoretical [242, 521, 961, 721, 777]. theories [441, 1341, 540, 521, 501, 1322]. Theory [1415, 1149, 670, 150, 666, 1191, 1044, 1425, 268, 619, 405, 333, 523, 627, 221, 1291, 543, 1046, 1232, 1351, 1405, 47, 1395, 410, 528, 256, 539, 1113, 391, 6, 190, 526, 360, 929, 1252, 252, 1142, 780, 637, 1148]. There [421]. thin [421]. third [793]. third-order [793]. threading [1176]. Three [1382, 446, 1264, 409, 96, 1091, 974, 585]. three-dimensional [1091]. Three-valued [446]. three-way [96]. three-word [409]. Threshold [1147, 681, 298, 583, 582, 8, 1390]. throw [600]. Thue [87, 866]. Thurston [353]. tidy [322]. Tight [133]. tile [816]. Tiling [1280, 355, 662, 353]. Tilings [1326, 165, 295, 1281, 352]. Time [1231, 291, 482, 368, 1117, 1044, 1425, 828, 1160, 1272, 1087, 706, 262, 157, 1213, 1337, 439, 136, 160, 95, 474, 132, 454, 1309, 974, 720, 1363, 338, 391, 1034, 190, 1385, 244, 515, 814, 927, 73, 306, 1121, 714, 971]. time-adaptive [1309]. Time-based [1231]. time-gaps [368]. time-space [157]. timed [238, 1186, 1046, 29, 1182, 27]. timers [830]. times [858]. Timing [1308]. token [652, 1318]. tolerance [775]. tolerant [1151, 1152, 960]. tool [907]. toolbox [881]. tools [39]. top [1089, 889, 297, 105]. top-down [1089, 889, 297, 105]. Topological [470, 217, 167, 587, 1330, 1299, 1300, 163, 219]. topologie [301]. Topology [214, 876, 301, 404, 513, 360, 224, 1228]. Tortoise [1166]. torus [57]. total [1356, 85]. trace [186, 1021, 529]. traces [752, 258, 559]. Tractability [518, 62]. trade [1323]. trade-offs [1323]. trajectories [965]. transactions [197, 845, 714]. Transcendence [1269]. transcendental [321]. transducers [1089, 852, 105]. transductions [908]. transfer [854]. transform [1423]. transformation [758, 1238, 919, 930, 1233]. Transformations [465, 719, 239, 889, 297, 269, 835, 943, 479]. transformer [148]. transformers [872, 604, 208]. Transforming [1101]. transforms [112, 115]. transient [1087]. transition [146, 39, 141, 991, 1334]. transitions [676]. translation [1320]. translational [51]. Translations [46, 1012, 884]. transmission [1343]. Transporting [572]. transpositions [1179]. Trapping [247]. treatment [222]. Tree [709, 455, 1177, 1089, 765, 457, 852, 889, 718, 1052, 908, 706, 1040, 297, 97, 834, 720, 300, 387, 61, 1174, 478, 105, 1079, 259, 971]. tree-like [1174]. tree-packing [720]. tree-to-graph-to-tree [908]. Trees [668, 1225, 177, 156, 1275, 1325, 656, 1054, 101, 309, 292, 1327, 1328, 1086, 507, 1083, 321, 438, 1318, 265, 117, 97, 695, 296, 1329, 119, 1214, 1278, 810, 778, 713, 1321, 121, 812, 1130, 1228, 63, 724, 1023]. treewidth [1143, 1062, 612]. triangles [432]. Triangular [958, 404]. tries [1222]. triple [612]. triple-free [612]. trois [585]. trolley [857]. trous [353]. true [260, 1105]. true-concurrency [1105]. Truly [1061, 582]. truth [342]. truth-table [342]. Turakainen [300]. Turing [172, 96, 1173, 492, 587, 1427, 491, 501, 1242, 685, 967]. tutorial [1044, 1425]. TWDM [891]. twice [557]. Two [648, 282, 262, 985, 428, 750, 1282, 1222, 270, 825, 700, 484, 402, 512, 67, 1359, 96, 1173, 495, 732, 1041, 558, 967, 611]. two-component [825]. two-counter [495]. Two-dimensional [428, 1282, 270, 96, 1041, 558, 611]. two-person [1222]. Two-sorted [750]. two-way [512]. type [518, 520, 633, 1380, 1335, 523, 627, 1197, 37, 573, 860, 229]. typed [227, 904, 89, 4, 528, 464]. types [1193, 282, 192, 577, 910, 774, 892, 925, 198, 802, 1233]. typing [899, 878].

UCT [402]. UET [402]. Ulam [732]. ultimate [1208]. ultimement [1208]. ultrametric [880, 526]. Unambiguous [892]. Unary [1109, 758, 1238, 688, 1239]. unavoidable [1036]. Unbordered [1098]. uncertainty [532, 56]. uncrossing [322]. Undecidability [188, 471, 689, 1115, 593]. undecidable [1066, 619, 588, 87, 1142]. underdiagonal [1277]. Unfolding [80, 249]. unicité [353]. Unification [1099, 911, 621, 471, 40, 1344, 508, 1000, 1140, 1368]. unified [304]. Uniform [1276, 958, 1055, 127, 1308, 1278, 1296, 1105, 1339, 1245]. uniforme [349]. uniformly [349, 266]. Unifying [628, 845, 755]. Union [262, 1057]. Union-Find [1057]. unions [765]. Uniqueness [283, 899, 353, 125, 1241]. unit [579, 653, 614, 1301]. unitary [653]. units [1242]. UNITY [34]. Univariate [846]. Universal [489, 20, 880, 405, 499, 719, 502, 1173, 494, 493, 491, 1035]. Universality [547, 495]. universelles [1035]. unknowns [314]. unnecessary [80]. unpredictable [835]. unreliable [1160]. unrestricted [689]. unsatisfied [1155]. update [968]. updates [982]. updating [1040, 1033]. upgrading [1329]. upon [939]. upward [1054]. Urquhart [1006]. use [703, 907, 322, 1303]. useful [276]. Using [1306, 1095, 901, 522, 130, 808, 957, 653, 1307, 173, 765, 649, 1276, 328, 89, 569, 1319, 1345, 1178, 48, 922, 1027, 1016].

V [1113]. Valentin [286, 285]. Valiant [109]. valid [25]. validation [825]. Valuations [12, 120]. value [1108, 1338, 1410, 11]. valued [1284, 392, 535, 1030, 446, 8]. values [1260, 305, 1209, 193]. Vapnik [9]. variable [1249, 435, 179, 654, 476, 340]. variable-constraint [340]. variable-length [435]. variables [566, 1155, 113, 1354, 1359, 553, 1295, 80]. variance [120]. Variations [1336]. Variétés [131]. varieties [406, 696, 1079, 131]. Various [592, 135, 1202, 1426]. VC [8]. VC-dimension [8]. vector [1118, 664, 412, 558]. Verification [1319, 647, 516, 48, 1235]. verify [522]. Verifying [588]. version [797, 109]. versus [255, 431, 519, 249, 140]. vertex [1372]. vertical [354, 288]. via [898, 1194, 876, 1063, 1291, 778, 534, 109, 365, 366]. Viability [23]. view [755, 356, 870, 31, 1415, 1294, 841, 1200]. viewing [948]. virtual [1125]. visibility [53]. Visualization [959]. VLSI [972, 1053, 973]. Volume [374, 530, 781]. Volumes [213, 211, 210, 701, 1206, 1204]. Voronoi [586, 51]. votes [1386]. voting [1386]. vs [859, 1021, 598].

W [850]. w.r.t [1377]. Wait [1313]. Wait-free [1313]. Walk [804]. walks [1277, 387]. warehouse [52]. watermarking [1387]. wavelength [1328]. way [319, 512, 96, 1352, 73, 306]. Weak [98, 148, 144, 1009, 261, 1154]. weakly [550, 1354]. weight [722]. weighted [1278]. weights [403]. well [167, 469, 1073, 147]. well-founded [469, 147]. well-nested [167]. wellorderings [523, 627]. where [1222]. which [409, 1377, 153, 181]. ‘While [1300]. whose [132]. Wide [729]. Wide-sense [729]. width [1114, 15]. Wilf [1268, 1264]. winning [581]. wireless [126]. with/ without [296]. within [1008]. without [130, 353, 64, 28, 855, 721, 1039, 422]. word [744, 409, 1364, 1070, 1374, 153, 1000, 1036, 422]. word-extension [1036]. words [656, 1268, 1264, 1373, 349, 1349, 545, 1265, 1428, 683, 1025, 273, 1267, 1270, 329, 578, 1262, 381, 555, 437, 976, 743, 1261]. work [439]. work-time [439]. Workshop [352, 489, 791, 886, 214, 310, 595]. workstations [960]. Worst [1323, 427, 1348]. Worst-case [1323, 427, 1348]. wp [148]. write [946]. writer [201]. writings [1208]. Wythoff [1093].

XI [1357]. XRAM [897].

Yekutieli [713]. yield [298]. Yoneda

[876, 1412].

zero [856]. zero-one [856]. ZFC [666, 1191]. Ziv [117]. zooming [157].

References

Matos:1994:PSI

[1] Armando B. Matos. Periodic sets of integers. Theoretical Computer Science, 127(2):287–312, May 23, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                              tronic). URL http://www.elsevier.
                                              com/cgi-bin/cas/tree/store/tcs/
                                              cas_sub/browse/browse.cgi?year=
                                              1994&volume=127&issue=2&aid=1479.

See remark [72].

Botelho:1994:BNN

[2] Fernando Botelho and Max Garzon. Boolean neural nets are observable. Theoretical Computer Science, 134(1):51–61, November 07, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                              tronic). URL http://www.elsevier.
                                              com/cgi-bin/cas/tree/store/tcs/
                                              cas_sub/browse/browse.cgi?year=
                                              1994&volume=134&issue=1&aid=1731.

See erratum [735].

Galmiche:1994:PNL

[3] Didier Galmiche and Guy Perrier. On proof normalization in linear logic. Theoretical Computer Science, 135(1):67–110, December 05, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                              tronic). URL http://www.elsevier.
                                              com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1994&volume=135&issue=1&aid=1746.

See corrigendum [1112].

Hagiya:1995:TCP

[4] Masami Hagiya. A typed λ-calculus for proving-by-example and bottom-up generalization procedure. Theoretical Computer Science, 137(1):3–23, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1786.

Jantke:1995:CBR

[5] Klaus P. Jantke and Steffen Lange. Case-based representation and learning of pattern languages. Theoretical Computer Science, 137(1):25–51, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1787.

Mukouchi:1995:TMT

[6] Yasuhito Mukouchi and Setsuo Arikawa. Towards a mathematical theory of machine discovery from facts. Theoretical Computer Science, 137(1):53–84, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1788.

Jain:1995:ATL

[7] Sanjay Jain and Arun Sharma. On aggregating teams of learning machines. Theoretical Computer Science, 137(1):85–108, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1789.

Sakurai:1995:VDD

[8] Akito Sakurai. On the VC-dimension of depth four threshold circuits and the complexity of Boolean-valued functions. Theoretical Computer Science, 137(1):109–127, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1790.

Shinohara:1995:CCV

[9] Ayumi Shinohara. Complexity of computing Vapnik–Chervonenkis dimension and some generalized dimensions. Theoretical Computer Science, 137(1):129–144, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1791.

Hasegawa:1995:ALS

[10] Susumu Hasegawa, Hiroshi Imai, and Masaki Ishiguro. ε-approximations of k-label spaces. Theoretical Computer Science, 137(1):145–157, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1792.

Nakamura:1995:ELL

[11] Atsuyoshi Nakamura and Naoki Abe. Exact learning of linear combinations of monotone terms from function value queries. Theoretical Computer Science, 137(1):159–176, January 09, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=1&aid=1793.

Fernau:1995:VLA

[12] Henning Fernau. Valuations of languages, with applications to fractal geometry. Theoretical Computer Science, 137(2):177–217, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=2&aid=1690.

Fabre:1995:SSN

[13] Stéphane Fabre. Substitutions et systèmes de numération. (French) [Substitutions and beta numbering systems]. Theoretical Computer Science, 137(2):219–236, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=2&aid=1692.

Esik:1995:EPK

[14] Z. Ésik and L. Bernátsky. Equational properties of Kleene algebras of relations with conversion. Theoretical Computer Science, 137(2):237– 251, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=137&issue=
2&aid=1682.

Takahashi:1995:MSP

[15] Atsushi Takahashi, Shuichi Ueno, and Yoji Kajitani. Mixed searching and proper-path-width. Theoretical Computer Science, 137(2):253–268, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=2&aid=1813.

Breslauer:1995:FPS

[16] Dany Breslauer. Fast parallel string prefix-matching. Theoretical Computer Science, 137(2):269–278, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=2&aid=1811.

Arvind:1995:INP

[17] Vikraman Arvind, Johannes Köbler, Uwe Schöning, and Rainer Schuler. If NP has polynomial-size circuits, then MA=AM. Theoretical Computer Science, 137(2):279–282, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1995&volume=137&issue=2&aid=1783.

Alur:1995:AAH

[18] R. Alur, C. Courcoubetis, N. Halbwachs, T. A. Henzinger, P.-H. Ho, X. Nicollin, A. Olivero, J. Sifakis, and S. Yovine. The algorithmic analysis of hybrid systems. Theoretical Computer Science, 138(1):3–34, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=1&aid=1836.

Asarin:1995:RAD

[19] Eugene Asarin, Oded Maler, and Amir Pnueli. Reachability analysis of dynamical systems having piecewiseconstant derivatives. Theoretical Computer Science, 138(1):35–65, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=1&aid=1862.

Branicky:1995:UCO

[20] Michael S. Branicky. Universal computation and other capabilities of hybrid and continuous dynamical systems. Theoretical Computer Science, 138(1):67–100, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=1&aid=1797.

Grossman:1995:AAH

[21] R. L. Grossman and R. G. Larson. An algebraic approach to hybrid systems. Theoretical Computer Science, 138(1):101–112, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=138&issue=1&aid=1796.

Hansen:1995:FD

[22] Michael R. Hansen, Paritosh K. Pandya, and Zhou Chaochen. Finite divergence. Theoretical Computer Science, 138(1):113–139, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=138&issue=1&aid=1795.

Kohn:1995:VHS

[23] Wolf Kohn, Anil Nerode, Jeffrey B. Remmel, and Alexander Yakhnis. Viability in hybrid systems. Theoretical Computer Science, 138(1):141–168, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=138&issue=1&aid=1800.

Lakhneche:1995:MTL

[24] Yassine Lakhneche and Jozef Hooman. Metric temporal logic with durations. Theoretical Computer Science, 138(1):169–199, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=1&aid=1801.

Lemmon:1995:IIV

[25] Michael Lemmon and Panos J. Antsaklis. Inductively inferring valid logical models of continuous-state dynamical systems. Theoretical Computer Science, 138(1):201–210, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=1&aid=1798.

Zhang:1995:CNS

[26] Ying Zhang and Alan K. Mackworth. Constraint nets: a semantic model for hybrid dynamic systems. Theoretical Computer Science, 138(1):211–239, February 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=1&aid=1799.

Davies:1995:BHT

[27] Jim Davies and Steve Schneider. A brief history of Timed CSP. Theoretical Computer Science, 138(2):243– 271, February 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=138&issue=
2&aid=1803.

Mislove:1995:FPC

[28] M. W. Mislove, A. W. Roscoe, and S. A. Schneider. Fixed points without completeness. Theoretical Computer Science, 138(2):273–314, February 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=138&issue=2&aid=1804.

Lowe:1995:PPM

[29] Gavin Lowe. Probabilistic and prioritized models of timed CSP. Theoretical Computer Science, 138(2):315– 352, February 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1995&volume=138&issue=
                                        2&aid=1805.

Hennessy:1995:SB

[30] M. Hennessy and H. Lin. Symbolic bisimulations. Theoretical Computer Science, 138(2):353–389, February 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=138&issue=2&aid=1806.

DeNicola:1995:PAV

[31] Rocco De Nicola and Roberto Segala. A process algebraic view of input/ output automata. Theoretical Computer Science, 138(2):391–423, February 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=2&aid=1807.

Pinna:1995:NEA

[32] G. Michele Pinna and Axel Poigné. On the nature of events: another perspective in concurrency. Theoretical Computer Science, 138(2):425–454, February 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=138&issue=2&aid=1808.

Hodkinson:1995:GTF

[33] Ian Hodkinson. On Gabbay’s temporal fixed point operator. Theoretical Computer Science, 139(1–2):1–25, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=139&issue=1-2&aid=1725.

Pappinghaus:1995:LU

[34] Peter Päppinghaus. On the logic of UNITY. Theoretical Computer Science, 139(1–2):27–67, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=139&issue=1-2&aid=1684.

Cockett:1995:SCD

[35] J. Robin B. Cockett and Dwight Spencer. Strong categorical datatypes. II. A term logic for categorical programming. Theoretical Computer Science, 139(1–2):69–113, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=139&issue=1-2&aid=1740.

Barr:1995:NAC

[36] Michael Barr. Nonsymmetric ∗- autonomous categories. Theoretical Computer Science, 139(1–2):115– 130, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1995&volume=139&issue=
                                        1-2&aid=1730.

Ghelli:1995:DTC

[37] Giorgio Ghelli. Divergence of F_{<} type checking. Theoretical Computer Science, 139(1–2):131–162, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=139&issue=1-2&aid=1678.

Bergeron:1995:SCD

[38] Anne Bergeron. Sharing out control in distributed processes. Theoretical Computer Science, 139(1–2): 163–186, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1995&volume=139&issue=
                                        1-2&aid=1773.

Malacaria:1995:SET

[39] Pasquale Malacaria. Studying equivalences of transition systems with algebraic tools. Theoretical Computer Science, 139(1–2):187–205, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=139&issue=1-2&aid=1688.

Dougherty:1995:CLA

[40] Daniel J. Dougherty and Patricia Johann. A combinatory logic approach to higher-order E-unification. Theoretical Computer Science, 139(1–2): 207–242, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=139&issue=
1-2&aid=1844.

Moschovakis:1995:CCP

[41] Yiannis N. Moschovakis. Computable concurrent processes. Theoretical Computer Science, 139(1–2): 243–273, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=139&issue=
1-2&aid=1758.

Bernot:1995:OSI

[42] Gilles Bernot, Michel Bidoit, and Teodor Knapik. Observational specifications and the indistinguishability assumption. Theoretical Computer Science, 139(1–2):275–314, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=139&issue=1-2&aid=1658.

Inverardi:1995:DOC

[43] P. Inverardi and M. Nesi. Deciding observational congruence of finitestate CCS expressions by rewriting. Theoretical Computer Science, 139(1–2):315–354, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=139&issue=1-2&aid=1715.

Weiermann:1995:TPT

[44] Andreas Weiermann. Termination proofs for term rewriting systems by lexicographic path orderings imply multiply recursive derivation lengths. Theoretical Computer Science, 139(1–2):355–362, March 06, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=139&issue=1-2&aid=1776.

Pigozzi:1995:LAA

[45] Don Pigozzi and Antonino Salibra. Lambda abstraction algebras: representation theorems. Theoretical Computer Science, 140(1):5–52, March 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1995&volume=140&issue=1&aid=1837.

Laroussinie:1995:TBM

[46] F. Laroussinie, S. Pinchinat, and Ph. Schnoebelen. Translations between modal logics of reactive systems. Theoretical Computer Science, 140(1):53–71, March 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=140&issue=1&aid=1838.

Gorrieri:1995:TPD

[47] Roberto Gorrieri, Marco Roccetti, and Enrico Stancampiano. A theory of processes with durational Actions. Theoretical Computer Science, 140(1):73–94, March 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=140&issue=1&aid=1839.

Mokkedem:1995:UTL

[48] Abdelillah Mokkedem and Dominique Méry. On using temporal logic for refinement and compositional verification of concurrent systems. Theoretical Computer Science, 140(1): 95–138, March 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=140&issue=
1&aid=1840.

Navarro:1995:CMS

[49] Marisa Navarro, Fernando Orejas, and Ana Sánchez. On the correctness of modular systems. Theoretical Computer Science, 140(1):139–177, March 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=140&issue=1&aid=1841.

Wagner:1995:RMO

[50] E. G. Wagner. On the role of memory in object-based and objectoriented languages. Theoretical Computer Science, 140(1):179–199, March 20, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=140&issue=1&aid=1842.

Dattasharma:1995:AVR

[51] Abhi Dattasharma and S. Sathiya Keerthi. An augmented Voronoi roadmap for 3D translational motion planning for a convex polyhedron moving amidst convex polyhedral obstacles. Theoretical Computer Science, 140(2):205–230, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=140&issue=2&aid=1825.

Sarrafzadeh:1995:DWP

[52] Majid Sarrafzadeh and Sanjeev R. Maddila. Discrete warehouse problem. Theoretical Computer Science, 140(2):231–247, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=140&issue=2&aid=1826.

Prasad:1995:NCS

[53] L. Prasad and S. S. Iyengar. A note on the combinatorial structure of the visibility graph in simple polygons. Theoretical Computer Science, 140(2):249–263, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=140&issue=2&aid=1827.

Rao:1995:FPS

[54] Nageswara S. V. Rao. On fast planning of suboptimal paths amidst polygonal obstacles in plane. Theoretical Computer Science, 140(2):265–289, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=140&issue=2&aid=1828.

Sridhar:1995:EAS

[55] R. Sridhar, K. Han, and N. Chandrasekharan. Efficient algorithms for shortest distance queries on special classes of polygons. Theoretical Computer Science, 140(2):291–300, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=140&issue=2&aid=1829.

deBerg:1995:RGD

[56] Mark de Berg, Leonidas Guibas, Dan Halperin, Mark Overmars, Otfried Schwarzkopf, Micha Sharir, and Monique Teillaud. Reaching a goal with directional uncertainty. Theoretical Computer Science, 140(2):301–317, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=140&issue=2&aid=1871.

Seetharaman:1995:SDS

[57] Guna Seetharaman. A simplified design strategy for mapping image processing algorithms on a SIMD torus. Theoretical Computer Science, 140(2):319–331, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=140&issue=2&aid=1830.

Bettayeb:1995:AH

[58] Saïd Bettayeb. On the k-ary hypercube. Theoretical Computer Science, 140(2):333–339, April 03, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=140&issue=2&aid=1831.

Gruau:1995:FSN

[59] Frédéric Gruau, Jean-Yves Ratajszczak, and Gilles Wiber. Fundamental study: a neural compiler. Theoretical Computer Science, 141(1– 2):1–52, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=141&issue=
1-2&aid=1834.

Johansen:1995:LSM

[60] Peter Johansen. On-line string matching with feedback. Theoretical Computer Science, 141(1–2):53–67, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1852.

Muller:1995:SAT

[61] David E. Muller and Paul E. Schupp. Simulating alternating tree automata by nondeterministic automata: New results and new proofs of the theorems of Rabin, McNaughton and Safra. Theoretical Computer Science, 141(1–2):69–107, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1848.

Downey:1995:FPT

[62] Rod G. Downey and Michael R. Fellows. Fixed-parameter tractability and completeness II: On completeness for W [1]. Theoretical Computer Science, 141(1–2):109–131, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1738.

Yao:1995:ADT

[63] Andrew Chi-Chih Yao. Algebraic decision trees and Euler characteristics. Theoretical Computer Science, 141(1–2):133–150, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1723.

Kapur:1995:LLO

[64] Shyam Kapur and Gianfranco Bilardi. Language learning without overgeneralization. Theoretical Computer Science, 141(1–2):151–162, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1716.

Apostolico:1995:PDA

[65] Alberto Apostolico, Dany Breslauer, and Zvi Galil. Parallel detection of all palindromes in a string. Theoretical Computer Science, 141(1–2):163–173, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1724.

Jenner:1995:CFP

[66] Birgit Jenner and Jacobo Torán. Computing functions with parallel queries to NP. Theoretical Computer Science, 141(1–2):175–193, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1721.

Gorrieri:1995:ICC

[67] Roberto Gorrieri and Ugo Montanari. On the implementation of concurrent calculi in net calculi: two case studies. Theoretical Computer Science, 141(1–2):195–252, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1714.

Kari:1995:MPL

[68] Lila Kari, Alexandru Mateescu, Gheorghe Păun, and Arto Salomaa. Multipattern languages. Theoretical Computer Science, 141(1–2):253–268, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1728.

Haastad:1995:SER

[69] Johan Håstad, Alexander Razborov, and Andrew Yao. On the shrinkage exponent for read-once formulae. Theoretical Computer Science, 141(1–2):269–282, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1722.

Sieling:1995:GDB

[70] Detlef Sieling and Ingo Wegener. Graph driven BDDs — a new data structure for Boolean functions. Theoretical Computer Science, 141(1–2):283–310, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1719.

Mizoguchi:1995:RGR

[71] Yoshihiro Mizoguchi and Yasuo Kawahara. Relational graph rewritings. Theoretical Computer Science, 141(1–2):311–328, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1717.

Petersen:1995:RPB

[72] H. Petersen. A remark on a paper by A. B. Matos. Theoretical Computer Science, 141(1–2):329–330, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=

1995&volume=141&issue=1-2&aid=1851.

See [1].

Terrier:1995:RTO

[73] Véronique Terrier. On real time oneway cellular array. Theoretical Computer Science, 141(1–2):331–335, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1846.

Dessmark:1995:MLL

[74] Anders Dessmark, Andrzej Lingas, and Anil Maheshwari. Multi-list layering: Complexity and applications. Theoretical Computer Science, 141(1–2):337–350, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=141&issue=1-2&aid=1847.

Justin:1995:QAF

[75] J. Justin and G. Pirillo. On a question about factorization forests. Theoretical Computer Science, 141(1–2): 351–355, April 17, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=141&issue=
1-2&aid=1843.

Maluszynski:1995:P

[76] J. Maluszynski and M. Wirsing. Preface. Theoretical Computer Science, 142(1):1–??, 1995. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aasa:1995:PSI

[77] Annika Aasa. Precedences in specifications and implementations of programming languages. Theoretical Computer Science, 142(1):3–26, May 01, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=1&aid=1860.

Alpuente:1995:ICS

[78] María Alpuente, Moreno Falaschi, and Giorgio Levi. Incremental constraint satisfaction for equational logic programming. Theoretical Computer Science, 142(1):27–57, May 01, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=1&aid=1858.

Loogen:1995:DDD

[79] Rita Loogen and Stephan Winkler. Dynamic detection of determinism in functional logic languages. Theoretical Computer Science, 142(1):59–87, May 01, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=1&aid=1859.

Proietti:1995:UDF

[80] Maurizio Proietti and Alberto Pettorossi. Unfolding–definition–folding, in this order, for avoiding unnecessary variables in logic programs. Theoretical Computer Science, 142(1):89–124, May 01, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=1&aid=1861.

Nilsson:1995:AIK

[81] Ulf Nilsson. Abstract interpretation: a kind of magic. Theoretical Computer Science, 142(1):125–138, May 01, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=1&aid=1857.

Kirchner:1995:E

[82] C. Kirchner. Editorial. Theoretical Computer Science, 142(2):139–??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lynch:1995:RCC

[83] Christopher Lynch and Wayne Snyder. Redundancy criteria for constrained completion. Theoretical Computer Science, 142(2):141–177, May 15, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=2&aid=1908.

Dershowitz:1995:NT

[84] Nachum Dershowitz and Charles Hoot. Natural termination. Theoretical Computer Science, 142(2):179–207, May 15, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=2&aid=1909.

Rubio:1995:TAC

[85] Albert Rubio and Robert Nieuwenhuis. A total AC-compatible ordering based on RPO. Theoretical Computer Science, 142(2):209–227, May 15, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=2&aid=1910.

Baader:1995:CTD

[86] Franz Baader and Klaus U. Schulz. Combination techniques and decision problems for disunification. Theoretical Computer Science, 142(2):229–255, May 15, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=2&aid=1911.

Senizergues:1995:SUT

[87] Géraud Sénizergues. Some undecidable termination problems for semi-Thue systems. Theoretical Computer Science, 142(2):257–276, May 15, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=2&aid=1912.

Asperti:1995:PCL

[88] Andrea Asperti and Cosimo Laneve. Paths, computations and labels in the lambda-calculus. Theoretical Computer Science, 142(2):277–297, May 15, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=2&aid=1913.

Gallier:1995:PPT

[89] Jean Gallier. Proving properties of typed λ-terms using realizability and covers, and sheaves. Theoretical Computer Science, 142(2):299–368, May 15, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=142&issue=2&aid=1914.

Van:1995:CRL

[90] Do Long Van, B. Le Saec, and I. Litovsky. Characterizations of rational ω-languages by means of right congruences. Theoretical Computer Science, 143(1):1–21, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1741.

Krithivasan:1995:NPA

[91] Kamala Krithivasan and Meena Mahajan. Nondeterministic, probabilistic and alternating computations on cellular array models. Theoretical Computer Science, 143(1):23–49, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1737.

Antimirov:1995:RER

[92] Valentin M. Antimirov and Peter D. Mosses. Rewriting extended regular expressions. Theoretical Computer Science, 143(1):51–72, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1726.

Bender:1995:PIO

[93] Michael A. Bender, Michel Gastaldo, and Michel Morvan. Parallel interval order recognition and construction of interval representations. Theoretical Computer Science, 143(1):73–91, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1739.

Goldstein:1995:CPG

[94] Arthur S. Goldstein and Edward M. Reingold. The complexity of pursuit on a graph. Theoretical Computer Science, 143(1):93–112, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1872.

Jiang:1995:SCS

[95] Tao Jiang and Vadim G. Timkovsky. Shortest consistent superstrings computable in polynomial time. Theoretical Computer Science, 143(1):113–122, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1886.

Ito:1995:OST

[96] Akira Ito, Katsushi Inoue, Itsuo Takanami, and Yue Wang. Optimal simulation of two-dimensional alternating finite automata by threeway nondeterministic Turing machines. Theoretical Computer Science, 143(1): 123–135, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=143&issue=
1&aid=1869.

Jiang:1995:ATA

[97] Tao Jiang, Lusheng Wang, and Kaizhong Zhang. Alignment of trees

— an alternative to tree edit. Theoretical Computer Science, 143(1):137–148, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1867.

Juedes:1995:WC

[98] David W. Juedes and Jack H. Lutz. Weak completeness in E and E_{2}. Theoretical Computer Science, 143(1):149–158, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1866.

Crescenzi:1995:RSS

[99] Pierluigi Crescenzi and Christos H. Papadimitriou. Reversible simulation of space-bounded computations. Theoretical Computer Science, 143(1):159–165, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1875.

Miltersen:1995:CPC

[100] Peter Bro Miltersen. On the cell probe complexity of polynomial evaluation. Theoretical Computer Science, 143(1):167–174, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1870.

DePrisco:1995:MPL

[101] Roberto De Prisco, Giuseppe Parlati, and Giuseppe Persiano. Minimal path length of trees with known fringe. Theoretical Computer Science, 143(1):175–188, May 29, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=1&aid=1877.

Malgouyres:1995:GGC

[102] Rémy Malgouyres. Graphs generalizing closed curves with linear construction of the Hamiltonian cycle. Parametrization of discretized curves. Theoretical Computer Science, 143(2):189–249, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=2&aid=1775.

Matamala:1995:RCP

[103] Martín Matamala. Recursive construction of periodic steady state for neural networks. Theoretical Computer Science, 143(2):251–267, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=2&aid=1755.

Harrison:1995:DPP

[104] John Harrison. Dynamical properties of PWD0L systems. Theoretical Computer Science, 143(2):269–284, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=2&aid=1750.

Slutzki:1995:DTT

[105] Giora Slutzki and Sándor Vágvölgyi. Deterministic top-down tree transducers with iterated look-ahead. Theoretical Computer Science, 143(2):285–308, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=143&issue=2&aid=1752.

Chen:1995:MDS

[106] Zhi-Zhong Chen. The maximal fdependent set problem for planar graphs is in NC. Theoretical Computer Science, 143(2):309–318, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=143&issue=2&aid=1753.

Fraenkel:1995:MNM

[107] Aviezri S. Fraenkel, Alan Jaffray, Anton Kotzig, and Gert Sabidussi. Modular Nim (mathematical game). Theoretical Computer Science, 143(2):319–333, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=143&issue=2&aid=1894.

Gubas:1995:NLB

[108] Xaver Gubáš, Juraj Hromkovič, and Juraj Waczulík. A nonlinear lower bound on the practical combinational complexity. Theoretical Computer Science, 143(2):335–342, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=143&issue=2&aid=1903.

Rytter:1995:CFR

[109] Wojciech Rytter. Context-free recognition via shortest paths computation: a version of Valiant’s algorithm. Theoretical Computer Science, 143(2):343–352, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=2&aid=1899.

Zhang:1995:ALC

[110] Louxin Zhang. On the approximation of longest common nonsupersequences and shortest common nonsubsequences. Theoretical Computer Science, 143(2):353–362, June 12, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=143&issue=2&aid=1891.

Prodinger:1995:P

[111] H. Prodinger and W. Szpankowski. Preface. Theoretical Computer Science, 144(1–2):1–??, 1995. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flajolet:1995:MTAa

[112] Philippe Flajolet, Xavier Gourdon, and Philippe Dumas. Mellin transforms and asymptotics: Harmonic sums. Theoretical Computer Science, 144(1–2):3–58, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=144&issue=1-2&aid=1958.

Bergeron:1995:CDF

[113] François Bergeron and Ulrike Sattler. Constructible differentially finite algebraic series in several variables. Theoretical Computer Science, 144(1–2):59–65, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=144&issue=1-2&aid=1931.

Drmota:1995:MCC

[114] Michael Drmota and Michèle Soria. Marking in combinatorial constructions: Generating functions and limiting distributions. Theoretical Computer Science, 144(1–2):67–99, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=144&issue=1-2&aid=1928.

Flajolet:1995:MTAb

[115] Philippe Flajolet and Robert Sedgewick. Mellin transforms and asymptotics: Finite differences and Rice’s integrals. Theoretical Computer Science, 144(1–2):101–124, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=144&issue=1-2&aid=1915.

Gardy:1995:DAS

[116] Danièle Gardy and Guy Louchard. Dynamic analysis of some relational databases parameters. Theoretical Computer Science, 144(1–2):125–159, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=144&issue=1-2&aid=1939.

Jacquet:1995:ABL

[117] Philippe Jacquet and Wojciech Szpankowski. Asymptotic behavior of the Lempel–Ziv parsing scheme and digital search trees. Theoretical Computer Science, 144(1–2):161–197, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=144&issue=1-2&aid=1932.

Kirschenhofer:1995:AOS

[118] Peter Kirschenhofer, Conrado Martínez, and Helmut Prodinger. Analysis of an optimized search algorithm for skip lists. Theoretical Computer Science, 144(1–2):199–220, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=144&issue=1-2&aid=1930.

Mahmoud:1995:PAB

[119] Hosam M. Mahmoud and Robert T. Smythe. Probabilistic analysis of bucket recursive trees. Theoretical Computer Science, 144(1–2):221–249, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=144&issue=1-2&aid=1942.

Schachinger:1995:VCI

[120] Werner Schachinger. On the variance of a class of inductive valuations of data structures for digital search. Theoretical Computer Science, 144(1–2):251–275, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=144&issue=1-2&aid=1940.

Schmid:1995:RTQ

[121] U. Schmid. Random trees in queueing systems with deadlines. Theoretical Computer Science, 144(1–2):277–314, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=144&issue=1-2&aid=1916.

Braga:1995:PGE

[122] G. Braga, G. Cattaneo, P. Flocchini, and C. Quaranta Vogliotti. Pattern growth in elementary cellular automata. Theoretical Computer Science, 145(1–2):1–26, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1782.

Srikrishna:1995:OPA

[123] G. Srikrishna and C. Pandu Rangan. Optimal parallel algorithms for path problems on planar graphs. Theoretical Computer Science, 145(1–2):27–43, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1823.

Breitbart:1995:SBD

[124] Y. Breitbart, H. Hunt, III, and D. Rosenkrantz. On the size of binary decision diagrams representing Boolean functions. Theoretical Computer Science, 145(1–2):45–69, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1815.

Leiss:1995:ILE

[125] Ernst L. Leiss. Implicit language equations: existence and uniqueness of solutions. Theoretical Computer Science, 145(1–2):71–93, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1810.

Diks:1995:AWR

[126] Krzysztof Diks, Evangelos Kranakis, Adam Malinowski, and Andrzej Pelc. Anonymous wireless rings. Theoretical Computer Science, 145(1–2):95–109, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1812.

Creignou:1995:CPL

[127] Nadia Creignou. The class of problems that are linearly equivalent to Satisfiability or a uniform method for proving NP-completeness. Theoretical Computer Science, 145(1–2):111–145, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1816.

Stewart:1995:CPM

[128] Iain A. Stewart. Complete problems for monotone NP. Theoretical Computer Science, 145(1–2):147–157, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1809.

Drewes:1995:GSA

[129] F. Drewes, A. Habel, H.-J. Kreowski, and S. Taubenberger. Generating self-affine fractals by collage grammars. Theoretical Computer Science, 145(1–2):159–187, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1759.

Cai:1995:UMD

[130] Jiazhen Cai and Robert Paige. Using multiset discrimination to solve language processing problems without hashing. Theoretical Computer Science, 145(1–2):189–228, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1817.

Reutenauer:1995:VFR

[131] Christophe Reutenauer and Marcel Paul Schützenberger. Variétés et fonctions rationnelles. (French) [Varieties and rational functions]. Theoretical Computer Science, 145(1–2): 229–240, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=145&issue=
1-2&aid=1814.

Ko:1995:PTC

[132] Ker-I Ko. A polynomial-time computable curve whose interior has a nonrecursive measure. Theoretical Computer Science, 145(1–2):241–270, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1781.

Biran:1995:TBR

[133] Ofer Biran, Shlomo Moran, and Shmuel Zaks. Tight bounds on the round complexity of distributed 1solvable tasks. Theoretical Computer Science, 145(1–2):271–290, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1784.

Ginsburg:1995:IQO

[134] Seymour Ginsburg and Katsumi Tanaka. Interval queries on object histories. Theoretical Computer Science, 145(1–2):291–316, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1774.

Middendorf:1995:FVM

[135] Martin Middendorf. On finding various minimal, maximal, and consistent sequences over a binary alphabet. Theoretical Computer Science, 145(1–2):317–327, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1954.

Jamison:1995:LTR

[136] B. Jamison and S. Olariu. A lineartime recognition algorithm for P_{4}- reducible graphs. Theoretical Computer Science, 145(1–2):329–344, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1956.

Zhang:1995:CRB

[137] Liang Zhang and Zhonghui Shen. Completion of recognizable bifix codes. Theoretical Computer Science, 145(1–2):345–355, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1934.

Benson:1995:SEA

[138] Gary Benson. A space efficient algorithm for finding the best nonoverlapping alignment score. Theoretical Computer Science, 145(1–2):357–369, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1943.

Hemaspaandra:1995:PSI

[139] Lane A. Hemaspaandra and Zhigen Jiang. P-selectivity: Intersections and indices. Theoretical Computer Science, 145(1–2):371–380, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=145&issue=1-2&aid=1924.

Teng:1995:ISV

[140] Shang-Hua Teng. Independent sets versus perfect matchings. Theoretical Computer Science, 145(1–2):381–390, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1923.

Maler:1995:DTP

[141] Oded Maler. A decomposition theorem for probabilistic transition systems. Theoretical Computer Science, 145(1–2):391–396, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=145&issue=1-2&aid=1944.

Scedrov:1995:MA

[142] Andre Scedrov, Dennis DeTurk, and Wolfgang Ziller. Moez Alimohamed, 1967–1994. Theoretical Computer Science, 146(1–2):1–3, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1941.

Alimohamed:1995:CLD

[143] Moez Alimohamed. A characterization of lambda definability in categorical models of implicit polymorphism. Theoretical Computer Science, 146(1–2):5–23, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1917.

Bloom:1995:SOS

[144] Bard Bloom. Structural operational semantics for weak bisimulations. Theoretical Computer Science, 146(1–2):25–68, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=146&issue=1-2&aid=1779.

Ariola:1995:PFO

[145] Zena M. Ariola and Arvind XXX. Properties of a first-order functional language with sharing. Theoretical Computer Science, 146(1–2):69–108, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=146&issue=1-2&aid=1819.

Cassez:1995:CER

[146] Franck Cassez and Olivier Roux. Compilation of the ELECTRE reactive language into finite transition systems. Theoretical Computer Science, 146(1–2):109–143, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=146&issue=1-2&aid=1777.

Kemp:1995:BEQ

[147] David B. Kemp, Divesh Srivastava, and Peter J. Stuckey. Bottom-up evaluation and query optimization of wellfounded models. Theoretical Computer Science, 146(1–2):145–184, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1780.

Berghammer:1995:FDP

[148] Rudolf Berghammer, Birgit Elbl, and Ulf Schmerl. Formalizing Dijkstra’s predicate transformer wp in weak second-order logic. Theoretical Computer Science, 146(1–2):185–197, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1818.

Bonacina:1995:TFC

[149] Maria Paola Bonacina and Jieh Hsiang. Towards a foundation of completion procedures as semidecision procedures. Theoretical Computer Science, 146(1–2):199–242, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1821.

Backofen:1995:CRF

[150] Rolf Backofen and Gert Smolka. A complete and recursive feature theory. Theoretical Computer Science, 146(1–2):243–268, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1822.

Naughton:1995:ARF

[151] J. F. Naughton, R. Ramakrishnan, Y. Sagiv, and J. D. Ullman. Argument reduction by factoring. Theoretical Computer Science, 146(1–2):269–310, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=146&issue=1-2&aid=1820.

Alessi:1995:FPT

[152] Fabio Alessi, Paolo Baldan, and Gianna Bellè. A fixed-point theorem in a category of compact metric spaces. Theoretical Computer Science, 146(1–2):311–320, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=146&issue=1-2&aid=1955.

Kobayashi:1995:FPM

[153] Yuji Kobayashi. A finitely presented monoid which has solvable word problem but has no regular complete presentation. Theoretical Computer Science, 146(1–2):321–329, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=146&issue=1-2&aid=1898.

Zhang:1995:MSF

[154] Guo-Qiang Zhang. On maximal stable functions. Theoretical Computer Science, 146(1–2):331–339, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1794.

Ingolfsdottir:1995:LES

[155] Anna Ingólfsdóttir. Late and early semantics coincide for testing. Theoretical Computer Science, 146(1–2):341–349, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=146&issue=1-2&aid=1938.

Bampis:1995:OPE

[156] E. Bampis, J.-C König, and D. Trystram. Optimal parallel execution of complete binary trees and grids into most popular interconnection networks. Theoretical Computer Science, 147(1–2):1–18, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=147&issue=1-2&aid=1901.

Gasieniec:1995:ZMR

[157] Leszek Gasieniec, Wojciech Plandowski, and Wojciech Rytter. The zooming method: a recursive approach to timespace efficient string-matching. Theoretical Computer Science, 147(1–2): 19–30, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1995&volume=147&issue=
                                        1-2&aid=1883.

Bodlaender:1995:PCS

[158] Hans L. Bodlaender, Rodney G. Downey, Michael R. Fellows, and Harold T. Wareham. The parameterized complexity of sequence alignment and consensus. Theoretical Computer Science, 147(1–2):31–54, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1885.

Sureson:1995:NCN

[159] Claude Sureson. NP≠co-NP and models of arithmetic. Theoretical Computer Science, 147(1–2):55–67, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1868.

Jansen:1995:MBT

[160] Klaus Jansen and Haiko Müller. The minimum broadcast time problem for several processor networks. Theoretical Computer Science, 147(1–2): 69–85, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=147&issue=
1-2&aid=1864.

Nishida:1995:QDS

[161] Taishin Y. Nishida. Quasi-deterministic 0L systems and their representation. Theoretical Computer Science, 147(1–2):87–116, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1884.

Cheng:1995:CRS

[162] Allan Cheng, Javier Esparza, and Jens Palsberg. Complexity results for 1-safe nets. Theoretical Computer Science, 147(1–2):117–136, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1865.

Zimand:1995:TSC

[163] Marius Zimand. On the topological size of p-m-complete degrees. Theoretical Computer Science, 147(1–2):137–147, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1874.

Schurz:1995:MGF

[164] Gerhard Schurz. Most general first order theorems are not recursively enumerable. Theoretical Computer Science, 147(1–2):149–163, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1863.

Aigrain:1995:PTC

[165] Philippe Aigrain and Danièle Beauquier. Polyomino tilings, cellular automata and codicity. Theoretical Computer Science, 147(1–2):165–180, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1835.

Amaldi:1995:CAF

[166] Edoardo Amaldi and Viggo Kann. The complexity and approximability of finding maximum feasible subsystems of linear relations. Theoretical Computer Science, 147(1–2):181–210, August 7, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1888.

Arques:1995:FSL

[167] Didier Arquès and Olivier Grange. A fast scan-line algorithm for topological filling of well-nested objects in 2.5D digital pictures. Theoretical Computer Science, 147(1–2):211–248, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1778.

Yu:1995:EPA

[168] Chang-Wu Yu and Gen-Huey Chen. Efficient parallel algorithms for doubly convex-bipartite graphs. Theoretical Computer Science, 147(1–2):

249–265, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=147&issue=
1-2&aid=1854.

Blot:1995:ACA

[169] Joël Blot, Wenceslas Fernandez de la Vega, Vangelis Th Paschos, and Rachid Saad. Average case analysis of greedy algorithms for optimisation problems on set systems. Theoretical Computer Science, 147(1–2):267–298, August 07, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=147&issue=1-2&aid=1876.

Diamond:1995:FCS

[170] P. Diamond, P. Kloeden, V. Kozyakin, and A. Pokrovskii. On the fragmentary complexity of symbolic sequences. Theoretical Computer Science, 148(1):1–17, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1895.

Durand:1995:RNC

[171] Bruno Durand. A Random NPcomplete problem for inversion of 2D cellular automata. Theoretical Computer Science, 148(1):19–32, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1873.

Cai:1995:IRM

[172] Liming Cai and Jianer Chen. On input read-modes of alternating Turing machines. Theoretical Computer Science, 148(1):33–55, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1887.

Shoudai:1995:UMI

[173] Takayoshi Shoudai and Satoru Miyano. Using maximal independent sets to solve problems in parallel. Theoretical Computer Science, 148(1):57–65, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1855.

Kumar:1995:ESI

[174] Ratnesh Kumar and Vijay K. Garg. Extremal solutions of inequations over lattices with applications to supervisory control. Theoretical Computer Science, 148(1):67–92, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1890.

Bodlaender:1995:RGP

[175] Hans L. Bodlaender and Klaus Jansen. Restrictions of graph partition problems. Part I. Theoretical Computer Science, 148(1):93–109, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=2016.

Althofer:1995:SPL

[176] Ingo Althöfer and Jörg Bültermann. Superlinear period lengths in some subtraction games. Theoretical Computer Science, 148(1):111–119, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1960.

Arnold:1995:ISC

[177] André Arnold. An initial semantics for the μ-calculus on trees and Rabin’s complementation theorem. Theoretical Computer Science, 148(1):121–132, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=2007.

Dubhashi:1995:FML

[178] Devdatt P. Dubhashi, Grammati E. Pantziou, Paul G. Spirakis, and Christos D. Zaroliagis. The fourth moment in Luby’s distribution. Theoretical Computer Science, 148(1): 133–140, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

cgi?year=1995&volume=148&issue=
1&aid=2015.

Kimber:1995:LLS

[179] Don Kimber and Philip M. Long. Online learning of smooth functions of a single variable. Theoretical Computer Science, 148(1):141–156, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1962.

Morita:1995:RSO

[180] Kenichi Morita. Reversible simulation of one-dimensional irreversible cellular automata. Theoretical Computer Science, 148(1):157–163, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1979.

Zwick:1995:SNW

[181] Uri Zwick. The smallest networks on which the Ford-Fulkerson maximum flow procedure may fail to terminate. Theoretical Computer Science, 148(1):165–170, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=1&aid=1963.

Pinotti:1995:PAP

[182] Maria Cristina Pinotti and Geppino Pucci. Parallel algorithms for priority queue operations. Theoretical Computer Science, 148(1):171–180, August 21, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=148&issue=1&aid=1980.

Enjalbert:1995:P

[183] P. Enjalbert, E. W. Mayr, and K. W. Wagner. Preface. Theoretical Computer Science, 148(2):181–??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alvarez:1995:ADP

[184] Carme Àlvarez and Birgit Jenner. On adaptive DLOGTIME and POLYLOG-TIME reductions. Theoretical Computer Science, 148(2):183–205, September 04, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=148&issue=2&aid=1974.

Borchert:1995:APR

[185] Bernd Borchert. On the acceptance power of regular languages. Theoretical Computer Science, 148(2):207–225, September 04, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=148&issue=2&aid=1973.

Bruyere:1995:SDP

[186] Véronique Bruyère, Clelia De Felice, and Giovanna Guaiana. On some decision problems for trace codings. Theoretical Computer Science, 148(2):227–260, September 04, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=2&aid=1977.

Etessami:1995:RPL

[187] Kousha Etessami and Neil Immerman. Reachability and the power of local ordering. Theoretical Computer Science, 148(2):261–279, September 04, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=2&aid=1975.

Jancar:1995:UBP

[188] Petr Jančar. Undecidability of bisimilarity for Petri nets and some related problems. Theoretical Computer Science, 148(2):281–301, September 04, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=2&aid=1978.

Laroussinie:1995:HTL

[189] F. Laroussinie and Ph. Schnoebelen. A hierarchy of temporal logics with past. Theoretical Computer Science, 148(2):303–324, September 04, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=148&issue=2&aid=1976.

Naik:1995:QTC

[190] Ashish V. Naik, Kenneth W. Regan, and D. Sivakumar. On quasilinear-time complexity theory. Theoretical Computer Science, 148(2):325–349, September 04, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=148&issue=2&aid=1972.

Hull:1995:F

[191] R. Hull. Foreword. Theoretical Computer Science, 149(1):1–??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buneman:1995:PPC

[192] Peter Buneman, Shamim Naqvi, Val Tannen, and Limsoon Wong. Principles of programming with complex objects and collection types. Theoretical Computer Science, 149(1):3–48, September 18, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=149&issue=1&aid=1965.

VandenBussche:1995:EPC

[193] Jan Van den Bussche and Dirk Van Gucht. The expressive power of cardinality-bounded set values in object-based data models. Theoretical Computer Science, 149(1):49– 66, September 18, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.

cgi?year=1995&volume=149&issue=
1&aid=1966.

Grumbach:1995:EPC

[194] Stéphane Grumbach and Christophe Tollu. On the expressive power of counting. Theoretical Computer Science, 149(1):67–99, September 18, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=1&aid=1967.

Abiteboul:1995:CIL

[195] Serge Abiteboul, Moshe Y. Vardi, and Victor Vianu. Computing with infinitary logic. Theoretical Computer Science, 149(1):101–128, September 18, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=1&aid=1968.

Kivinen:1995:AIF

[196] Jyrki Kivinen and Heikki Mannila. Approximate inference of functional dependencies from relations. Theoretical Computer Science, 149(1):129–149, September 18, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=1&aid=1969.

Fekete:1995:HAN

[197] Alan Fekete, Nancy Lynch, and William E. Weihl. Hybrid atomicity for nested transactions. Theoretical Computer Science, 149(1):151–178, September 18, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=1&aid=1970.

Wong:1995:CSS

[198] Man Hon Wong and Divyakant Agrawal. Context-specific synchronization for atomic data types in objectbased databases. Theoretical Computer Science, 149(1):179–199, September 18, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=1&aid=1971.

Brogi:1995:FAC

[199] Antonio Brogi and Franco Turini. Fully abstract compositional semantics for an algebra of logic programs. Theoretical Computer Science, 149(2):201–229, October 02, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=2&aid=1952.

Costantini:1995:CSM

[200] Stefania Costantini. Contributions to the stable model semantics of logic programs with negation. Theoretical Computer Science, 149(2):231–255, October 02, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=2&aid=1893.

Abraham:1995:ICI

[201] Uri Abraham. On interprocess communication and the implementation of multi-writer atomic registers. Theoretical Computer Science, 149(2):257–298, October 02, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=2&aid=1953.

Montanari:1995:LEP

[202] Ugo Montanari and Daniel Yankelevich. Location equivalence in a parametric setting. Theoretical Computer Science, 149(2):299–332, October 02, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=2&aid=1853.

Desharnais:1995:EDS

[203] Jules Desharnais, Nadir Belkhiter, Salah Ben Mohamed Sghaier, Fairouz Tchier, Ali Jaoua, Ali Mili, and Nejib Zaguia. Embedding a demonic semilattice in a relation algebra. Theoretical Computer Science, 149(2):333–360, October 02, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=149&issue=2&aid=1905.

Schmidt-Schauss:1995:MTC

[204] Manfred Schmidt-Schauß, Massimo Marchiori, and Sven Eric Panitz. Modular termination of r-consistent and left-linear term rewriting systems. Theoretical Computer Science, 149(2):361–374, October 02, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=149&issue=2&aid=2021.

Ausiello:1995:ASN

[205] G. Ausiello, P. Crescenzi, and M. Protasi. Approximate solution of NP optimization problems. Theoretical Computer Science, 150(1):1–55, October 16, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=150&issue=1&aid=1925.

Adamek:1995:GFP

[206] Jiří Adámek and Václav Koubek. On the greatest fixed point of a set functor. Theoretical Computer Science, 150(1):57–75, October 16, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=150&issue=1&aid=1951.

Droste:1995:RLC

[207] Manfred Droste. Recognizable languages in concurrency monoids. Theoretical Computer Science, 150(1):77– 109, October 16, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=150&issue=
1&aid=1900.

Naumann:1995:PTH

[208] David A. Naumann. Predicate transformers and higher-order programs. Theoretical Computer Science, 150(1):111–159, October 16, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=150&issue=1&aid=1881.

Gardiner:1995:APC

[209] P. H. B. Gardiner. Algebraic proofs of consistency and completeness. Theoretical Computer Science, 150(1):161–191, October 16, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=150&issue=1&aid=1845.

Hazewinkel:1995:ESI

[210] M. Hazewinkel. Editorial to the Subject Index Volumes 101–150. Theoretical Computer Science, 150(2):195–??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1995:SIV

[211] Anonymous. Subject Index Volumes 101–150. Theoretical Computer Science, 150(2):197–??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1995:RLI

[212] Anonymous. Reference List of Indexed Articles. Theoretical Computer Science, 150(2):315–??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1995:CIV

[213] Anonymous. Cumulative Index Volumes 101–150. Theoretical Computer Science, 150(2):339–??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1995:WTC

[214] Anonymous. Workshop on Topology and Completion in Semantics. Theoretical Computer Science, 151(1):??, November 13, 1995. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gastin:1995:P

[215] P. Gastin and J. J. M. M. Rutten. Preface. Theoretical Computer Science, 151(1):1–??, 1995. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambler:1995:DCM

[216] Simon Ambler, Marta Kwiatkowska, and Nicholas Measor. Duality and the completeness of the modal μcalculus. Theoretical Computer Science, 151(1):3–27, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1995&volume=151&issue=1&aid=1986.

Arnold:1995:TPR

[217] André Arnold. A topological property of rational ω-languages. Theoretical Computer Science, 151(1):29– 36, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=151&issue=
1&aid=1987.

deBoer:1995:NIC

[218] Frank S. de Boer, Alessandra Di Pierro, and Catuscia Palamidessi. Nondeterminism and infinite computations in constraint programming. Theoretical Computer Science, 151(1):37–78, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=1&aid=1988.

Bonsangue:1995:DBS

[219] Marcello M. Bonsangue, Bart Jacobs, and Joost N. Kok. Duality beyond sober spaces: Topological spaces and observation frames. Theoretical Computer Science, 151(1):79–124, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=1&aid=1989.

Courcelle:1995:MSO

[220] Bruno Courcelle. The monadic secondorder logic of graphs IX: Machines and their behaviours. Theoretical Computer Science, 151(1):125–162, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=1&aid=1990.

Edalat:1995:DTI

[221] Abbas Edalat. Domain theory and integration. Theoretical Computer Science, 151(1):163–193, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=1&aid=1991.

Matthews:1995:ETL

[222] S. G. Matthews. An extensional treatment of lazy data flow deadlock. Theoretical Computer Science, 151(1):195–205, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=1&aid=1992.

Mislove:1995:FAR

[223] Michael W. Mislove and Frank J. Oles. Full abstraction and recursion. Theoretical Computer Science, 151(1):207–256, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=1&aid=1993.

Smyth:1995:SMC

[224] M. B. Smyth. Semi-metrics, closure spaces and digital topology. Theoretical Computer Science, 151(1):257– 276, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=151&issue=
1&aid=1994.

Sunderhauf:1995:FCM

[225] Philipp Sünderhauf. A faithful computational model of the real numbers. Theoretical Computer Science, 151(1):277–294, November 13, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=1&aid=1995.

Shyamasundar:1995:F

[226] R. K. Shyamasundar. Foreword. Theoretical Computer Science, 151(2):295– ??, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castagna:1995:MLT

[227] Giuseppe Castagna. A metalanguage for typed object-oriented languages. Theoretical Computer Science, 151(2):297–352, November 27, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=2&aid=2009.

Ait-Kaci:1995:LSC

[228] Hassan Aït-Kaci and Jacques Garrigue. Label-selective λ-calculus syntax and confluence. Theoretical Computer Science, 151(2):353–383, November 27, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=2&aid=2010.

vanBakel:1995:ITA

[229] Steffen van Bakel. Intersection type assignment systems. Theoretical Computer Science, 151(2):385–435, November 27, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=2&aid=2011.

Honda:1995:RBP

[230] Kohei Honda and Nobuko Yoshida. On reduction-based process semantics. Theoretical Computer Science, 151(2):437–486, November 27, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=2&aid=2012.

Rao:1995:MPC

[231] M. R. K. Krishna Rao. Modular proofs for completeness of hierarchical term rewriting systems. Theoretical Computer Science, 151(2):487–512, November 27, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=151&issue=2&aid=2013.

Fredholm:1995:IAF

[232] Daniel Fredholm. Intensional aspects of function definitions. Theoretical Computer Science, 152(1):1–66, December 11, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=1&aid=1902.

Hodges:1995:MSD

[233] Wilfrid Hodges. The meaning of specifications I: Domains and initial models. Theoretical Computer Science, 152(1):67–89, December 11, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=1&aid=1879.

Astesiano:1995:FOE

[234] Egidio Astesiano and Maura Cerioli. Free objects and equational deduction for partial conditional specifications. Theoretical Computer Science, 152(1):91–138, December 11, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=1&aid=1880.

Kurihara:1995:MNT

[235] Masahito Kurihara and Azuma Ohuchi. Modularity in noncopying term rewriting. Theoretical Computer Science, 152(1):139–169, December 11, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=1&aid=1882.

Benveniste:1995:CSS

[236] Albert Benveniste, Bernard C. Levy, Eric Fabre, and Paul Le Guernic. A calculus of stochastic systems for the specification, simulation, and hidden state estimation of mixed stochastic/nonstochastic systems. Theoretical Computer Science, 152(2):171–217, December 25, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=2&aid=1929.

Seidel:1995:PCP

[237] Karen Seidel. Probabilistic communicating processes. Theoretical Computer Science, 152(2):219–249, December 25, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=2&aid=1920.

Aceto:1995:CAT

[238] Luca Aceto and Alan Jeffrey. A complete axiomatization of timed bisimulation for a class of timed regular behaviours. Theoretical Computer Science, 152(2):251–268, December 25, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=2&aid=1919.

Verma:1995:TCR

[239] Rakesh M. Verma. Transformations and confluence for rewrite systems. Theoretical Computer Science, 152(2):269–283, December 25, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=2&aid=1889.

Inverardi:1995:INF

[240] Paola Inverardi and Monica Nesi. Infinite normal forms for non-linear term rewriting systems. Theoretical Computer Science, 152(2):285–303, December 25, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=2&aid=1878.

Banach:1995:LCD

[241] R. Banach. Locating the contractum in the double pushout approach. Theoretical Computer Science, 152(2):305–320, December 25, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1995&volume=152&issue=2&aid=1935.

Mislove:1995:EEN

[242] M. Mislove, M. Nivat, and C. Papadimitnou. Editorial on the Electronic Notes in Theoretical Computer Science. Theoretical Computer Science, 152(2):321–??, 1995. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:1996:P

[243] G. Rozenberg. Preface. Theoretical Computer Science, 153(1–2):1–??, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Petri:1996:NTS

[244] C. A. Petri. Nets, time and space. Theoretical Computer Science, 153(1–2):3–48, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=153&issue=1-2&aid=2057.

Desel:1996:PNM

[245] J. Desel, K.-P. Neuendorf, and M.-D. Radola. Proving nonreachability by modulo-invariants. Theoretical Computer Science, 153(1–2):49–64, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=153&issue=1-2&aid=2058.

Engelfriet:1996:MSP

[246] Joost Engelfriet. A multiset semantics for the pi-calculus with replication. Theoretical Computer Science, 153(1–2):65–94, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=153&issue=1-2&aid=2059.

Esparza:1996:TME

[247] Javier Esparza and Glenn Bruns. Trapping mutual exclusion in the box calculus. Theoretical Computer Science, 153(1–2):95–128, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=153&issue=1-2&aid=2060.

Hoogers:1996:ESS

[248] P. W. Hoogers, H. C. M. Kleijn, and P. S. Thiagarajan. An event structure semantics for general Petri nets. Theoretical Computer Science, 153(1–2):129–170, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=153&issue=1-2&aid=2061.

Meseguer:1996:PVU

[249] José Meseguer, Ugo Montanari, and Vladimiro Sassone. Process versus unfolding semantics for Place/Transition Petri nets. Theoretical Computer Science, 153(1–2):171–210, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=153&issue=1-2&aid=2062.

Nielsen:1996:PNB

[250] Mogens Nielsen and Glynn Winskel. Petri nets and bisimulation. Theoretical Computer Science, 153(1–2):211–244, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=153&issue=1-2&aid=2063.

Smith:1996:BCC

[251] Einar Smith. On the border of causality: Contact and confusion. Theoretical Computer Science, 153(1–2):245–270, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=153&issue=1-2&aid=2064.

Teruel:1996:STE

[252] Enrique Teruel and Manuel Silva. Structure theory of equal conflict systems. Theoretical Computer Science, 153(1–2):271–300, January 08, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=153&issue=1-2&aid=2065.

Lingas:1996:P

[253] A. Lingas. Preface. Theoretical Computer Science, 154(1):1–??, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czumaj:1996:GPN

[254] Artur Czumaj and Alan Gibbons. Guthrie’s problem: new equivalences and rapid reductions. Theoretical Computer Science, 154(1):3–22, January 22, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=1&aid=2067.

Karpinski:1996:RVD

[255] Marek Karpinski and Rutger Verbeek. On randomized versus deterministic computation. Theoretical Computer Science, 154(1):23–39, January 22, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=1&aid=2068.

Jaromczyk:1996:TEF

[256] Jerzy W. Jaromczyk and Grzegorz Świaţek. A theory of even functionals and their algorithmic applications. Theoretical Computer Science, 154(1):41–56, January 22, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=1&aid=2069.

Cohen:1996:PFS

[257] David M. Cohen and Michael L. Fredman. Products of finite state machines with full coverage. Theoretical Computer Science, 154(1):57–65, January 22, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=1&aid=2070.

Ebinger:1996:LDI

[258] Werner Ebinger and Anca Muscholl. Logical definability on infinite traces. Theoretical Computer Science, 154(1): 67–84, January 22, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=154&issue=
                                        1&aid=2071.

Wilke:1996:ACF

[259] Thomas Wilke. An algebraic characterization of frontier testable tree languages. Theoretical Computer Science, 154(1):85–106, January 22, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=154&issue=1&aid=2072.

Jategaonkar:1996:DTC

[260] Lalita Jategaonkar and Albert R. Meyer. Deciding true concurrency equivalences on safe, finite nets. Theoretical Computer Science, 154(1):107–143, January 22, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=154&issue=1&aid=2073.

Sureson:1996:PNC

[261] Claude Sureson. P, NP, Co-NP and weak systems of arithmetic. Theoretical Computer Science, 154(2):145–163, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=2040.

Fiorio:1996:TLT

[262] Christophe Fiorio and Jens Gustedt. Two linear time Union Find strategies for image processing. Theoretical Computer Science, 154(2):165–181, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=1896.

Mitrana:1996:PS

[263] Victor Mitrana, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Pattern systems. Theoretical Computer Science, 154(2):183–201, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=1892.

Idury:1996:MMP

[264] Ramana M. Idury and Alejandro A. Schäffer. Multiple matching of parameterized patterns. Theoretical Computer Science, 154(2):203–224, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=1904.

JaJa:1996:SSC

[265] Joseph F. JáJá, Kwan Woo Ryu, and Uzi Vishkin. Sorting strings and constructing digital search trees in parallel. Theoretical Computer Science, 154(2):225–245, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=154&issue=2&aid=1897.

Engelfriet:1996:CCU

[266] J. Engelfriet, T. Harju, A. Proskurowski, and G. Rozenberg. Characterization and complexity of uniformly nonprimitive labeled 2-structures. Theoretical Computer Science, 154(2):247– 282, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=154&issue=
                                        2&aid=1906.

Blundo:1996:IRS

[267] Carlo Blundo, Alfredo De Santis, Luisa Gargano, and Ugo Vaccaro. On the information rate of secret sharing schemes. Theoretical Computer Science, 154(2):283–306, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=154&issue=2&aid=2003.

Calude:1996:ECM

[268] Cristian Calude and Marius Zimand. Effective category and measure in abstract complexity theory. Theoretical Computer Science, 154(2):307–327, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=2004.

LafayedeMicheaux:1996:CGT

[269] N. Lafaye de Micheaux and C. Rambaud. Confluence for graph transformations. Theoretical Computer Science, 154(2):329–348, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=1947.

Barua:1996:GGT

[270] Rana Barua and S. Ramakrishnan. σ-game, σ^{+}-game and twodimensional additive cellular automata. Theoretical Computer Science, 154(2):349–366, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=154&issue=2&aid=2032.

Hemaspaandra:1996:OA

[271] Lane A. Hemaspaandra and Leen Torenvliet. Optimal advice. Theoretical Computer Science, 154(2):367– 377, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

                                        cgi?year=1996&volume=154&issue=
                                        2&aid=2017.

Ramesh:1996:DCC

[272] S. Ramesh and Bommadevara N. Srinivas. A direct characterization of completion. Theoretical Computer Science, 154(2):379–385, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=154&issue=2&aid=2029.

Justin:1996:CPS

[273] J. Justin and G. Pirillo. On a combinatorial property of Sturmian words. Theoretical Computer Science, 154(2):387–394, February 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=154&issue=2&aid=2018.

Bloom:1996:FPO

[274] Stephen L. Bloom and Zoltán Ésik. Fixed point operations on CCC’s. Part I. Theoretical Computer Science, 155(1):1–38, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=155&issue=1&aid=1950.

Sangiorgi:1996:LIS

[275] Davide Sangiorgi. Locality and interleaving semantics in calculi for mobile processes. Theoretical Computer Science, 155(1):39–83, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=1&aid=1961.

Kamareddine:1996:UN

[276] Fairouz Kamareddine and Rob Nederpelt. A useful λ-notation. Theoretical Computer Science, 155(1):85– 109, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=155&issue=
1&aid=2042.

Immerman:1996:EFF

[277] Neil Immerman, Sushant Patnaik, and David Stemple. The expressiveness of a family of finite set languages. Theoretical Computer Science, 155(1):111–140, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=1&aid=1921.

Fourneau:1996:NMC

[278] Jean-Michel Fourneau, Erol Gelenbe, and Rina Suros. G-networks with multiple classes of negative and positive customers. Theoretical Computer Science, 155(1):141–156, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=1&aid=1959.

Kagan:1996:CMM

[279] Vadim Kagan, Anil Nerode, and V. S. Subrahmanian. Computing minimal models by partial instantiation. Theoretical Computer Science, 155(1):157–177, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=155&issue=1&aid=1850.

Nielson:1996:CPA

[280] Flemming Nielson and Hanne Riis Nielson. From CML to its process algebra. Theoretical Computer Science, 155(1):179–219, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=155&issue=1&aid=1957.

Zhang:1996:QPA

[281] Guo-Qiang Zhang. Quasi-prime algebraic domains. Theoretical Computer Science, 155(1):221–264, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=155&issue=1&aid=2074.

Bunder:1996:TTN

[282] M. W. Bunder and J. R. Hindley. Two β = λ-terms with no types in common. Theoretical Computer Science, 155(1):265–266, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=1&aid=2088.

Drakengren:1996:USR

[283] Thomas Drakengren. Uniqueness of Scott’s reflexive domain in P ω. Theoretical Computer Science, 155(1):267– 276, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=155&issue=
1&aid=2043.

Zhang:1996:NMS

[284] Wenhui Zhang. Number of models and satisfiability of sets of clauses. Theoretical Computer Science, 155(1):277–288, February 26, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=1&aid=2085.

Kuoherov:1996:VA

[285] G. Kuoherov, P. Lescanne, and P. Mosses. Valentin Antimirov (1961– 1995). Theoretical Computer Science, 155(2):239–??, 1996. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kucherov:1996:VA

[286] Gregory Kucherov, Pierre Lescanne, and Peter Mosses. Valentin Antimirov (1961–1995). Theoretical Computer Science, 155(2):289–290, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=2122.

Antimirov:1996:PDR

[287] Valentin Antimirov. Partial derivatives of regular expressions and finite automaton constructions. Theoretical Computer Science, 155(2): 291–319, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=155&issue=
2&aid=2123.

Barcucci:1996:RCP

[288] Elena Barcucci, Alberto Del Lungo, Maurice Nivat, and Renzo Pinzani. Reconstructing convex polyominoes from horizontal and vertical projections. Theoretical Computer Science, 155(2):321–347, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=1927.

Keller:1996:FRP

[289] Jörg Keller. Fast rehashing in PRAM emulations. Theoretical Computer Science, 155(2):349–363, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=1926.

Lange:1996:MDM

[290] Steffen Lange, Thomas Zeugmann, and Shyam Kapur. Monotonic and dual monotonic language learning. Theoretical Computer Science, 155(2):365–410, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=1918.

Iwama:1996:TLB

[291] Kazuo Iwama, Chuzo Iwamoto, and Manzur Morshed. Time lower bounds do not exist for CRCW PRAMs. Theoretical Computer Science, 155(2):411–424, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=1933.

Ehrenfeucht:1996:NBG

[292] Andrzej Ehrenfeucht, Paulien ten Pas, and Grzegorz Rozenberg. A note on binary grammatical codes of trees. Theoretical Computer Science, 155(2):425–438, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=2034.

Berstel:1996:LLB

[293] Jean Berstel and Jean-Eric Pin. Local languages and the Berry–Sethi algorithm. Theoretical Computer Science, 155(2):439–446, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=2045.

Hemaspaandra:1996:RCS

[294] Lane A. Hemaspaandra, Albrecht Hoene, and Mitsunori Ogihara. Reducibility classes of P -selective sets. Theoretical Computer Science, 155(2):447–457, March 11, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=155&issue=2&aid=2035.

See erratum [1424].

Nehlig:1996:AQA

[295] Philippe Nehlig. Applications quasi affines: pavages par images réciproques. (French) [Quasi-affine applications: tilings for reciprocal images]. Theoretical Computer Science, 156(1–2): 1–38, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=156&issue=
1-2&aid=1981.

Kemp:1996:BST

[296] Rainer Kemp. Binary search trees constructed from nondistinct keys with/without specified probabilities. Theoretical Computer Science, 156(1–2):39–70, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=1936.

Gyenizse:1996:CDB

[297] Pál Gyenizse and Sándor Vágvölgyi. Compositions of deterministic bottomup, top-down, and regular look-ahead tree transformations. Theoretical Computer Science, 156(1–2):71–97, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=1937.

Krause:1996:GAY

[298] Matthias Krause. Geometric arguments yield better bounds for threshold circuits and distributed computing. Theoretical Computer Science, 156(1–2):99–117, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=1945.

Bedon:1996:FAO

[299] Nicolas Bedon. Finite automata and ordinals. Theoretical Computer Science, 156(1–2):119–144, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=1946.

Louscou-Bozapalidou:1996:SCT

[300] Olympia Louscou-Bozapalidou. Stochastically costed tree automata: Turakainen’s theorem. Theoretical Computer Science, 156(1–2):145–158, March 25, 1996. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=156&issue=
                                        1-2&aid=1949.

Francon:1996:TPA

[301] Jean Françon. Sur la topologie d’un plan arithmétique. (French) [On the topology of an arithmetic plan]. Theoretical Computer Science, 156(1–2):159–176, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=156&issue=1-2&aid=1997.

Schwartzbach:1996:SCH

[302] Michael I. Schwartzbach. Static correctness of hierarchical procedures. Theoretical Computer Science, 156(1–2):177–201, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=156&issue=1-2&aid=2027.

Forti:1996:GCH

[303] Marco Forti and Furio Honsell. A general construction of hyperuniverses. Theoretical Computer Science, 156(1–2):203–215, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=2028.

Takahashi:1996:UCN

[304] Yoshikane Takahashi. A unified constructive network model for problemsolving. Theoretical Computer Science, 156(1–2):217–261, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=2075.

Kim:1996:NVD

[305] Yonghoan Kim. New values in Domineering. Theoretical Computer Science, 156(1–2):263–280, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=2091.

Terrier:1996:LRR

[306] Véronique Terrier. Language not recognizable in real time by one-way cellular automata. Theoretical Computer Science, 156(1–2):281–287, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=156&issue=1-2&aid=2052.

Arnold:1996:ACO

[307] André Arnold and Ilaria Castellani. An algebraic characterization of observational equivalence. Theoretical Computer Science, 156(1–2):289– 299, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=156&issue=
                                        1-2&aid=2082.

Manzini:1996:OSL

[308] Giovanni Manzini. On the ordering of sparse linear systems. Theoretical Computer Science, 156(1–2): 301–313, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=156&issue=
                                        1-2&aid=2092.

DePrisco:1996:NLB

[309] Roberto De Prisco and Alfredo De Santis. New lower bounds on the cost of binary search trees. Theoretical Computer Science, 156(1–2): 315–325, March 25, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=156&issue=
                                        1-2&aid=2094.

Anonymous:1996:WAC

[310] Anonymous. Workshop on Algorithmic Complexity of Algebraic and Geometric Models. Theoretical Computer Science, 157(1):??, April 09, 1996.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bronstein:1996:IPL

[311] Manuel Bronstein and Marko Petkovšek. An introduction to pseudo-linear algebra. Theoretical Computer Science, 157(1):3–33, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=157&issue=
1&aid=2114.

Devillers:1996:IRC

[312] Olivier Devillers. An introduction to randomization in computational geometry. Theoretical Computer Science, 157(1):35–52, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=1&aid=2115.

Elihai:1996:FHO

[313] Y. Elihai and Y. Yomdin. Flexible high-order discretization of geometric data for global motion planning. Theoretical Computer Science, 157(1):53–77, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=1&aid=2116.

Grigoriev:1996:NSS

[314] D. Grigoriev. NC solving of a system of linear ordinary differential equations in several unknowns. Theoretical Computer Science, 157(1):79–90, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=1&aid=2117.

Grigoriev:1996:CAC

[315] Dima Grigoriev and Marek Karpinski. Computability of the additive complexity of algebraic circuits with root extracting. Theoretical Computer Science, 157(1):91–99, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=1&aid=2118.

Laumond:1996:NSC

[316] Jean-Paul Laumond and Jean-Jacques Risler. Nonholonomic systems: Controllability and complexity. Theoretical Computer Science, 157(1): 101–114, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=157&issue=
1&aid=2119.

Ollagnier:1996:AMD

[317] Jean Moulin Ollagnier. Algorithms and methods in differential algebra. Theoretical Computer Science, 157(1):115–127, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=1&aid=2120.

Risler:1996:BDN

[318] Jean-Jacques Risler. A bound for the degree of nonholonomy in the plane. Theoretical Computer Science, 157(1):129–136, April 09, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=1&aid=2121.

Ablayev:1996:LBO

[319] Farid Ablayev. Lower bounds for one-way probabilistic communication complexity and their application to space complexity. Theoretical Computer Science, 157(2):139–159, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2098.

Burago:1996:CPO

[320] Dima Burago, Michel de Rougemont, and Anatol Slissenko. On the complexity of partially observed Markov decision processes. Theoretical Computer Science, 157(2):161–183, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2099.

Grigoriev:1996:CLB

[321] D. Grigoriev and N. Vorobjov. Complexity lower bounds for computation trees with elementary transcendental function gates. Theoretical Computer Science, 157(2):185–214, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2100.

Karzanov:1996:HTS

[322] Alexander V. Karzanov. How to tidy up a symmetric set-system by use of uncrossing operations. Theoretical Computer Science, 157(2):215–225, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2101.

Muchnik:1996:GMC

[323] Andrei A. Muchnik and Nikolai K. Vereshchagin. A general method to construct oracles realizing given relationships between complexity classes. Theoretical Computer Science, 157(2): 227–258, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=157&issue=
2&aid=2102.

Karpinski:1996:SAP

[324] Marek Karpinski and Igor Shparlinski. On some approximation problems concerning sparse polynomials over finite fields. Theoretical Computer Science, 157(2):259–266, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2103.

Levin:1996:CCF

[325] Leonid A. Levin. Computational complexity of functions. Theoretical Computer Science, 157(2):267–271, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2104.

Shparlinski:1996:FPR

[326] Igor Shparlinski. On finding primitive roots in finite fields. Theoretical Computer Science, 157(2):273–275, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2105.

Verbitsky:1996:TPR

[327] Oleg Verbitsky. Towards the parallel repetition conjecture. Theoretical Computer Science, 157(2):277–282, May 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=157&issue=2&aid=2106.

Gabarro:1996:DPD

[328] Joaquim Gabarró, Conrado Martínez, and Xavier Messeguer. A design of a parallel dictionary using skip lists. Theoretical Computer Science, 158(1–2):1–33, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=1922.

Koskas:1996:APW

[329] Michel Koskas. About the ppaperfolding words. Theoretical Computer Science, 158(1–2):35–51, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=1948.

Ferreira:1996:RRG

[330] Afonso Ferreira and Miltos Grammatikakis. Randomized routing on generalized hypercubes. Theoretical Computer Science, 158(1–2):53–64, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=1982.

Fabre:1996:DSN

[331] Stéphane Fabre. Dépendance de systèmes de numération associés à des puissances d’un nombre de Pisot. (French) [Dependence of number systems associated with the influences of a Pisot number]. Theoretical Computer Science, 158(1–2):65–79, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=1983.

Jonoska:1996:SSS

[332] Nataša Jonoska. Sofic shifts with synchronizing presentations. Theoretical Computer Science, 158(1–2):81–115, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=1996.

Demange:1996:AMF

[333] Marc Demange and Vangelis Th Paschos. On an approximation measure founded on the links between optimization and polynomial approximation theory. Theoretical Computer Science, 158(1–2):117–141, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=1998.

Hirshfeld:1996:PAD

[334] Yoram Hirshfeld, Mark Jerrum, and Faron Moller. A polynomial algorithm for deciding bisimilarity of normed context-free processes. Theoretical Computer Science, 158(1–2):143–159, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=2002.

Nishida:1996:SL

[335] Taishin Y. Nishida and Arto Salomaa. Slender 0L languages. Theoretical Computer Science, 158(1–2):161–176, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=158&issue=1-2&aid=2005.

Breslauer:1996:SCC

[336] Dany Breslauer. Saving comparisons in the Crochemore-Perrin string matching algorithm. Theoretical Computer Science, 158(1–2):177–192, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=158&issue=1-2&aid=2006.

Agrawal:1996:GSL

[337] M. Agrawal and V. Arvind. Geometric sets of low information content. Theoretical Computer Science, 158(1–2):193–219, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=158&issue=1-2&aid=2014.

Mocas:1996:SCE

[338] Sarah E. Mocas. Separating classes in the exponential-time hierarchy from classes in PH. Theoretical Computer Science, 158(1–2):221–231, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=158&issue=1-2&aid=2019.

Ramalingam:1996:CCD

[339] G. Ramalingam and Thomas Reps. On the computational complexity of dynamic graph problems. Theoretical Computer Science, 158(1–2):233–277, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=2020.

Takahashi:1996:SOP

[340] Yoshikane Takahashi. Solving optimization problems with variableconstraint by an extended Cohen– Grossberg model. Theoretical Computer Science, 158(1–2):279–341, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=2026.

Zwick:1996:CMP

[341] Uri Zwick and Mike S. Paterson. The complexity of mean payoff games on graphs. Theoretical Computer Science, 158(1–2):343–359, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=158&issue=1-2&aid=2129.

Agrawal:1996:QLT

[342] M. Agrawal and V. Arvind. Quasilinear truth-table reductions to pselective sets. Theoretical Computer Science, 158(1–2):361–370, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=158&issue=1-2&aid=2108.

Burckel:1996:CIC

[343] Serge Burckel. Closed iterative calculus. Theoretical Computer Science, 158(1–2):371–378, May 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=158&issue=1-2&aid=2112.

Anonymous:1996:CEBa

[344] Anonymous. Contents EATCS Bulletin Number 58. Theoretical Computer Science, 158(1–2):379–??, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gouyou-Beauchamps:1996:PSI

[345] D. Gouyou-Beauchamps and J.-G. Penaud. Preface to the Special Issue on GASCOM ’94. Theoretical Computer Science, 159(1):3–??, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wilf:1996:RDP

[346] Herbert S. Wilf. Recents dévéloppements et problèmes dans le domaine de la génération aléatoire. (French) [recent developments and problems in the area of random generation]. Theoretical Computer Science, 159(1):5–13, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=1&aid=2138.

Alonso:1996:PAG

[347] Laurent Alonso and René Schott. A parallel algorithm for the generation of a permutation and applications. Theoretical Computer Science, 159(1):15–28, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=1&aid=2139.

Barcucci:1996:DPP

[348] Elena Barcucci, Alberto Del Lungo, and Renzo Pinzani. “Deco” polyominoes, permutations and random generation. Theoretical Computer Science, 159(1):29–42, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=1&aid=2140.

Denise:1996:GAU

[349] Alain Denise. Génération aléatoire uniforme de mots de langages rationnels. (French) [Generating uniformly random words of rational languages]. Theoretical Computer Science, 159(1):43–63, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=1&aid=2141.

Louchard:1996:PAS

[350] G. Louchard. Probabilistic analysis of some (un)directed animals. Theoretical Computer Science, 159(1):65–79, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=159&issue=1&aid=2142.

Mosbah:1996:PHR

[351] Mohamed Mosbah. Probabilistic hyperedge replacement grammars. Theoretical Computer Science, 159(1): 81–102, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=159&issue=
                                        1&aid=2143.

Aigrain:1996:PSI

[352] P. Aigrain. Preface to the Special Issue on the 3rd International Workshop on Polyominoes and Tilings. Theoretical Computer Science, 159(1):103– 104, May 28, 1996. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fournier:1996:PFP

[353] J. C. Fournier. Pavage des figures planes sans trous par des dominos: Fondement graphique de l’algorithme de Thurston, parallelisation, unicité et décomposition. (French) [Tiling pictures of the plane without holes by dominoes: graphical basis of the Thurston algorithm, parallelisation, uniqueness and decomposition]. Theoretical Computer Science, 159(1):105–128, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=1&aid=2145.

Barcucci:1996:PDT

[354] Elena Barcucci, Alberto Del Lungo, Renzo Pinzani, and Renzo Sprugnoli. Polyominoes defined by their vertical and horizontal projections. Theoretical Computer Science, 159(1):129–136, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=1&aid=2146.

Chaboud:1996:DTP

[355] Thomas Chaboud. Domino tiling in planar graphs with regular and bipartite dual. Theoretical Computer Science, 159(1):137–142, May 28, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=1&aid=2147.

Bundgen:1996:BAT

[356] Reinhard Bündgen. Buchberger’s algorithm: The term rewriter’s point of view. Theoretical Computer Science, 159(2):143–190, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=2&aid=1999.

Asperti:1996:ISI

[357] Andrea Asperti and Cosimo Laneve. Interaction systems II: The practice of optimal reductions. Theoretical Computer Science, 159(2):191–244, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=159&issue=2&aid=2000.

Gritzner:1996:RAM

[358] Thomas F. Gritzner and Rudolf Berghammer. A relation algebraic model of robust correctness. Theoretical Computer Science, 159(2):245–270, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=159&issue=2&aid=2001.

Erdogmus:1996:OSN

[359] Hakan Erdogmus, Robert Johnston, and Michael Ferguson. On the operational semantics of nondeterminism and divergence. Theoretical Computer Science, 159(2):271–317, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=159&issue=2&aid=2107.

Sambin:1996:CDT

[360] Giovanni Sambin, Silvio Valentini, and Paolo Virgili. Constructive domain theory as a branch of intuitionistic pointfree topology. Theoretical Computer Science, 159(2):319–341, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=2&aid=2110.

Arai:1996:PHP

[361] Noriko H. Arai. A proper hierarchy of propositional sequent calculi. Theoretical Computer Science, 159(2):343–354, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=2&aid=2134.

Ying:1996:WIC

[362] Mingsheng Ying. When is the ideal completion of abstract basis algebraic. Theoretical Computer Science, 159(2):355–356, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=2&aid=2124.

Maddux:1996:RAS

[363] Roger D. Maddux. Relation-algebraic semantics. Theoretical Computer Science, 160(1–2):1–85, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2023.

Courcelle:1996:MSO

[364] Bruno Courcelle. The monadic second-order logic of graphs X: Linear orderings. Theoretical Computer Science, 160(1–2):87–143, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2024.

Tronci:1996:EPCa

[365] Enrico Tronci. Equational programming in λ-calculus via SL-systems. Part 1. Theoretical Computer Science, 160(1–2):145–184, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2046.

Tronci:1996:EPCb

[366] Enrico Tronci. Equational programming in λ-calculus via SL-systems. Part 2. Theoretical Computer Science, 160(1–2):185–216, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2047.

Tuijn:1996:CCG

[367] Chris Tuijn and Marc Gyssens. CGOOD, a categorical graph-oriented object data model. Theoretical Computer Science, 160(1–2):217–239, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2030.

Baaz:1996:CFO

[368] Matthias Baaz, Alexander Leitsch, and Richard Zach. Completeness of a first-order temporal logic with time-gaps. Theoretical Computer Science, 160(1–2):241–270, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2048.

Kaminski:1996:PAO

[369] Michael Kaminski and Chung Kei Wong. The power of the “always” operator in first-order temporal logic. Theoretical Computer Science, 160(1–2):271–281, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2049.

Yamasaki:1996:SRN

[370] Susumu Yamasaki. SLDNF resolution with non-safe rule and fixpoint semantics for general logic programs. Theoretical Computer Science, 160(1–2):283–303, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=160&issue=1-2&aid=2050.

Durand:1996:FOS

[371] Arnaud Durand and Solomampionona Ranaivoson. First-order spectra with one binary predicate. Theoretical Computer Science, 160(1–2):305–320, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=160&issue=1-2&aid=2053.

Bonatti:1996:QDD

[372] Piero A. Bonatti and Thomas Eiter. Querying disjunctive databases through nonmonotonic logics. Theoretical Computer Science, 160(1–2):321–363, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=160&issue=1-2&aid=2133.

Marriott:1996:NBC

[373] Kim Marriott and Martin Odersky. Negative Boolean constraints. Theoretical Computer Science, 160(1–2): 365–380, June 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=160&issue=
                                        1-2&aid=2150.

Anonymous:1996:MIVa

[374] Anonymous. Master Index Volume 151–160. Theoretical Computer Science, 160(1–2):383–??, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:1996:PCM

[375] Zhi-Zhong Chen. Parallel constructions of maximal path sets and applications to short superstrings. Theoretical Computer Science, 161(1–2):1–21, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2051.

Carrere:1996:KRA

[376] F. Carrere. On the Kleijn–Rozenberg k-adjacent languages. Theoretical Computer Science, 161(1–2):23–68, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2025.

Labhalla:1996:ACR

[377] Salah Labhalla, Henri Lombardi, and Roger Marlin. Algorithmes de calcul de la réduction de Hermite d’une matrice à coefficients polynomiaux. (French) [Algorithms for computing an Hermitian reduction of a matrix with polynomial coefficients]. Theoretical Computer Science, 161(1–2):69–92, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2031.

Harju:1996:FPC

[378] Tero Harju, Marjo Lipponen, and Alexandru Mateescu. Flatwords and Post Correspondence Problem. Theoretical Computer Science, 161(1–2):93–108, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=161&issue=1-2&aid=2033.

Mayer:1996:CPI

[379] Alain J. Mayer and Larry J. Stockmeyer. The complexity of PDL with interleaving. Theoretical Computer Science, 161(1–2):109–122, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=161&issue=1-2&aid=2036.

Fortnow:1996:RBI

[380] Lance Fortnow and Martin Kummer. On resource-bounded instance complexity. Theoretical Computer Science, 161(1–2):123–140, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=161&issue=1-2&aid=2038.

Petersen:1996:LPW

[381] H. Petersen. On the language of primitive words. Theoretical Computer Science, 161(1–2):141–156, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=161&issue=1-2&aid=2039.

Selmi:1996:TL

[382] Carla Selmi. Over testable languages. Theoretical Computer Science, 161(1–2):157–190, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2041.

Carton:1996:CA

[383] Olivier Carton. Chain automata. Theoretical Computer Science, 161(1–2): 191–203, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=161&issue=
1-2&aid=2044.

Hotz:1996:PCC

[384] Günter Hotz and Gisela Pitsch. On parsing coupled-context-free languages. Theoretical Computer Science, 161(1–2):205–233, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2055.

Fulk:1996:LPI

[385] Mark Fulk and Sanjay Jain. Learning in the presence of inaccurate information. Theoretical Computer Science, 161(1–2):235–261, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2076.

Regan:1996:ISP

[386] Kenneth W. Regan. Index sets and presentations of complexity classes. Theoretical Computer Science, 161(1–2):263–287, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2087.

Maruyama:1996:ITW

[387] Osamu Maruyama and Satoru Miyano. Inferring a tree from walks. Theoretical Computer Science, 161(1–2):289–300, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2097.

Cucker:1996:GKP

[388] Felipe Cucker and Michael Shub. Generalized Knapsack problems and fixed degree separations. Theoretical Computer Science, 161(1–2):301–306, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2205.

Andreev:1996:CHD

[389] Alexander E. Andreev, Andrea E. F. Clementi, and José D. P. Rolim. Constructing the highest degree subgraph for dense graphs is in NCAS. Theoretical Computer Science, 161(1–2):307–314, July 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1996&volume=161&issue=1-2&aid=2217.

Vigna:1996:RBD

[390] Sebastiano Vigna. On the relations between distributive computability and the BSS model. Theoretical Computer Science, 162(1):5–21, August 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=1&aid=2188.

Moore:1996:RTR

[391] Cristopher Moore. Recursion theory on the reals and continuoustime computation. Theoretical Computer Science, 162(1):23–44, August 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=1&aid=2189.

Brattka:1996:RCC

[392] Vasco Brattka. Recursive characterization of computable real-valued functions and relations. Theoretical Computer Science, 162(1):45–77, August 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=1&aid=2190.

Escardo:1996:PER

[393] Martín Hötzel Escardó. PCF extended with real numbers. Theoretical Computer Science, 162(1):79–115, August 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=162&issue=1&aid=2191.

Pelibossian:1996:LLC

[394] Hratchia Pélibossian. On the linearity of on-line computable functions. Theoretical Computer Science, 162(1):117–132, August 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=162&issue=1&aid=2192.

Revol:1996:PEA

[395] Nathalie Revol and Jean-Louis Roch. Parallel evaluation of arithmetic circuits. Theoretical Computer Science, 162(1):133–150, August 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=162&issue=1&aid=2193.

Mikov:1996:LSA

[396] Alexander I. Mikov. Large-scale addition of machine real numbers: Accuracy estimates. Theoretical Computer Science, 162(1):151–170, August 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=162&issue=1&aid=2194.

Pan:1996:PCP

[397] Victor Y. Pan. Parallel computation of polynomial GCD and some related parallel computations over abstract fields. Theoretical Computer Science, 162(2):173–223, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2233.

Baker:1996:MES

[398] Brenda S. Baker and Edward G. Coffman, Jr. Mutual exclusion scheduling. Theoretical Computer Science, 162(2):225–243, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2234.

MeyeraufderHeide:1996:ESR

[399] Friedhelm Meyer auf der Heide, Christian Scheideler, and Volker Stemann. Exploiting storage redundancy to speed up randomized shared memory simulations. Theoretical Computer Science, 162(2):245–281, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2235.

Mayr:1996:DCA

[400] Ernst W. Mayr and Ralph Werchner. Divide-and-conquer algorithms on the hypercube. Theoretical Computer Science, 162(2):283–296, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2236.

Hsu:1996:EMP

[401] Tsan sheng Hsu and Vijaya Ramachandran. Efficient massively parallel implementation of some combinatorial algorithms. Theoretical Computer Science, 162(2):297–322, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2237.

Finta:1996:SUU

[402] Lucian Finta, Zhen Liu, Ioannis Milis, and Evripidis Bampis. Scheduling UET–UCT series–parallel graphs on two processors. Theoretical Computer Science, 162(2):323–340, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2238.

Chakrabarti:1996:RAJ

[403] Soumen Chakrabarti. Random allocation of jobs with weights and precedence. Theoretical Computer Science, 162(2):341–349, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2239.

Garcia:1996:MPA

[404] Miguel Angel García. Massively parallel approximation of irregular triangular meshes of arbitrary topology with smooth parametric surfaces. Theoretical Computer Science, 162(2):351–369, August 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=162&issue=2&aid=2240.

Courcelle:1996:BNU

[405] Bruno Courcelle. Basic notions of universal algebra for language theory and graph grammars. Theoretical Computer Science, 163(1–2):1–54, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2086.

Bloom:1996:FSA

[406] Stephen L. Bloom and Zoltán Ésik. Free shuffle algebras in language varieties. Theoretical Computer Science, 163(1–2):55–98, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2171.

Niwinski:1996:GC

[407] Damian Niwiński and Igor Walukiewicz. Games for the μ-calculus. Theoretical Computer Science, 163(1–2): 99–116, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=163&issue=
1-2&aid=2077.

Colussi:1996:HCC

[408] Livio Colussi and Laura Toniolo. How the character comparison order shapes the shift function of on-line pattern matching algorithms. Theoretical Computer Science, 163(1–2):117–144, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2078.

Derencourt:1996:TWC

[409] Denis Derencourt. A three-word code which is not prefix–suffix composed. Theoretical Computer Science, 163(1–2):145–160, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2080.

Hausen-Tropper:1996:FTA

[410] E. Hausen-Tropper. A framework for a theory of automated learning. Theoretical Computer Science, 163(1–2):161–176, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2084.

Beigel:1996:FCB

[411] Richard Beigel, William Gasarch, and Efim Kinber. Frequency computation and bounded queries. Theoretical Computer Science, 163(1–2):177–192, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2090.

Lehr:1996:VSA

[412] Siegfried Lehr, Jeffrey Shallit, and John Tromp. On the vector space of the automatic reals. Theoretical Computer Science, 163(1–2):193–210, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2175.

Levcopoulos:1996:EFI

[413] Christos Levcopoulos and Ola Petersson. Exploiting few inversions when sorting: Sequential and parallel algorithms. Theoretical Computer Science, 163(1–2):211–238, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2197.

Gu:1996:OHA

[414] Xunrang Gu and Yuzhang Zhu. Optimal heapsort algorithm. Theoretical Computer Science, 163(1–2):239– 243, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=163&issue=
1-2&aid=2174.

Vollmer:1996:RTC

[415] Heribert Vollmer and Klaus W. Wagner. Recursion theoretic characterizations of complexity classes of counting functions. Theoretical Computer Science, 163(1–2):245–258, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2178.

Long:1996:NGC

[416] Dongyang Long. Note on group codes. Theoretical Computer Science, 163(1–2):259–267, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2179.

Fredholm:1996:CMP

[417] Daniel Fredholm. Computing minimum with primitive recursion over lists. Theoretical Computer Science, 163(1–2):269–276, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2184.

Gilman:1996:SLI

[418] Robert H. Gilman. A shrinking lemma for indexed languages. Theoretical Computer Science, 163(1–2): 277–281, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=163&issue=
1-2&aid=2185.

Tsukiji:1996:SCB

[419] Tatsuie Tsukiji. On a small class of Boolean sums. Theoretical Computer Science, 163(1–2):283–289, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2210.

Blanchard:1996:DBC

[420] F. Blanchard and A. Maass. Dynamical behaviour of Coven’s aperiodic cellular automata. Theoretical Computer Science, 163(1–2):291–302, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=163&issue=1-2&aid=2221.

Malgouyres:1996:TNL

[421] Rémy Malgouyres. There is no local characterization of separating and thin objects in Z^{3}. Theoretical Computer Science, 163(1–2):303–308, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=163&issue=1-2&aid=2224.

Ziadi:1996:REL

[422] Djelloul Ziadi. Regular expression for a language without empty word. Theoretical Computer Science, 163(1–2):309–315, August 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=163&issue=1-2&aid=2231.

Carlsson:1996:HB

[423] Svante Carlsson, Jingsen Chen, and Christer Mattsson. Heaps with bits. Theoretical Computer Science, 164(1–2):1–12, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2093.

Case:1996:ALH

[424] John Case, Sanjay Jain, and Arun Sharma. Anomalous learning helps succinctness. Theoretical Computer Science, 164(1–2):13–28, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2095.

Harrison:1996:ACL

[425] John Harrison. On almost cylindrical languages and the decidability of the D0L and PWD0L primitivity problems. Theoretical Computer Science, 164(1–2):29–40, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2109.

Sirvent:1996:RBD

[426] Víctor F. Sirvent. Relationships between the dynamical systems associated to the Rauzy substitutions. Theoretical Computer Science, 164(1–2):41–57, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2111.

Picouleau:1996:WCA

[427] C. Picouleau. Worst-case analysis of fast heuristics for packing squares into a square. Theoretical Computer Science, 164(1–2):59–72, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2125.

Middendorf:1996:TDP

[428] Martin Middendorf. Two-dimensional partitioning problems. Theoretical Computer Science, 164(1–2):73– 106, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

                                        cgi?year=1996&volume=164&issue=
                                        1-2&aid=2126.

Diks:1996:RCF

[429] Krzysztof Diks and Andrzej Pelc. Reliable computations on faulty EREW PRAM. Theoretical Computer Science, 164(1–2):107–122, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2127.

Salomaa:1996:SEG

[430] Kai Salomaa, Derick Wood, and Sheng Yu. Structural equivalence and ET0L grammars. Theoretical Computer Science, 164(1–2):123–140, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2128.

Lutz:1996:CVK

[431] Jack H. Lutz and Elvira Mayordomo. Cook versus Karp-Levin: Separating completeness notions if NP is not small. Theoretical Computer Science, 164(1–2):141–163, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2130.

Hubert:1996:PCS

[432] Pascal Hubert. Propriétés combinatoires des suites définies par le billard dans les triangles pavants. (French)

[Combinatorial properties of sequences defined by a billiard in paved triangles]. Theoretical Computer Science, 164(1–2):165–183, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2149.

Fill:1996:LRC

[433] James Allen Fill. Limits and rates of convergence for the distribution of search cost under the move-to-front rule. Theoretical Computer Science, 164(1–2):185–206, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2151.

Finkel:1996:PAM

[434] Alain Finkel and Isabelle Tellier. A polynomial algorithm for the membership problem with categorial grammars. Theoretical Computer Science, 164(1–2):207–221, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2152.

DeFelice:1996:AHF

[435] Clelia De Felice. An application of Hajós factorizations to variablelength codes. Theoretical Computer Science, 164(1–2):223–252, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2164.

Moews:1996:CSG

[436] David Moews. Coin-sliding and Go. Theoretical Computer Science, 164(1–2):253–276, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2291.

Tomescu:1996:AAL

[437] Ioan Tomescu. On the asymptotic average length of a maximum common subsequence for words over a finite alphabet. Theoretical Computer Science, 164(1–2):277–285, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2200.

Gupta:1996:CSI

[438] Arvind Gupta and Naomi Nishimura. The complexity of subgraph isomorphism for classes of partial ktrees. Theoretical Computer Science, 164(1–2):287–298, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=164&issue=1-2&aid=2254.

Iliopoulos:1996:WTO

[439] Costas S. Iliopoulos and Kunsoo Park. A work-time optimal algorithm for computing all string covers. Theoretical Computer Science, 164(1–2):299–310, September 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=164&issue=1-2&aid=2255.

Anonymous:1996:CEBb

[440] Anonymous. Contents EATCS Bulletin Number 59. Theoretical Computer Science, 164(1–2):311–??, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bidoit:1996:BTP

[441] Michel Bidoit and Rolf Hennicker. Behavioural theories and the proof of behavioural properties. Theoretical Computer Science, 165(1):3–55, September 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=165&issue=1&aid=2242.

Codish:1996:PIA

[442] Michael Codish and Grigory Mashevitzky. Proving implications by algebraic approximation. Theoretical Computer Science, 165(1):57–74, September 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=165&issue=1&aid=2243.

Antoy:1996:SRS

[443] Sergio Antoy and Aart Middeldorp. A sequential reduction strategy. Theoretical Computer Science, 165(1):75–95, September 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=1&aid=2244.

Gramlich:1996:TCP

[444] Bernhard Gramlich. On termination and confluence properties of disjoint and constructor-sharing conditional rewrite systems. Theoretical Computer Science, 165(1):97–131, September 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=1&aid=2245.

Alpuente:1996:CSB

[445] María Alpuente, Moreno Falaschi, and Germán Vidal. A compositional semantic basis for the analysis of equational Horn programs. Theoretical Computer Science, 165(1):133–169, September 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=1&aid=2246.

Teusink:1996:TVC

[446] Frank Teusink. Three-valued completion for abductive logic programs. Theoretical Computer Science, 165(1):171–200, September 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=165&issue=1&aid=2247.

Miller:1996:FMC

[447] Dale Miller. Forum: A multipleconclusion specification logic. Theoretical Computer Science, 165(1):201– 232, September 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=165&issue=
1&aid=2248.

Fraser:1996:CSS

[448] Campbell Fraser. Consistent subsequences and supersequences. Theoretical Computer Science, 165(2):233– 246, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=165&issue=
2&aid=2079.

Maurin:1996:ECB

[449] Françoise Maurin. Exact complexity bounds for ordinal addition. Theoretical Computer Science, 165(2):247–273, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=165&issue=2&aid=2162.

Pelecq:1996:AGC

[450] Laurent Pélecq. Automorphism groups of context-free graphs. Theoretical Computer Science, 165(2):275– 293, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=165&issue=
                                        2&aid=2163.

Berthe:1996:FFS

[451] Valérie Berthé. Frequences des facteurs des suites sturmiennes. (French) [Frequencies of factors of Sturmian series]. Theoretical Computer Science, 165(2):295–309, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=2&aid=2165.

Munro:1996:SRO

[452] J. Ian Munro and Venkatesh Raman. Selection from read-only memory and sorting with minimum data movement. Theoretical Computer Science, 165(2):311–323, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=2&aid=2166.

Morcrette:1996:EDF

[453] Michelle Morcrette. Sur l’équivalence de descriptions de figures iterées. (French) [On the equivalence of descriptions of iterative figures]. Theoretical Computer Science, 165(2):325– 354, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=165&issue=
2&aid=2167.

Kulkarni:1996:LTP

[454] Sulekha R. Kulkarni and Priti Shankar. Linear time parsers for classes of non context free languages. Theoretical Computer Science, 165(2):355–390, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=165&issue=2&aid=2172.

Habib:1996:TSD

[455] Michel Habib and Lhouari Nourine. Tree structure for distributive lattices and its applications. Theoretical Computer Science, 165(2):391–405, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=165&issue=2&aid=2173.

Blundo:1996:FDS

[456] Carlo Blundo, Antonella Cresti, Alfredo De Santis, and Ugo Vaccaro. Fully dynamic secret sharing schemes. Theoretical Computer Science, 165(2):407–440, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=2&aid=2206.

Auletta:1996:DSA

[457] Vincenzo Auletta, Domenico Parente, and Giuseppe Persiano. Dynamic and static algorithms for optimal placement of resources in a tree. Theoretical Computer Science, 165(2):441–461, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=2&aid=2292.

Dumitrescu:1996:NPG

[458] Sorina Dumitrescu. Nonreturning PC grammar systems can be simulated by returning systems. Theoretical Computer Science, 165(2):463–474, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=165&issue=2&aid=2199.

Imai:1996:FSS

[459] Katsunobu Imai and Kenichi Morita. Firing squad synchronization problem in reversible cellular automata. Theoretical Computer Science, 165(2): 475–482, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=165&issue=
                                        2&aid=2219.

Brimkov:1996:SNC

[460] V. Brimkov, B. Codenotti, M. Leoncini, and G. Resta. Strong NP-completeness of a matrix similarity problem. Theoretical Computer Science, 165(2):483– 490, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=165&issue=
2&aid=2306.

Ramakrishna:1996:ILTa

[461] Y. S. Ramakrishna, P. M. Melliar-Smith, L. E. Moser, L. K. Dillon, and G. Kutty. Interval logics and their decision procedures. part I: An interval logic. Theoretical Computer Science, 166(1–2):1–47, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=166&issue=1-2&aid=2195.

Park:1996:AFA

[462] Uchang Park. An algebraic formulation of the aggregative closure query. Theoretical Computer Science, 166(1–2):49–62, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=166&issue=1-2&aid=2022.

Shen:1996:MPL

[463] Enshao Shen and Qijia Tian. Monadic partition logics and finite automata. Theoretical Computer Science, 166(1– 2):63–81, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=166&issue=
                                        1-2&aid=2054.

Hoofman:1996:CMI

[464] R. Hoofman. Comparing models of the intensional typed λcalculus. Theoretical Computer Science, 166(1–2):83–99, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=166&issue=1-2&aid=2081.

Etalle:1996:TCM

[465] Sandro Etalle and Maurizio Gabbrielli. Transformations of CLP modules. Theoretical Computer Science, 166(1–2):101–146, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=166&issue=1-2&aid=2089.

Fouks:1996:MRA

[466] Jean-Dénis Fouks and Jean-Claude Spehner. Meta-resolution: An algorithmic formalisation. Theoretical Computer Science, 166(1–2):147–172, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=166&issue=1-2&aid=2096.

Demri:1996:LAD

[467] Stéphane Demri and Ewa Orlowska. Logical analysis of demonic nondeterministic programs. Theoretical Computer Science, 166(1–2):173–202, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=166&issue=1-2&aid=2131.

Zhang:1996:LCC

[468] Guo-Qiang Zhang. The largest Cartesian closed category of stable domains. Theoretical Computer Science, 166(1–2):203–219, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=166&issue=1-2&aid=2132.

Gottlob:1996:NGR

[469] Georg Gottlob, Sherry Marcus, Anil Nerode, Gernot Salzer, and V. S. Subrahmanian. A non-ground realization of the stable and well-founded semantics. Theoretical Computer Science, 166(1–2):221–262, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=166&issue=1-2&aid=2148.

Meinke:1996:TMA

[470] Karl Meinke. Topological methods for algebraic specification. Theoretical Computer Science, 166(1–2):263– 290, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

                                        cgi?year=1996&volume=166&issue=
                                        1-2&aid=2202.

Degtyarev:1996:USR

[471] Anatoli Degtyarev and Andrei Voronkov. The undecidability of simultaneous rigid E-unification. Theoretical Computer Science, 166(1–2):291–300, October 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=166&issue=1-2&aid=2295.

Mosses:1996:F

[472] Peter D. Mosses, Mogens Nielsen, and Michael I. Schwartzbach. Foreword. Theoretical Computer Science, 167(1– 2):1, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=167&issue=
                                        1-2&aid=2269.

Hofmann:1996:BAB

[473] Martin Hofmann and Donald Sannella. On behavioural abstraction and behavioural satisfaction in higherorder logic. Theoretical Computer Science, 167(1–2):3–45, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=167&issue=1-2&aid=2270.

Jonsson:1996:AGS

[474] Bengt Jonsson and Yih-Kuen Tsay. Assumption/guarantee specifications in linear-time temporal logic. Theoretical Computer Science, 167(1–2): 47–72, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=167&issue=
1-2&aid=2271.

Kozen:1996:RSS

[475] Dexter Kozen. Rational spaces and set constraints. Theoretical Computer Science, 167(1–2):73–94, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=167&issue=1-2&aid=2272.

Middeldorp:1996:LNS

[476] Aart Middeldorp, Satoshi Okui, and Tetsuo Ida. Lazy narrowing: Strong completeness and eager variable elimination. Theoretical Computer Science, 167(1–2):95–130, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=167&issue=1-2&aid=2273.

Sagiv:1996:PID

[477] Mooly Sagiv, Thomas Reps, and Susan Horwitz. Precise interprocedural dataflow analysis with applications to constant propagation. Theoretical Computer Science, 167(1–2):131–170, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=167&issue=1-2&aid=2274.

Salomaa:1996:DED

[478] Kai Salomaa. Decidability of equivalence for deterministic synchronized tree automata. Theoretical Computer Science, 167(1–2):171–192, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=167&issue=1-2&aid=2275.

Sands:1996:PCR

[479] David Sands. Proving the correctness of recursion-based automatic program transformations. Theoretical Computer Science, 167(1–2):193–233, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=167&issue=1-2&aid=2276.

Sangiorgi:1996:CIM

[480] Davide Sangiorgi. π-calculus, internal mobility, and agent-passing calculi. Theoretical Computer Science, 167(1–2):235–274, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=167&issue=1-2&aid=2277.

Rovan:1996:I

[481] Brànislav Rovan. Introduction. Theoretical Computer Science, 168(1): 1, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=168&issue=
1&aid=2261.

Ambos-Spies:1996:GME

[482] Klaus Ambos-Spies, Hans-Christian Neis, and Sebastiaan A. Terwijn. Genericity and measure for exponential time. Theoretical Computer Science, 168(1):3–19, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=1&aid=2262.

Baeza-Yates:1996:BDE

[483] Ricardo A. Baeza-Yates. Bounded disorder: The effect of the index. Theoretical Computer Science, 168(1):21–38, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=1&aid=2263.

Dietzfelbinger:1996:CTL

[484] Martin Dietzfelbinger, Juraj Hromkovič, and Georg Schnitger. A comparison of two lower-bound methods for communication complexity. Theoretical Computer Science, 168(1):39–51, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=1&aid=2264.

Ferrari:1996:CES

[485] Gian-Luigi Ferrari, Ugo Montanari, and Paola Quaglia. A π-calculus with explicit substitutions. Theoretical Computer Science, 168(1):53–103, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=1&aid=2265.

Karhumaki:1996:SIS

[486] Juhani Karhumäki and Wojciech Plandowski. On the size of independent systems of equations in semigroups. Theoretical Computer Science, 168(1):105–119, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=1&aid=2266.

Kavvadias:1996:HED

[487] Dimitris J. Kavvadias, Grammati E. Pantziou, Paul G. Spirakis, and Christos D. Zaroliagis. Hammock-on-ears decomposition: A technique for the efficient parallel solution of shortest paths and other problems. Theoretical Computer Science, 168(1):121–154, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=1&aid=2267.

Sieber:1996:FAS

[488] Kurt Sieber. Full abstraction for the second order subset of an ALGOLlike language. Theoretical Computer Science, 168(1):155–212, November 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=1&aid=2268.

Anonymous:1996:IWU

[489] Anonymous. International Workshop on Universal Machines and Computations. Theoretical Computer Science, 168(2):??, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Margenstern:1996:F

[490] Maurice Margenstern. Foreword. Theoretical Computer Science, 168(2):213– 214, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=168&issue=
2&aid=2278.

Rogozhin:1996:SUT

[491] Yurii Rogozhin. Small universal Turing machines. Theoretical Computer Science, 168(2):215–240, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=2&aid=2279.

Kudlek:1996:SDT

[492] Manfred Kudlek. Small deterministic Turing machines. Theoretical Computer Science, 168(2):241–255, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=2&aid=2280.

Pavlotskaya:1996:MUE

[493] Liudmila Pavlotskaya. On machines, universal by extensions. Theoretical Computer Science, 168(2):257– 266, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=168&issue=
                                        2&aid=2281.

Korec:1996:SUR

[494] Ivan Korec. Small universal register machines. Theoretical Computer Science, 168(2):267–301, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=2&aid=2282.

Morita:1996:URT

[495] Kenichi Morita. Universality of a reversible two-counter machine. Theoretical Computer Science, 168(2):303– 320, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=168&issue=
                                        2&aid=2283.

Paun:1996:CS

[496] Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Computing by splicing. Theoretical Computer Science, 168(2):321–336, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=2&aid=2284.

Morita:1996:SRR

[497] Kenichi Morita and Katsunobu Imai. Self-reproduction in a reversible cellular space. Theoretical Computer Science, 168(2):337–366, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=2&aid=2285.

Mazoyer:1996:OSF

[498] Jacques Mazoyer. On optimal solutions to the firing squad synchronization problem. Theoretical Computer Science, 168(2):367–404, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=168&issue=2&aid=2286.

Goles:1996:SDU

[499] Eric Goles and Martín Matamala. Symmetric discrete universal neural networks. Theoretical Computer Science, 168(2):405–416, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=2&aid=2287.

Bournez:1996:CPD

[500] Olivier Bournez and Michel Cosnard. On the computational power of dynamical systems and hybrid systems. Theoretical Computer Science, 168(2):417–459, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=2&aid=2288.

Siegelmann:1996:SDS

[501] Hava T. Siegelmann. The simple dynamics of super Turing theories. Theoretical Computer Science, 168(2):461–472, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=2&aid=2289.

Koiran:1996:FUR

[502] Pascal Koiran. A family of universal recurrent networks. Theoretical Computer Science, 168(2):473–480, November 20, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=168&issue=2&aid=2290.

Jay:1996:P

[503] Barry Jay and John Staples. Preface. Theoretical Computer Science, 169 (1):1, November 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=169&issue=
1&aid=2313.

Bunder:1996:LTD

[504] M. W. Bunder. Lambda terms definable as combinators. Theoretical Computer Science, 169(1):3–21, November 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=169&issue=1&aid=2314.

Eades:1996:LER

[505] Peter Eades and Sue Whitesides. The logic engine and the realization problem for nearest neighbor graphs. Theoretical Computer Science, 169(1):23–37, November 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=169&issue=1&aid=2315.

Farrell:1996:FSP

[506] C. A. Farrell and D. H. Kieronska. Formal specification of parallel SIMD execution. Theoretical Computer Science, 169(1):39–65, November 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=169&issue=1&aid=2316.

Gibbons:1996:CDA

[507] Jeremy Gibbons. Computing downwards accumulations on trees quickly. Theoretical Computer Science, 169(1): 67–80, November 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=169&issue=
                                        1&aid=2317.

Nickolas:1996:QPU

[508] Peter Nickolas and Peter J. Robinson. The Qu-Prolog unification algorithm: Formalisation and correctness. Theoretical Computer Science, 169(1):81–112, November 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=169&issue=1&aid=2318.

Yarroll:1996:SMT

[509] La Monte H. Yarroll. See more through lenses than bananas [recursive data structures]. Theoretical Computer Science, 169(1):113–121, November 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=169&issue=1&aid=2319.

Muscholl:1996:CAC

[510] Anca Muscholl. On the complementation of asynchronous cellular Buchi automata. Theoretical Computer Science, 169(2):123–145, December 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=169&issue=2&aid=2320.

Feige:1996:FRL

[511] Uriel Feige. A fast randomized LOGSPACE algorithm for graph connectivity. Theoretical Computer Science, 169(2):147–160, December 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=169&issue=2&aid=2321.

Globerman:1996:CRT

[512] Noa Globerman and David Harel. Complexity results for two-way and multi-pebble automata and their logics. Theoretical Computer Science, 169(2):161–184, December 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=169&issue=2&aid=2322.

Pin:1996:PCG

[513] Jean-Eric Pin. Polynomial closure of group languages and open sets of the Hall topology. Theoretical Computer Science, 169(2):185–200, December 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=169&issue=2&aid=2323.

DiCosmo:1996:CAR

[514] Roberto Di Cosmo and Delia Kesner. Combining algebraic rewriting, extensional lambda calculi, and fixpoints. Theoretical Computer Science, 169(2):201–220, December 05, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=169&issue=2&aid=2324.

Ramakrishna:1996:ILTb

[515] Y. S. Ramakrishna, P. M. Melliar-Smith, L. E. Moser, L. K. Dillon, and G. Kutty. Interval logics and their decision procedures. Part II: A realtime interval logic. Theoretical Computer Science, 170(1–2):1–46, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2196.

Groote:1996:CPV

[516] J. F. Groote and M. P. A. Sellink. Confluence for process verification. Theoretical Computer Science, 170(1–2):47–81, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2388.

Dezani-Ciancaglini:1996:FMC

[517] Mariangiola Dezani-Ciancaglini, Ugo de’Liguoro, and Adolfo Piperno. Filter models for conjunctive-disjunctive λ-calculi. Theoretical Computer Science, 170(1–2):83–128, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=170&issue=1-2&aid=2176.

Arai:1996:TCF

[518] Noriko H. Arai. Tractability of cut-free Gentzen type propositional calculus with permutation inference. Theoretical Computer Science, 170(1–2):129–144, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=170&issue=1-2&aid=2182.

Majster-Cederbaum:1996:MCV

[519] Mila E. Majster-Cederbaum and Christel Baier. Metric completion versus ideal completion. Theoretical Computer Science, 170(1–2):145– 171, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=170&issue=
1-2&aid=2203.

Barbanera:1996:ITA

[520] Franco Barbanera and Maribel Fernández. Intersection type assignment systems with higher-order algebraic rewriting. Theoretical Computer Science, 170(1–2):173–207, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2204.

Dimopoulos:1996:GTS

[521] Yannis Dimopoulos and Alberto Torres. Graph theoretical structures in logic programs and default theories. Theoretical Computer Science, 170(1–2):209–244, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2207.

Bouhoula:1996:UIR

[522] Adel Bouhoula. Using induction and rewriting to verify and complete parameterized specifications. Theoretical Computer Science, 170(1–2):245–276, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2208.

Dybjer:1996:RID

[523] Peter Dybjer. Representing inductively defined sets by wellorderings in Martin-Löf’s type theory. Theoretical Computer Science, 170(1–2):245– 276, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1996&volume=170&issue=
                                        1-2&aid=2348.

Sassone:1996:AAP

[524] Vladimiro Sassone. An axiomatization of the algebra of Petri net concatenable processes. Theoretical Computer Science, 170(1–2):277–296, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=170&issue=1-2&aid=2212.

Sassone:1996:MCT

[525] Vladimiro Sassone, Mogens Nielsen, and Glynn Winskel. Models for concurrency: Towards a classification. Theoretical Computer Science, 170(1–2):297–348, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=170&issue=1-2&aid=2214.

Rutten:1996:EGU

[526] J. J. M. M. Rutten. Elements of generalized ultrametric domain theory. Theoretical Computer Science, 170(1–2):349–381, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1996&volume=170&issue=1-2&aid=2216.

You:1996:IBR

[527] Jia-Huai You, Robert Cartwright, and Ming Li. Iterative belief revision in extended logic programming. Theoretical Computer Science, 170(1–2):383–406, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2293.

Hilken:1996:TPT

[528] Barnaby P. Hilken. Towards a proof theory of rewriting: The simply typed 2λ-calculus. Theoretical Computer Science, 170(1–2):407–444, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2300.

Yen:1996:DBT

[529] Hsu-Chun Yen, Shi-Tsuen Jian, and Ta-Pang Lao. Deciding bisimulation and trace equivalences for systems with many identical processes. Theoretical Computer Science, 170(1–2):445–464, December 15, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1996&volume=170&issue=1-2&aid=2389.

Anonymous:1996:MIVb

[530] Anonymous. Master Index Volume 161–170. Theoretical Computer Science, 170(1–2):467–??, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lakshmanan:1997:P

[531] Laks V. S. Lakshmanan. Preface. Theoretical Computer Science, 171(1–2):1–2, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2325.

Ausiello:1997:LAS

[532] Giorgio Ausiello and Roberto Giaccio. On-line algorithms for satisfiability problems with uncertainty. Theoretical Computer Science, 171(1–2):3–24, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2326.

Koubarakis:1997:CQE

[533] Manolis Koubarakis. The complexity of query evaluation in indefinite temporal constraint databases. Theoretical Computer Science, 171(1–2):25–60, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2327.

Ruet:1997:CEN

[534] Paul Ruet and François Fages. Combining explicit negation and negation by failure via Belnap’s logic. Theoretical Computer Science, 171(1–2):61–75, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2328.

Mobasher:1997:MVL

[535] Bamshad Mobasher, Don Pigozzi, and Giora Slutzki. Multi-valued logic programming semantics: An algebraic approach. Theoretical Computer Science, 171(1–2):77–109, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=171&issue=1-2&aid=2329.

Nerode:1997:ANR

[536] A. Nerode, J. B. Remmel, and V. S. Subrahmanian. Annotated nonmonotonic rule systems. Theoretical Computer Science, 171(1–2):111–146, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=171&issue=1-2&aid=2330.

Ngo:1997:AQC

[537] Liem Ngo and Peter Haddawy. Answering queries from context-sensitive probabilistic knowledge bases. Theoretical Computer Science, 171(1–2): 147–177, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=171&issue=
                                        1-2&aid=2331.

Zimanyi:1997:QEP

[538] Esteban Zimányi. Query evaluation in probabilistic relational databases. Theoretical Computer Science, 171(1–2):179–219, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2332.

Kohlas:1997:AAE

[539] Jürg Kohlas. Allocation of arguments and evidence theory. Theoretical Computer Science, 171(1–2):221–246, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2333.

Chatalic:1997:GHT

[540] Philippe Chatalic, Christine Froidevaux, and Camilla Schwind. Graded hypothesis theories. Theoretical Computer Science, 171(1–2):247–280, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2334.

Bosc:1997:FQR

[541] Patrick Bosc, Didier Dubois, Olivier Pivert, and Henri Prade. Flexible queries in relational databases — The example of the division operator. Theoretical Computer Science, 171(1–2):281–302, January 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=171&issue=1-2&aid=2335.

Anonymous:1997:CEBa

[542] Anonymous. Contents EATCS Bulletin Number 60. Theoretical Computer Science, 171(1–2):303–??, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Edmonds:1997:FSR

[543] Jeff Edmonds. Fundamental study: Removing Ramsey theory: lower bounds with smaller domain size. Theoretical Computer Science, 172(1–2): 1–41, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=172&issue=
                                        1-2&aid=2223.

Fischer:1997:PLG

[544] Paul Fischer, Klaus-Uwe Höffgen, and Hanno Lefmann. PAC-learning from general examples. Theoretical Computer Science, 172(1–2):43–65, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=172&issue=1-2&aid=2177.

Fagnot:1997:FMA

[545] Isabelle Fagnot. Sur les facteurs des mots automatiques. (French) [On the factors of automatic words]. Theoretical Computer Science, 172(1–2): 67–89, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.

cgi?year=1997&volume=172&issue=
1-2&aid=2180.

Apolloni:1997:PLC

[546] B. Apolloni and S. Chiaravalli. PAC learning of concept classes through the boundaries of their items. Theoretical Computer Science, 172(1–2):91–120, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2181.

Goles:1997:UCF

[547] Eric Goles and Maurice Margenstern. Universality of the chip-firing game. Theoretical Computer Science, 172(1–2):121–134, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2183.

Handley:1997:DSM

[548] W. G. Handley. Deterministic summation modulo B_{n}, the semigroup of binary relations on 0, 1, …, n − 1. Theoretical Computer Science, 172(1–2):135–174, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2186.

Kant:1997:REL

[549] Goos Kant and Xin He. Regular edge labeling of 4-connected plane graphs and its applications in graph drawing problems. Theoretical Computer Science, 172(1–2):175–193, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2198.

Ambos-Spies:1997:RBR

[550] Klaus Ambos-Spies, Sebastiaan A. Terwijn, and Xizhong Zheng. Resource bounded randomness and weakly complete problems. Theoretical Computer Science, 172(1–2):195–207, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2201.

Brandstadt:1997:HOG

[551] Andreas Brandstädt, Feodor F. Dragan, and Falk Nicolai. Homogeneously orderable graphs. Theoretical Computer Science, 172(1–2):209– 232, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=172&issue=
1-2&aid=2294.

Dendris:1997:FSG

[552] Nick D. Dendris, Lefteris M. Kirousis, and Dimitrios M. Thilikos. Fugitivesearch games on graphs and related parameters. Theoretical Computer Science, 172(1–2):233–254, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2390.

Hohl:1997:MPF

[553] Jean-Christophe Hohl. Massively parallel factorizations of polynomials with many non-commuting variables. Theoretical Computer Science, 172(1–2):255–263, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2311.

Shimozono:1997:FOS

[554] Shinichi Shimozono. Finding optimal subgraphs by local search. Theoretical Computer Science, 172(1–2): 265–271, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=172&issue=
1-2&aid=2338.

Rystsov:1997:RWC

[555] Igor Rystsov. Reset words for commutative and solvable automata. Theoretical Computer Science, 172(1–2): 273–279, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=172&issue=
1-2&aid=2339.

Iliopoulos:1997:CSF

[556] Costas S. Iliopoulos, Dennis Moore, and W. F. Smyth. A characterization of the squares in a Fibonacci string. Theoretical Computer Science, 172(1–2):281–291, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=172&issue=1-2&aid=2344.

Savicky:1997:LBB

[557] Petr Savický and Stanislav Žák. A lower bound on branching programs reading some bits twice. Theoretical Computer Science, 172(1–2):293– 301, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=172&issue=
                                        1-2&aid=2394.

Rote:1997:FSV

[558] Günter Rote. Finding a shortest vector in a two-dimensional lattice modulo m. Theoretical Computer Science, 172(1–2):303–308, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=172&issue=1-2&aid=2397.

Hoogeboom:1997:CPT

[559] Hendrik Jan Hoogeboom and Anca Muscholl. The code problem for traces — improving the boundaries. Theoretical Computer Science, 172(1–2):309–321, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=172&issue=1-2&aid=2419.

Montanari:1997:P

[560] Ugo Montanari and Francesca Rossi. Preface. Theoretical Computer Science, 173(1):1, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=1&aid=2370.

Michel:1997:HML

[561] Laurent Michel and Pascal Van Hentenryck. Helios: A modeling language for global optimization and its implementation in Newton. Theoretical Computer Science, 173(1):3–48, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=1&aid=2361.

Bjorner:1997:AGI

[562] Nikolaj Bjørner, Anca Browne, and Zohar Manna. Automatic generation of invariants and intermediate assertions. Theoretical Computer Science, 173(1):49–87, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=1&aid=2362.

Koubarakis:1997:LGC

[563] Manolis Koubarakis. From local to global consistency in temporal constraint networks. Theoretical Computer Science, 173(1):89–112, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=173&issue=1&aid=2363.

Maher:1997:CD

[564] Michael J. Maher. Constrained dependencies. Theoretical Computer Science, 173(1):113–149, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=173&issue=1&aid=2364.

Grumbach:1997:QAC

[565] Stéphane Grumbach and Jianwen Su. Queries with arithmetical constraints. Theoretical Computer Science, 173(1):151–181, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=173&issue=1&aid=2365.

Ajili:1997:ASV

[566] Farid Ajili and Evelyne Contejean. Avoiding slack variables in the solving of linear Diophantine equations and inequations. Theoretical Computer Science, 173(1):183–208, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=1&aid=2366.

Marriott:1997:CCC

[567] Kim Marriott and Martin Odersky. A confluent calculus for concurrent constraint programming. Theoretical Computer Science, 173(1):209–233, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=1&aid=2367.

Podelski:1997:SS

[568] Andreas Podelski and Gert Smolka. Situated simplification. Theoretical Computer Science, 173(1):235– 252, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=173&issue=
1&aid=2368.

Girodias:1997:SLM

[569] Pierre Girodias, Eduard Cerny, and William J. Older. Solving linear, min and max constraint systems using CLP based on relational interval arithmetic. Theoretical Computer Science, 173(1):253–281, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=1&aid=2369.

Dechter:1997:LGR

[570] Rina Dechter and Peter van Beek. Local and global relational consistency. Theoretical Computer Science, 173(1):283–308, February 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=173&issue=1&aid=2371.

Astesiano:1997:P

[571] Egidio Astesiano. Preface. Theoretical Computer Science, 173(2):309– 310, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=173&issue=
                                        2&aid=2372.

Cerioli:1997:MBY

[572] Maura Cerioli and José Meseguer. May I borrow your logic? (transporting logical structures along maps). Theoretical Computer Science, 173(2):311– 347, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=173&issue=
                                        2&aid=2373.

Jouannaud:1997:ADT

[573] Jean-Pierre Jouannaud and Mitsuhiro Okada. Abstract data type systems. Theoretical Computer Science, 173(2):349–391, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=2&aid=2374.

Hennicker:1997:PSS

[574] Rolf Hennicker, Martin Wirsing, and Michel Bidoit. Proof systems for structured specifications with observability operators. Theoretical Computer Science, 173(2):393–443, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=2&aid=2375.

Kahrs:1997:DEM

[575] Stefan Kahrs, Donald Sannella, and Andrzej Tarlecki. The definition of Extended ML: A gentle introduction. Theoretical Computer Science, 173(2):445–484, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=2&aid=2376.

Orejas:1997:ILP

[576] Fernando Orejas, Elvira Pino, and Hartmut Ehrig. Institutions for logic programming. Theoretical Computer Science, 173(2):485–511, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=173&issue=2&aid=2377.

Costa:1997:SAD

[577] Gerardo Costa and Gianna Reggio. Specification of abstract dynamicdata types: A temporal logic approach. Theoretical Computer Science, 173(2):513–554, February 28, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=173&issue=2&aid=2378.

Litovsky:1997:FAI

[578] Igor Litovsky and Ludwig Staiger. Finite acceptance of infinite words. Theoretical Computer Science, 174(1– 2):1–21, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=174&issue=
                                        1-2&aid=2209.

Marathe:1997:HSU

[579] Madhav V. Marathe, Venkatesh Radhakrishnan, Harry B. Hunt III, and S. S. Ravi. Hierarchically specified unit disk graphs. Theoretical Computer Science, 174(1–2):23–65, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=174&issue=1-2&aid=2211.

Bracho:1997:RCC

[580] Felipe Bracho, Manfred Droste, and Dietrich Kuske. Representation of computations in concurrent automata by dependence orders. Theoretical Computer Science, 174(1–2):67–96, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=174&issue=1-2&aid=2215.

Arora:1997:WSE

[581] Sanjeev Arora and Ronald Fagin. On winning strategies in Ehrenfeucht– Fraïssé; games. Theoretical Computer Science, 174(1–2):97–121, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=174&issue=1-2&aid=2218.

Orponen:1997:CTA

[582] Pekka Orponen. Computing with truly asynchronous threshold logic networks. Theoretical Computer Science, 174(1–2):123–136, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=174&issue=1-2&aid=2220.

Krause:1997:CPD

[583] Matthias Krause and Pavel Pudlák. On the computational power of depth-2 circuits with threshold and modulo gates. Theoretical Computer Science, 174(1–2):137–156, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=174&issue=1-2&aid=2222.

Favati:1997:AOD

[584] Paola Favati, Grazia Lotti, and Luciano Margara. Additive onedimensional cellular automata are chaotic according to Devaney’s definition of chaos. Theoretical Computer Science, 174(1–2):157–170, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=174&issue=1-2&aid=2225.

Santini-Bouchard:1997:ETI

[585] Marie-Line Santini-Bouchard. Echanges de trois intervalles et suites minimales. (French) [Exchanges of three intervals and minimal sequences]. Theoretical Computer Science, 174(1–2): 171–191, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=174&issue=
                                        1-2&aid=2226.

Berman:1997:NOP

[586] Piotr Berman and Andrzej Lingas. A nearly optimal parallel algorithm for the Voronoi diagram of a convex polygon. Theoretical Computer Science, 174(1–2):193–202, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=174&issue=1-2&aid=2227.

Kurka:1997:TDT

[587] Petr Kůrka. On topological dynamics of Turing machines. Theoretical Computer Science, 174(1–2): 203–216, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=174&issue=

1-2&aid=2228. See erratum [1427].

Finkel:1997:VIC

[588] Alain Finkel and Pierre McKenzie. Verifying identical communicating processes is undecidable. Theoretical Computer Science, 174(1–2): 217–230, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=174&issue=
1-2&aid=2229.

Mullins:1997:EHC

[589] John Mullins. On an effective hierarchy of communicating processes: Separation principle and testing. Theoretical Computer Science, 174(1–2):231–246, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=174&issue=1-2&aid=2230.

Yamasaki:1997:MGA

[590] Yõhei Yamasaki. Mathematical games. the arithmetic of reversed positional games. Theoretical Computer Science, 174(1–2):247–249, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=174&issue=1-2&aid=2359.

Kobayashi:1997:LAR

[591] Satoshi Kobayashi and Takashi Yokomori. Learning approximately regular languages with reversible languages. Theoretical Computer Science, 174(1–2):251–257, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=174&issue=1-2&aid=2427.

Takahashi:1997:VHR

[592] Nobuyuki Takahashi. Various hierarchies of ω-regular sets. Theoretical Computer Science, 174(1–2): 259–268, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=174&issue=
                                        1-2&aid=2447.

Turakainen:1997:USE

[593] P. Turakainen. The undecidability of some equivalence problems concerning ngsm’s and finite substitutions. Theoretical Computer Science, 174(1–2):269–274, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=174&issue=1-2&aid=2448.

Zhang:1997:EP

[594] Guo-Qiang Zhang and Rodney E. Canfield. The end of pumping? Theoretical Computer Science, 174(1–2):275–279, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=174&issue=1-2&aid=2449.

Anonymous:1997:WNS

[595] Anonymous. Workshop on Non-Standard Logics and Logical Aspects of Computer Science. Theoretical Computer Science, 175(1):??, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ono:1997:F

[596] Hiroakira Ono. Foreword. Theoretical Computer Science, 175(1):1, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=1&aid=2379.

Ershov:1997:BCH

[597] Yu. L. Ershov. The bounded-complete hull of an α-space. Theoretical Computer Science, 175(1):3–13, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=1&aid=2380.

Shilov:1997:PSV

[598] N. V. Shilov. Program schemata vs. automata for decidability of program logics. Theoretical Computer Science, 175(1):15–27, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=1&aid=2381.

Kobayashi:1997:MM

[599] Satoshi Kobayashi. Monad as modality. Theoretical Computer Science, 175(1):29–74, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=1&aid=2382.

Sato:1997:ICN

[600] Masahiko Sato. Intuitionistic and classical natural deduction systems with the catch and the throw rules. Theoretical Computer Science, 175(1): 75–92, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=175&issue=
1&aid=2383.

Kennaway:1997:ILC

[601] J. R. Kennaway, J. W. Klop, M. R. Sleep, and F. J. de Vries. Infinitary lambda calculus. Theoretical Computer Science, 175(1):93–125, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=1&aid=2384.

Middeldorp:1997:STR

[602] Aart Middeldorp and Hans Zantema. Simple termination of rewrite systems. Theoretical Computer Science, 175(1):127–158, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=1&aid=2385.

vanOostrom:1997:DDR

[603] Vincent van Oostrom. Developing developments [rewriting system confluence]. Theoretical Computer Science, 175(1):159–181, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=1&aid=2386.

Lisitsa:1997:LSL

[604] Alexei Lisitsa and Vladimir Sazonov. Δ-languages for sets and LOGSPACE computable graph transformers. Theoretical Computer Science, 175(1): 183–222, March 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=175&issue=
1&aid=2387.

Bouchitte:1997:P

[605] Vincent Bouchitté, Michel Habib, and Michel Morvan. Preface. Theoretical Computer Science, 175(2):223, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=175&issue=2&aid=2401.

Bouchitte:1997:LAO

[606] Vincent Bouchitté and Jean-Xavier Rampon. On-line algorithms for orders. Theoretical Computer Science, 175(2):225–238, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=175&issue=2&aid=2402.

Bauernoppel:1997:PSG

[607] Frank Bauernöppel, Evangelos Kranakis, Danny Krizanc, Anil Maheshwari, Jörg-Rüdiger Sack, and Jorge Urrutia. Planar stage graphs: Characterizations and applications. Theoretical Computer Science, 175(2):239–255, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=175&issue=2&aid=2403.

Ekin:1997:HFS

[608] Oya Ekin, Peter L. Hammer, and Uri N. Peled. Horn functions and submodular Boolean functions. Theoretical Computer Science, 175(2): 257–270, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=175&issue=
2&aid=2404.

Ewacha:1997:ANL

[609] Kevin Ewacha, Ivan Rival, and Nejib Zaguia. Approximating the number of linear extensions. Theoretical Computer Science, 175(2):271–282, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=2&aid=2405.

Felsner:1997:LCP

[610] Stefan Felsner. On-line chain partitions of orders. Theoretical Computer Science, 175(2):283–292, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=2&aid=2406.

delaHiguera:1997:DET

[611] Colin de la Higuera and Lhouari Nourine. Drawing and encoding twodimensional posets. Theoretical Computer Science, 175(2):293–308, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=2&aid=2407.

Kloks:1997:TMF

[612] Ton Kloks, Dieter Kratsch, and Jeremy Spinrad. On treewidth and minimum fill-in of asteroidal triplefree graphs. Theoretical Computer Science, 175(2):309–335, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=175&issue=2&aid=2408.

Mitas:1997:CPE

[613] Jutta Mitas and Klaus Reuter. Cover-preserving embeddings of bipartite orders into Boolean lattices. Theoretical Computer Science, 175(2):337–347, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=175&issue=2&aid=2409.

Peer:1997:SPI

[614] Itsik Pe’er and Ron Shamir. Satisfiability problems on intervals and unit intervals. Theoretical Computer Science, 175(2):349–372, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=175&issue=2&aid=2410.

Talamo:1997:DSL

[615] M. Talamo and P. Vocca. A data structure for lattice representation. Theoretical Computer Science, 175(2):373–392, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=2&aid=2411.

Viennot:1997:PFO

[616] Laurent Viennot. Parallel N-free order recognition. Theoretical Computer Science, 175(2):393–406, April 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=175&issue=2&aid=2412.

Wu:1997:CBP

[617] Sue-Hwey Wu, Scott A. Smolka, and Eugene W. Stark. Composition and behaviors of probabilistic I/O automata. Theoretical Computer Science, 176(1–2):1–38, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=176&issue=1-2&aid=2258.

Chiola:1997:SRG

[618] G. Chiola, C. Dutheillet, G. Franceschinis, and S. Haddad. A symbolic reachability graph for coloured Petri nets. Theoretical Computer Science, 176(1–2):39–65, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=176&issue=1-2&aid=2213.

Comon:1997:FOT

[619] Hubert Comon and Ralf Treinen. The first-order theory of lexicographic path orderings is undecidable. Theoretical Computer Science, 176(1–2):67–87, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=176&issue=1-2&aid=2249.

Hoover:1997:LSN

[620] D. N. Hoover. Limiting semantics of numerical programs. Theoretical Computer Science, 176(1–2):89–110, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=176&issue=1-2&aid=2250.

Hermann:1997:UIS

[621] Miki Hermann and Roman Galbavý. Unification of infinite sets of terms schematized by primal grammars. Theoretical Computer Science, 176(1–2):111–158, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=176&issue=1-2&aid=2252.

Martini:1997:ELN

[622] Simone Martini and Andrea Masini. Experiments in linear natural deduction. Theoretical Computer Science, 176(1–2):159–173, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=176&issue=1-2&aid=2257.

Matos:1997:MLP

[623] Armando B. Matos. Monadic logic programs and functional complexity. Theoretical Computer Science, 176(1–2):175–204, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=176&issue=1-2&aid=2296.

Hartonas:1997:SFD

[624] Chrysafis Hartonas. Semantics for finite delay. Theoretical Computer Science, 176(1–2):205–234, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=176&issue=1-2&aid=2298.

Pierce:1997:HOS

[625] Benjamin Pierce and Martin Steffen. Higher-order subtyping. Theoretical Computer Science, 176(1–2): 235–282, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=176&issue=

1-2&aid=2299. See corrigendum [761].

Suciu:1997:BFC

[626] Dan Suciu. Bounded fixpoints for complex objects. Theoretical Computer Science, 176(1–2):283–328, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=176&issue=1-2&aid=2530.

Dybjer:1997:RID

[627] P. Dybjer. Representing inductively defined sets by wellorderings in Martin-Löf’s type theory. Theoretical Computer Science, 176(1–2):329–335, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castagna:1997:UOA

[628] Giuseppe Castagna. Unifying overloading and λ-abstraction: λ^{{}}. Theoretical Computer Science, 176(1–2): 337–345, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=176&issue=
                                        1-2&aid=2424.

Massol:1997:MSS

[629] A. Massol. Minimality of the system of seven equations for the category of finite sets. Theoretical Computer Science, 176(1–2):347–353, April 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=176&issue=1-2&aid=2444.

Mislove:1997:F

[630] Michael W. Mislove and David A. Schmidt. Foreword. Theoretical Computer Science, 177(1):1, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=1&aid=2433.

Anderson:1997:RAF

[631] S. O. Anderson and A. J. Power. A representable approach to finite nondeterminism. Theoretical Computer Science, 177(1):3–25, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=1&aid=2434.

Brauner:1997:GAR

[632] Torben Braüner. A general adequacy result for a linear functional language. Theoretical Computer Science, 177(1):27–58, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=1&aid=2435.

Bucciarelli:1997:DPC

[633] Antonio Bucciarelli. Degrees of parallelism in the continuous type hierarchy. Theoretical Computer Science, 177(1):59–71, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=1&aid=2436.

Cockett:1997:CPC

[634] J. R. B. Cockett and David A. Spooner. Constructing process categories. Theoretical Computer Science, 177(1):73–109, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=177&issue=1&aid=2437.

Flagg:1997:CSR

[635] Bob Flagg and Ralph Kopperman. Continuity spaces: Reconciling domains and metric spaces. Theoretical Computer Science, 177(1): 111–138, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=177&issue=
                                        1&aid=2438.

Nolte:1997:AFS

[636] Doris Nolte and Lutz Priese. Abstract fairness and semantics. Theoretical Computer Science, 177(1): 139–153, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=177&issue=
                                        1&aid=2439.

Zhang:1997:DDT

[637] Guo-Qiang Zhang and William C. Rounds. Defaults in domain theory. Theoretical Computer Science, 177(1):155–182, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=1&aid=2440.

Leavens:1997:BRA

[638] Gray T. Leavens and Don Pigozzi. The behavior–realization adjunction and generalized homomorphic relations. Theoretical Computer Science, 177(1):183–216, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=1&aid=2441.

Esik:1997:CPI

[639] Z. Ésik. Completeness of Park induction. Theoretical Computer Science, 177(1):217–283, April 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=1&aid=2442.

Ponse:1997:P

[640] Alban Ponse, Chris Verhoef, and Bas van Vlijmen. Preface. Theoretical Computer Science, 177(2):285–286, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=2&aid=2451.

Vrancken:1997:ACP

[641] J. L. M. Vrancken. The algebra of communicating processes with empty process. Theoretical Computer Science, 177(2):287–328, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=2&aid=2452.

vanGlabbeek:1997:NMC

[642] R. J. van Glabbeek. Notes on the methodology of CCS and CSP. Theoretical Computer Science, 177(2): 329–349, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=177&issue=
2&aid=2453.

DArgenio:1997:GCE

[643] Pedro R. D’Argenio and Chris Verhoef. A general conservative extension theorem in process algebras with inequalities. Theoretical Computer Science, 177(2):351–380, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=2&aid=2506.

Baeten:1997:PAP

[644] J. C. M. Baeten and J. A. Bergstra. Process algebra with propositional signals. Theoretical Computer Science, 177(2):381–405, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/

browse.cgi?year=1997&volume=177&
issue=2&aid=2455.

Fokkink:1997:TME

[645] Wan Fokkink and Hans Zantema. Termination module equations by abstract commutation with an application to iteration. Theoretical Computer Science, 177(2):407–423, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=2&aid=2456.

vanWamel:1997:PAL

[646] Jos van Wamel. Process algebra with language matching. Theoretical Computer Science, 177(2):425–458, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=2&aid=2457.

Fredlund:1997:FVL

[647] Lars åke Fredlund, Jan Friso Groote, and Henri Korver. Formal verification of a leader election protocol in process algebra. Theoretical Computer Science, 177(2):459–486, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=2&aid=2458.

Bezem:1997:TFS

[648] Marc Bezem and Alban Ponse. Two finite specifications of a queue. Theoretical Computer Science, 177(2):487–507, May 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=177&issue=2&aid=2459.

Chang:1997:RED

[649] Chia-Hsiang Chang and Robert Paige. From regular expressions to DFA’s using compressed NFA’s. Theoretical Computer Science, 178(1–2):1–36, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2343.

Clote:1997:NSR

[650] P. Clote. Nondeterministic stack register machines. Theoretical Computer Science, 178(1–2):37–76, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2251.

Cattaneo:1997:MRS

[651] G. Cattaneo and C. Quaranta Vogliotti. The “magic” rule spaces of neural-like elementary cellular automata. Theoretical Computer Science, 178(1–2): 77–102, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=178&issue=
1-2&aid=2253.

Atkinson:1997:PGT

[652] M. D. Atkinson, M. J. Livesey, and D. Tulley. Permutations generated by token passing in graphs. Theoretical Computer Science, 178(1–2):103–118, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2259.

Munier:1997:UDS

[653] A. Munier and C. Hanen. Using duplication for scheduling unitary tasks on m processors with unit communication delays. Theoretical Computer Science, 178(1–2):119–127, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2260.

Kucherov:1997:MSS

[654] Gregory Kucherov and Michaël Rusinowitch. Matching a set of strings with variable length don’t cares. Theoretical Computer Science, 178(1–2):129–154, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2297.

Strauss:1997:NNS

[655] Martin Strauss. Normal numbers and sources for BPP. Theoretical Computer Science, 178(1–2):155–169, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=178&issue=1-2&aid=2302.

Berstel:1997:SWL

[656] Jean Berstel and Aldo de Luca. Sturmian words, Lyndon words and trees. Theoretical Computer Science, 178(1–2):171–203, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=178&issue=1-2&aid=2304.

deLuca:1997:SSM

[657] Aldo de Luca. Standard Sturmian morphisms. Theoretical Computer Science, 178(1–2):205–224, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=178&issue=1-2&aid=2256.

Damaschke:1997:OPA

[658] Peter Damaschke. An optimal parallel algorithm for digital curve segmentation. Theoretical Computer Science, 178(1–2):225–236, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=178&issue=1-2&aid=2430.

Browne:1997:IAE

[659] A. Browne, E. M. Clarke, S. Jha, D. E. Long, and W. Marrero. An improved algorithm for the evaluation of fixpoint expressions. Theoretical Computer Science, 178(1–2):237–255, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2431.

Higgins:1997:PST

[660] Peter M. Higgins. A proof of Simon’s theorem on piecewise testable languages. Theoretical Computer Science, 178(1–2):257–264, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2413.

Kopidakis:1997:ASS

[661] Y. Kopidakis and V. Zissimopoulos. An approximation scheme for scheduling independent jobs into subcubes of a hypercube of fixed dimension. Theoretical Computer Science, 178(1–2):265–273, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=178&issue=1-2&aid=2484.

Latteux:1997:RPL

[662] Michel Latteux and David Simplot. Recognizable picture languages and domino tiling. Theoretical Computer Science, 178(1–2):275–283, May 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=178&issue=1-2&aid=2485.

Bloom:1997:ELF

[663] Stephen L. Bloom and Zoltán Ésik. The equational logic of fixed points. Theoretical Computer Science, 179(1– 2):1–60, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=179&issue=
                                        1-2&aid=2450.

Keesmaat:1997:RRV

[664] N. W. Keesmaat and H. C. M. Kleijn. Restrictions and representations of vector controlled concurrent system behaviours. Theoretical Computer Science, 179(1–2):61–102, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=179&issue=1-2&aid=2308.

Doornbos:1997:CAM

[665] Henk Doornbos, Roland Backhouse, and Jaap van der Woude. A calculational approach to mathematical induction. Theoretical Computer Science, 179(1–2):103–135, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=179&issue=1-2&aid=2357.

Berline:1997:DSM

[666] Chantal Berline and Klaus Grue. A κ-denotational semantics for map theory in ZFC + SI. Theoretical Computer Science, 179(1–2):137–202, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2502.

Castellani:1997:PPE

[667] Ilaria Castellani and Guo-Qiang Zhang. Parallel product of event structures. Theoretical Computer Science, 179(1–2):203–215, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2307.

Baier:1997:TS

[668] Christel Baier. Trees and semantics. Theoretical Computer Science, 179(1–2):217–250, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2310.

Gruska:1997:SDS

[669] Jozef Gruska, Angelo Monti, Margherita Napoli, and Domenico Parente. Succinctness of descriptions of SBTAlanguages. Theoretical Computer Science, 179(1–2):251–271, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2336.

Alagic:1997:OSM

[670] Suad Alagić and Mara Alagić. Ordersorted model theory for temporal executable specifications. Theoretical Computer Science, 179(1–2):273–299, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=179&issue=1-2&aid=2337.

Yen:1997:REB

[671] Hsu-Chun Yen. On reachability equivalence for BPP-nets. Theoretical Computer Science, 179(1–2):301–317, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=179&issue=1-2&aid=2350.

Sigstam:1997:RLC

[672] Inger Sigstam and Viggo Stoltenberg-Hansen. Representability of locally compact regular spaces by domains and formal spaces. Theoretical Computer Science, 179(1–2):319–331, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=179&issue=1-2&aid=2355.

Lamport:1997:PEB

[673] Leslie Lamport. Processes are in the eye of the beholder. Theoretical Computer Science, 179(1–2):333–351, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2356.

Gurevich:1997:EEB

[674] Yuri Gurevich and James K. Huggins. Equivalence is in the eye of the beholder. Theoretical Computer Science, 179(1–2):353–380, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2516.

Ibarra:1997:PCL

[675] Oscar H. Ibarra, Nicholas Q. Tran, and Tao Yang. On the parallel complexity of loops. Theoretical Computer Science, 179(1–2):381–395, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2396.

Eloranta:1997:ETB

[676] Jaana Eloranta, Martti Tienari, and Antti Valmari. Essential transitions to bisimulation equivalences. Theoretical Computer Science, 179(1–2):397–419, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2483.

Koslowski:1997:NFA

[677] Jürgen Koslowski. Note on free algebras over continuous domains. Theoretical Computer Science, 179(1–2):

421–425, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=179&issue=
1-2&aid=2489.

Duntsch:1997:LRS

[678] Ivo Düntsch. A logic for rough sets. Theoretical Computer Science, 179(1–2):427–436, June 01, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=179&issue=1-2&aid=2536.

Madonia:1997:SDP

[679] Marina Madonia and Stefano Varricchio. Some decisional problems on rational relations. Theoretical Computer Science, 180(1–2):1–15, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2301.

Arvind:1997:SBB

[680] V. Arvind and N. V. Vinodchandran. Solvable black-box group problems are low for PP. Theoretical Computer Science, 180(1–2):17–45, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2303.

Hosaka:1997:SOB

[681] K. Hosaka, Y. Takenaga, T. Kaneda, and S. Yajima. Size of ordered binary decision diagrams representing threshold functions. Theoretical Computer Science, 180(1–2):47–60, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2305.

Bassino:1997:NCM

[682] Frédérique Bassino. Nonnegative companion matrices and star-height of N-rational series. Theoretical Computer Science, 180(1–2):61–80, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2309.

Garel:1997:SMI

[683] E. Garel. Séparateurs dans les mots infinis engendrés par morphismes. (French) [Separators in infinite words generated by morphisms]. Theoretical Computer Science, 180(1–2):81–113, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2312.

Sagot:1997:MSC

[684] Marie-France Sagot, Alain Viari, and Henri Soldano. Multiple sequence comparison — A peptide matching approach. Theoretical Computer Science, 180(1–2):115–137, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2340.

Yamamoto:1997:PAR

[685] Hiroaki Yamamoto. On the power of alternation on reversal-bounded alternating Turing machines with a restriction. Theoretical Computer Science, 180(1–2):139–154, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2346.

Chen:1997:LCF

[686] Zhixiang Chen and Steven Homer. Learning counting functions with queries. Theoretical Computer Science, 180(1–2):155–168, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2347.

Shen:1997:EEA

[687] Hong Shen and Weifa Liang. Efficient enumeration of all minimal separators in a graph. Theoretical Computer Science, 180(1–2):169–180, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2351.

Pomerance:1997:AID

[688] Carl Pomerance, John Michael Robson, and Jeffrey Shallit. Automaticity II: Descriptional complexity in the unary case. Theoretical Computer Science, 180(1–2):181–201, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2360.

Leung:1997:UUM

[689] Vitus J. Leung. The undecidability of the unrestricted modified edit distance. Theoretical Computer Science, 180(1–2):203–215, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2399.

Grigoriev:1997:TSE

[690] D. Grigoriev. Testing shift-equivalence of polynomials by deterministic, probabilistic and quantum machines. Theoretical Computer Science, 180(1–2): 217–228, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=180&issue=
1-2&aid=2400.

Matamala:1997:ACA

[691] Martín Matamala. Alternation on cellular automata. Theoretical Computer Science, 180(1–2):229–241, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2417.

Andreev:1997:OBA

[692] Alexander E. Andreev, Andrea E. F. Clementi, and José D. P. Rolim. Optimal bounds for the approximation of Boolean functions and some applications. Theoretical Computer Science, 180(1–2):243–268, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2420.

Giakoumakis:1997:ERE

[693] Vassilis Giakoumakis and Jean-Marie Vanherpe. On extended P_{4}-reducible and extended P_{4}-sparse graphs. Theoretical Computer Science, 180(1–2): 269–286, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=180&issue=
1-2&aid=2423.

Codenotti:1997:PAC

[694] Bruno Codenotti, Biswa N. Datta, Karabi Datta, and Mauro Leoncini. Parallel algorithms for certain matrix computations. Theoretical Computer Science, 180(1–2):287–308, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2446.

Karpinski:1997:CCO

[695] Marek Karpinski, Lawrence L. Larmore, and Wojciech Rytter. Correctness of constructing optimal alphabetic trees revisited. Theoretical Computer Science, 180(1–2):309–324, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2496.

Peladeau:1997:FSV

[696] Pierre Péladeau, Howard Straubing, and Denis Thérien. Finite semigroup varieties defined by programs. Theoretical Computer Science, 180(1–2):325–339, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2497.

Kudlek:1997:DC

[697] Manfred Kudlek and Alexandru Mateescu. On distributed catenation. Theoretical Computer Science, 180(1–2):341–352, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=180&issue=1-2&aid=2486.

Dassow:1997:CCF

[698] Jürgen Dassow and Victor Mitrana. Cooperation in context-free grammars. Theoretical Computer Science, 180(1–2):353–361, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=180&issue=1-2&aid=2532.

Sirvent:1997:SDS

[699] Víctor F. Sirvent. On some dynamical subsets of the Rauzy fractal. Theoretical Computer Science, 180(1–2):363–370, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=180&issue=1-2&aid=2533.

David:1997:SPO

[700] René David and Karim Nour. A syntactical proof of the operational equivalence of two λ-terms. Theoretical Computer Science, 180(1–2):371–375, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=180&issue=1-2&aid=2540.

Anonymous:1997:MIV

[701] Anonymous. Master index volumes 171-180. Theoretical Computer Science, 180(1–2):379–??, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeza-Yates:1997:P

[702] Ricardo Baeza-Yates and Eric Goles. Preface. Theoretical Computer Science, 181(1):1–2, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=181&issue=
1&aid=2460.

Asano:1997:SFC

[703] Tetsuo Asano, Desh Ranjan, Thomas Roos, Emo Welzl, and Peter Widmayer. Space-filling curves and their use in the design of geometric data structures. Theoretical Computer Science, 181(1):3–15, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=1&aid=2461.

Bruyere:1997:BNS

[704] Véronique Bruyère and Georges Hansel. Bertrand numeration systems and recognizability. Theoretical Computer Science, 181(1):17–43, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=1&aid=2462.

Chaudhuri:1997:PRR

[705] Shiva Chaudhuri and Devdatt Dubhashi. Probabilistic recurrence relations revisited. Theoretical Computer Science, 181(1):45–56, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=1&aid=2463.

Fernandez-Baca:1997:LTA

[706] David Fernández-Baca and Giora Slutzki. Linear-time algorithms for parametric minimum spanning tree problems on planar graphs. Theoretical Computer Science, 181(1):57–74, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=181&issue=1&aid=2464.

Feuerstein:1997:PMT

[707] Esteban Feuerstein. Paging more than one page. Theoretical Computer Science, 181(1):75–90, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=181&issue=1&aid=2465.

deFigueiredo:1997:ECI

[708] Celina M. H. de Figueiredo, João Meidanis, and Célia Picinin de Mello. On edge-colouring indifference graphs. Theoretical Computer Science, 181 (1):91–106, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=181&issue=
                                        1&aid=2466.

Galbiati:1997:ASM

[709] Giulia Galbiati, Angelo Morzenti, and Francesco Maffioli. On the approximability of some Maximum Spanning Tree Problems. Theoretical Computer Science, 181(1):107–118, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=1&aid=2467.

Gasarch:1997:BSR

[710] William I. Gasarch and Katia S. Guimarães. Binary search and recursive graph problems. Theoretical Computer Science, 181(1):119–139, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=1&aid=2468.

Herzog:1997:PAB

[711] Christian Herzog. Pushdown automata with bounded nondeterminism and bounded ambiguity. Theoretical Computer Science, 181(1):141–157, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=1&aid=2469.

Lopresti:1997:BEM

[712] Daniel Lopresti and Andrew Tomkins. Block edit models for approximate string matching. Theoretical Computer Science, 181(1):159–179, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=1&aid=2470.

Prodinger:1997:PYM

[713] Helmut Prodinger. On a problem of Yekutieli and Mandelbrot about the bifurcation ratio of binary trees. Theoretical Computer Science, 181(1):181–194, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=181&issue=1&aid=2471.

Wang:1997:TLR

[714] Farn Wang. A temporal logic for real-time partial ordering with named transactions. Theoretical Computer Science, 181(1):195–225, July 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=181&issue=1&aid=2472.

Du:1997:F

[715] Dingzhu Du and Ming Li. Foreword. Theoretical Computer Science, 181(2):227, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=181&issue=
                                        2&aid=2473.

Belanger:1997:NNP

[716] Jay Belanger and Jie Wang. No NP problems averaging over ranking of distributions are harder. Theoretical Computer Science, 181(2):229–245, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=2&aid=2474.

Chen:1997:AGE

[717] Jianer Chen. Algorithmic graph embeddings. Theoretical Computer Science, 181(2):247–266, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=2&aid=2475.

Diaz:1997:PAM

[718] Josep Díaz, Alan Gibbons, Grammati E. Pantziou, Maria J. Serna, Paul G. Spirakis, and Jacobo Toran. Parallel algorithms for the minimum cut and the minimum length tree layout problems. Theoretical Computer Science, 181(2):267–287, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=2&aid=2476.

Kobayashi:1997:TPM

[719] Kojiro Kobayashi. Transformations that preserve malignness of universal distributions. Theoretical Computer Science, 181(2):289–306, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=2&aid=2477.

Lingas:1997:MTP

[720] Andrzej Lingas. Maximum treepacking in time O(n^{5}/2). Theoretical Computer Science, 181(2):307–316, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=181&issue=2&aid=2478.

Sakurai:1997:PPK

[721] Kouichi Sakurai. Practical proofs of knowledge without relying on theoretical proofs of membership on languages. Theoretical Computer Science, 181(2):317–335, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=181&issue=2&aid=2479.

Tromp:1997:SWB

[722] John Tromp, Louxin Zhang, and Ying Zhao. Small weight bases for Hamming codes. Theoretical Computer Science, 181(2):337–345, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=181&issue=2&aid=2480.

Wan:1997:CAC

[723] Peng-Jun Wan, Qifan Yang, and Dean Kelley. A (3/2) log 3-competitive algorithm for the counterfeit coin problem. Theoretical Computer Science, 181(2):347–356, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=2&aid=2481.

Yu:1997:STT

[724] Xiangdong Yu and Moti Yung. Scheduling task-trees with additive scales on parallel/distributed machines. Theoretical Computer Science, 181(2):357–378, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=2&aid=2482.

Krumke:1997:CLP

[725] S. O. Krumke, M. V. Marathe, H. Noltemeier, V. Radhakrishnan, S. S. Ravi, and D. J. Rosenkrantz. Compact location problems. Theoretical Computer Science, 181(2):379–404, July 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=181&issue=2&aid=2504.

Czech:1997:PH

[726] Zbigniew J. Czech, George Havas, and Bohdan S. Majewski. Perfect hashing. Theoretical Computer Science, 182(1–2):1–143, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=182&issue=1-2&aid=2349.

Atkinson:1997:BCP

[727] M. D. Atkinson and D. Tulley. Bounded capacity priority queues. Theoretical Computer Science, 182(1–2): 145–157, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=182&issue=
                                        1-2&aid=2352.

Abello:1997:CPS

[728] James Abello and Shlomi Dolev. On the computational power of selfstabilizing systems. Theoretical Computer Science, 182(1–2):159–170, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=182&issue=1-2&aid=2353.

Gao:1997:WSN

[729] B. Gao and F. K. Hwang. Widesense nonblocking for multirate 3stage Clos networks. Theoretical Computer Science, 182(1–2):171–182, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=182&issue=1-2&aid=2354.

Cadoli:1997:CRP

[730] Marco Cadoli, Francesco M. Donini, Marco Schaerf, and Riccardo Silvestri. On compact representations of propositional circumscription. Theoretical Computer Science, 182(1–2):183–202, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=182&issue=1-2&aid=2395.

Koole:1997:ASS

[731] Ger Koole. Assigning a single server to inhomogeneous queues with switching costs. Theoretical Computer Science, 182(1–2):203–216, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=182&issue=1-2&aid=2398.

Mundici:1997:OCS

[732] Daniele Mundici and Alberto Trombetta. Optimal comparison strategies in Ulam’s searching game with two errors. Theoretical Computer Science, 182(1–2):217–232, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=182&issue=1-2&aid=2566.

Bafna:1997:AAM

[733] Vineet Bafna, Eugene L. Lawler, and Pavel A. Pevzner. Approximation algorithms for multiple sequence alignment. Theoretical Computer Science, 182(1–2):233–244, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=182&issue=1-2&aid=2559.

Lam:1997:HFC

[734] Nguyen Huong Lam. Hajós factorizations and completion of codes. Theoretical Computer Science, 182(1–2): 245–256, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=182&issue=
                                        1-2&aid=2568.

Botelho:1997:EBN

[735] Fernanda Botelho and Max Garzon. Erratum to “Boolean neural nets are observable” [Theoret. Comput. Sci. 134 (1994) 51–61]. Theoretical Computer Science, 182(1–2):257, August 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=182&issue=1-2&aid=2649.

See [2].

Anonymous:1997:CEBb

[736] Anonymous. Contents EATCS Bulletin Number 61. Theoretical Computer Science, 182(1–2):259–??, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1997:CEBc

[737] Anonymous. Contents EATCS Bulletin Number 62. Theoretical Computer Science, 182(1–2):263–??, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:1997:C

[738] G. Rozenberg and A. Salomaa. Contributions. Theoretical Computer Science, 183(1):1–??, 1997. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:1997:P

[739] Grzegorz Rozenberg and Arto Salomaa. Preface. Theoretical Computer Science, 183(1):1, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=1&aid=2507.

Ito:1997:IDC

[740] Masami Ito, Lila Kari, and Gabriel Thierrin. Insertion and deletion closure of languages. Theoretical Computer Science, 183(1):3–19, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=1&aid=2508.

Raz:1997:LCC

[741] Danny Raz. Length considerations in context-free languages. Theoretical Computer Science, 183(1):21–32, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=1&aid=2509.

Ilie:1997:CCC

[742] Lucian Ilie. On computational complexity of contextual languages. Theoretical Computer Science, 183(1): 33–44, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=183&issue=
1&aid=2510.

deLuca:1997:SWS

[743] Aldo de Luca. Sturmian words: Structure, combinatorics, and their arithmetics. Theoretical Computer Science, 183(1):45–82, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=1&aid=2511.

Choffrut:1997:NDQ

[744] C. Choffrut, T. Harju, and J. Karhumäki. A note on decidability questions on presentations of word semigroups. Theoretical Computer Science, 183(1): 83–92, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=183&issue=
1&aid=2512.

Maler:1997:SCL

[745] Oded Maler and Ludwig Staiger. On syntactic congruences for ωlanguages. Theoretical Computer Science, 183(1):93–112, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=1&aid=2513.

Honkala:1997:LAP

[746] Juha Honkala and Werner Kuich. On Lindenmayerian algebraic power series. Theoretical Computer Science, 183(1):113–142, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=1&aid=2514.

Honkala:1997:LAS

[747] Juha Honkala. On Lindenmayerian algebraic sequences. Theoretical Computer Science, 183(1):143–154, August 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=1&aid=2515.

Alagar:1997:F

[748] V. S. Alagar. Foreword. Theoretical Computer Science, 183(2):155– 156, September 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=183&issue=
2&aid=2523.

Haxthausen:1997:OSA

[749] Anne Elisabeth Haxthausen. Ordersorted algebraic specifications with higher-order functions. Theoretical Computer Science, 183(2):157– 185, September 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=183&issue=
                                        2&aid=2524.

Montanari:1997:TSM

[750] Angelo Montanari and Maarten de Rijke. Two-sorted metric temporal logics. Theoretical Computer Science, 183(2):187–214, September 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=183&issue=2&aid=2525.

Dam:1997:DPE

[751] Mads Dam. On the decidability of process equivalences for the pi— calculus. Theoretical Computer Science, 183(2):215–228, September 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=183&issue=2&aid=2526.

Cheng:1997:PNT

[752] Allan Cheng. Petri nets, traces, and local model checking. Theoretical Computer Science, 183(2):229–251, September 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=183&issue=2&aid=2527.

Collette:1997:FMR

[753] Pierre Collette and Edgar Knapp. A foundation for modular reasoning about safety and progress properties of state-based concurrent programs. Theoretical Computer Science, 183(2):253–279, September 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=2&aid=2528.

Falaschi:1997:CCC

[754] Moreno Falaschi, Maurizio Gabbrielli, Kim Marriott, and Catuscia Palamidessi. Confluence in concurrent constraint programming. Theoretical Computer Science, 183(2):281–315, September 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=183&issue=2&aid=2529.

Brogi:1997:UVL

[755] Antonio Brogi, Evelina Lamma, Paolo Mancarella, and Paola Mello. A unifying view for logic programming with non-monotonic reasoning. Theoretical Computer Science, 184(1–2):1–59, September 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=184&issue=1-2&aid=2425.

Wagner:1997:LCC

[756] Kim Ritter Wagner. Liminf convergence in Ω-categories. Theoretical Computer Science, 184(1–2):61– 104, September 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/

                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=184&issue=
                                        1-2&aid=2426.

Andrews:1997:LSD

[757] James Andrews. A logical semantics for depth-first Prolog with ground negation. Theoretical Computer Science, 184(1–2):105–143, September 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=184&issue=1-2&aid=2341.

Burmeister:1997:ATU

[758] P. Burmeister, F. Rossello, J. Torrens, and G. Valiente. Algebraic transformation of unary partial algebras. I. Double-pushout approach. Theoretical Computer Science, 184(1–2):145–193, September 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=184&issue=1-2&aid=2342.

Su:1997:DCO

[759] Jianwen Su. Dynamic constraints and object migration. Theoretical Computer Science, 184(1–2):195– 236, September 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=184&issue=
                                        1-2&aid=2345.

Lacoste:1997:LP

[760] Thierry Lacoste. 0–1 laws by preservation. Theoretical Computer Science, 184(1–2):237–245, September 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=184&issue=1-2&aid=2622.

Pierce:1997:CHO

[761] Benjamin Pierce and Martin Steffen. Corrigendum to “Higher-order subtyping” [Theoret. Comput. Sci. 176(1–2) (1997) 235–282]. Theoretical Computer Science, 184(1–2):247, September 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=184&issue=1-2&aid=2648.

See [625].

Zeugmann:1997:F

[762] Thomas Zeugmann. Foreword. Theoretical Computer Science, 185(1): 1, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=185&issue=
1&aid=2548.

Kececioglu:1997:IDS

[763] John Kececioglu, Ming Li, and John Tromp. Inferring a DNA sequence from erroneous copies. Theoretical Computer Science, 185(1):3–13, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=1&aid=2549.

Sakakibara:1997:RAG

[764] Yasubumi Sakakibara. Recent advances of grammatical inference. Theoretical Computer Science, 185(1): 15–45, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=185&issue=
1&aid=2550.

Arimura:1997:LUT

[765] Hiroki Arimura, Hiroki Ishizaka, and Takeshi Shinohara. Learning unions of tree patterns using queries. Theoretical Computer Science, 185(1):47–62, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=1&aid=2551.

Koshiba:1997:LDE

[766] Takeshi Koshiba, Erkki Mäkinen, and Yuji Takada. Learning deterministic even linear languages from positive examples. Theoretical Computer Science, 185(1):63–79, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=1&aid=2552.

Meyer:1997:PLL

[767] Léa Meyer. Probabilistic language learning under monotonicity constraints. Theoretical Computer Science, 185(1):81–128, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=1&aid=2553.

Stephan:1997:NIO

[768] Frank Stephan. Noisy inference and oracles. Theoretical Computer Science, 185(1):129–157, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=1&aid=2554.

Auer:1997:LND

[769] Peter Auer. Learning nested differences in the presence of malicious noise. Theoretical Computer Science, 185(1):159–175, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=1&aid=2555.

Takimoto:1997:LOH

[770] Eiji Takimoto, Akira Miyashiro, Akira Maruoka, and Yoshifumi Sakai. Learning orthogonal F -Horn formulas. Theoretical Computer Science, 185(1):177–190, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=185&issue=1&aid=2556.

Rao:1997:FIL

[771] M. R. K. Krishna Rao. A framework for incremental learning of logic programs. Theoretical Computer Science, 185(1):191–213, October 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=185&issue=1&aid=2557.

Staples:1997:P

[772] John Staples. Preface. Theoretical Computer Science, 185(2):215, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=185&issue=2&aid=2579.

Albrecht:1997:NCH

[773] David Albrecht, John N. Crossley, and John S. Jeavons. New Curry–Howard terms for full linear logic. Theoretical Computer Science, 185(2):217–235, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=185&issue=2&aid=2580.

Jay:1997:CT

[774] C. Barry Jay. Covariant types. Theoretical Computer Science, 185(2):237– 258, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=185&issue=
2&aid=2581.

Lin:1997:FDQ

[775] Xuemin Lin. A fully distributed quorum consensus method with high fault-tolerance and low communication overhead. Theoretical Computer Science, 185(2):259–275, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=2&aid=2582.

Mason:1997:FOL

[776] Ian A. Mason. A first order logic of effects. Theoretical Computer Science, 185(2):277–318, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=185&issue=2&aid=2583.

Orgun:1997:MDL

[777] Mehmet A. Orgun and Weichang Du. Multi-dimensional logic programming: Theoretical foundations. Theoretical Computer Science, 185(2):319– 345, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=185&issue=
2&aid=2584.

Pantziou:1997:MMR

[778] Grammati E. Pantziou, Alan Roberts, and Antonios Symvonis. Many-to-many routing on trees via matchings. Theoretical Computer Science, 185(2):347–377, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=185&issue=2&aid=2585.

Vincent:1997:CDR

[779] Millist W. Vincent. A corrected 5NF definition for relational database design. Theoretical Computer Science, 185(2):379–391, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=185&issue=2&aid=2586.

Weibel:1997:OSR

[780] Trudy Weibel. An order-sorted resolution in theory and practice. Theoretical Computer Science, 185(2):393– 410, October 20, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=185&issue=
                                        2&aid=2587.

Anonymous:1997:AIV

[781] Anonymous. Author index volume 185. Theoretical Computer Science, 185(2):411–??, 1997. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malgouyres:1997:DSN

[782] Rémy Malgouyres. A definition of surfaces of Z^{3}: a new 3D discrete Jordan theorem. Theoretical Computer Science, 186(1–2):1–41, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=186&issue=1-2&aid=2416.

Taentzer:1997:PHL

[783] Gabriele Taentzer. Parallel high-level replacement systems. Theoretical Computer Science, 186(1–2):43–81, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=186&issue=1-2&aid=2418.

Leiss:1997:SSE

[784] Ernst L. Leiss. Solving systems of explicit language relations. Theoretical Computer Science, 186(1–2):83–105, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=186&issue=1-2&aid=2421.

Badouel:1997:SPE

[785] Eric Badouel, Luca Bernardinello, and Philippe Darondeau. The synthesis problem for elementary net systems is NP-complete. Theoretical Computer Science, 186(1–2):107–134, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=186&issue=1-2&aid=2422.

Peled:1997:PSP

[786] Doron Peled. On projective and separable properties. Theoretical Computer Science, 186(1–2):135–156, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=186&issue=1-2&aid=2428.

Kim:1997:HEF

[787] Changwook Kim. A hierarchy of eNCE families of graph languages. Theoretical Computer Science, 186(1–2):157–169, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=186&issue=1-2&aid=2429.

Flammini:1997:DBR

[788] Michele Flammini and Giorgio Gambosi. On devising Boolean Routing Schemes. Theoretical Computer Science, 186(1–2):171–198, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=186&issue=1-2&aid=2487.

Afek:1997:LDP

[789] Yehuda Afek, Shay Kutten, and Moti Yung. The local detection paradigm and its applications to selfstabilization. Theoretical Computer Science, 186(1–2):199–230, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=186&issue=1-2&aid=2488.

Ohlebusch:1997:EPP

[790] Enno Ohlebusch and Esko Ukkonen. On the equivalence problem for Epattern languages. Theoretical Computer Science, 186(1–2):231–248, October 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=186&issue=1-2&aid=2491.

Anonymous:1997:CAR

[791] Anonymous. Computer algebra. 5th Rhine Workshop (RWCA). Theoretical Computer Science, 187(1–2):??, November 15, 1997. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calmet:1997:F

[792] Jacques Calmet and Alain Carrière. Foreword. Theoretical Computer Science, 187(1–2):1, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2588.

Geiselman:1997:CTO

[793] Willi Geiselman and Felix Ulmer. Constructing a third-order linear differential equation. Theoretical Computer Science, 187(1–2):3–6, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2589.

Hubert:1997:DDB

[794] Evelyne Hubert. Detecting degenerate behaviors in first order algebraic differential equations. Theoretical Computer Science, 187(1–2):7–25, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2590.

Fakler:1997:SOH

[795] Winfried Fakler. On second order homogeneous linear differential equations with Liouvillian solutions. Theoretical Computer Science, 187(1–2):27–48, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2591.

Thomas:1997:PDS

[796] G. Thomas. The problem of defining the singular points of quasi-linear differential-algebraic systems. Theoretical Computer Science, 187(1–2): 49–79, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

cgi?year=1997&volume=187&issue=
1-2&aid=2592.

Pflugel:1997:LVD

[797] E. Pflügel. On the latest version of DESIR-II. Theoretical Computer Science, 187(1–2):81–86, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2593.

Scheen:1997:IPT

[798] Christian Scheen. Implementation of the Painlevé test for ordinary differential systems. Theoretical Computer Science, 187(1–2):87–104, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2594.

Bernardin:1997:SFF

[799] Laurent Bernardin. On square-free factorization of multivariate polynomials over a finite field. Theoretical Computer Science, 187(1–2):105–116, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2595.

deGraaf:1997:ADS

[800] W. A. de Graaf. An algorithm for the decomposition of semisimple Lie algebras. Theoretical Computer Science, 187(1–2):117–122, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=187&issue=1-2&aid=2596.

Makhlouf:1997:AAC

[801] Abdenacer Makhlouf. Algèbres associatives et calcul formel. (French) [Associative algebras and computer algebra]. Theoretical Computer Science, 187(1–2):123–145, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=187&issue=1-2&aid=2597.

Zenger:1997:IT

[802] Christoph Zenger. Indexed types. Theoretical Computer Science, 187(1–2): 147–165, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=187&issue=
                                        1-2&aid=2598.

Mall:1997:CFP

[803] Daniel Mall. Covers and fans of polynomial ideals. Theoretical Computer Science, 187(1–2):167–178, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=187&issue=1-2&aid=2599.

Amrhein:1997:W

[804] Beatrice Amrhein, Oliver Gloor, and Wolfgang Küchlin. On the Walk. Theoretical Computer Science, 187(1–2):179–202, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2600.

Karczmarczuk:1997:GPL

[805] Jerzy Karczmarczuk. Generating power of lazy semantics. Theoretical Computer Science, 187(1–2):203– 219, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1997&volume=187&issue=
1-2&aid=2601.

Calmet:1997:TMS

[806] Jacques Calmet and Karsten Homann. Towards the Mathematics Software Bus. Theoretical Computer Science, 187(1–2):221–230, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=187&issue=1-2&aid=2602.

Korelc:1997:AGF

[807] Jože Korelc. Automatic generation of finite-element code by simultaneous optimization of expressions. Theoretical Computer Science, 187(1–2):231–248, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=187&issue=1-2&aid=2603.

ElHamidi:1997:UMA

[808] A. El Hamidi and M. Garbey. Using MAPLE for the analysis of bifurcation phenomena in gas combustion. Theoretical Computer Science, 187(1–2):249–262, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=187&issue=1-2&aid=2604.

Carriere:1997:ADC

[809] Alain Carrière and Louis-Rémi Oudin. Applications du calcul formel à la balistique. (French) [Applications of formal calculus to ballistics]. Theoretical Computer Science, 187(1–2):263–284, November 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=187&issue=1-2&aid=2605.

Nagaraj:1997:OBS

[810] S. V. Nagaraj. Optimal binary search trees. Theoretical Computer Science, 188(1–2):1–44, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=188&issue=1-2&aid=2521.

Heen:1997:LSC

[811] Olivier Heen. Linear speed-up for cellular automata synchronizers and applications. Theoretical Computer Science, 188(1–2):45–57, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=188&issue=1-2&aid=2414.

Sen:1997:LBP

[812] Sandeep Sen. Lower bounds for parallel algebraic decision trees: parallel complexity of convex hulls and related problems. Theoretical Computer Science, 188(1–2):59–78, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=188&issue=1-2&aid=2415.

Romanik:1997:ATR

[813] Kathleen Romanik. Approximate testing and its relationship to learning. Theoretical Computer Science, 188(1–2):79–99, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=188&issue=1-2&aid=2443.

Regan:1997:GLL

[814] Kenneth W. Regan and Heribert Vollmer. Gap-languages and logtime complexity classes. Theoretical Computer Science, 188(1–2):101– 116, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.

                                        cgi?year=1997&volume=188&issue=
                                        1-2&aid=2490.

Grolmusz:1997:PCG

[815] Vince Grolmusz. On the power of circuits with gates of low L_{1} norms. Theoretical Computer Science, 188(1–2):117–128, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=188&issue=1-2&aid=2492.

Goles:1997:CTR

[816] Eric Goles and Iván Rapaport. Complexity of tile rotation problems. Theoretical Computer Science, 188(1–2):129–159, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=188&issue=1-2&aid=2493.

Cerny:1997:SRI

[817] Anton Černý. On sequences resulting from iteration of modified quadratic and palindromic mappings. Theoretical Computer Science, 188(1–2):161–174, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=188&issue=1-2&aid=2494.

Freivalds:1997:KNM

[818] Rūsin¸š Freivalds and Sanjay Jain. Kolmogorov numberings and minimal identification. Theoretical Computer Science, 188(1–2):175–194, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=188&issue=1-2&aid=2495.

Allouche:1997:ADS

[819] J.-P. Allouche, F. von Haeseler, H.-O. Peitgen, A. Petersen, and G. Skordev. Automaticity of double sequences generated by one-dimensional linear cellular automata. Theoretical Computer Science, 188(1–2):195–209, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=188&issue=1-2&aid=2498.

Gyuris:1997:SPR

[820] Viktor Gyuris. A short proof of representability of fork algebras. Theoretical Computer Science, 188(1–2):211–220, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=188&issue=1-2&aid=2499.

Shen:1997:OAG

[821] Hong Shen. Optimal algorithms for generalized searching in sorted matrices. Theoretical Computer Science, 188(1–2):221–230, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=188&issue=1-2&aid=2563.

Long:1997:SIO

[822] Dong Yang Long, Jian Ma, and Duanning Zhou. Structure of 3-infix–outfix maximal codes. Theoretical Computer Science, 188(1–2):231–240, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=188&issue=1-2&aid=2633.

Liu:1997:ISA

[823] Renren Liu. An improved Shellsort algorithm. Theoretical Computer Science, 188(1–2):241–247, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=188&issue=1-2&aid=2636.

Niwinski:1997:FPC

[824] Damian Niwiński. Fixed point characterization of infinite behavior of finite-state systems. Theoretical Computer Science, 189(1–2):1–69, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=189&issue=1-2&aid=2575.

Bouge:1997:FVD

[825] Luc Bougé, David Cachera, Yann Le Guyadec, Gil Utard, and Bernard Virot. Formal validation of dataparallel programs: a two-component assertional proof system for a simple language. Theoretical Computer Science, 189(1–2):71–107, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=189&issue=1-2&aid=2577.

Burkhard:1997:FCM

[826] Hans-Dieter Burkhard. Fairness and control in multi-agent systems. Theoretical Computer Science, 189(1–2):109–127, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=189&issue=1-2&aid=2391.

Eiter:1997:ALP

[827] Thomas Eiter, Georg Gottlob, and Nicola Leone. Abduction from logic programs: semantics and complexity. Theoretical Computer Science, 189(1–2):129–177, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1997&volume=189&issue=1-2&aid=2392.

Bremond-Gregoire:1997:PAC

[828] Patrice Brémond-Grégoire and Insup Lee. A process algebra of communicating shared resources with dense time and priorities. Theoretical Computer Science, 189(1–2):179–219, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=189&issue=1-2&aid=2432.

Mezghiche:1997:MR

[829] Mohamed Mezghiche. cβ-machine with λβ-reduction. Theoretical Computer Science, 189(1–2):221–228, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1997&volume=189&issue=1-2&aid=2576.

Bergeron:1997:RBC

[830] Anne Bergeron. On the rational behaviors of concurrent timers. Theoretical Computer Science, 189(1–2): 229–237, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=189&issue=
                                        1-2&aid=2610.

Cicerone:1997:CSM

[831] Serafino Cicerone and Francesco Parisi-Presicce. On the complexity of specification morphisms. Theoretical Computer Science, 189(1–2):239– 248, December 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1997&volume=189&issue=
                                        1-2&aid=2642.

Kirchner:1998:E

[832] H. Kirchner. Editorial. Theoretical Computer Science, 190(1):1–2, January 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
190/1/2617.pdf.

Fernandez:1998:INT

[833] Maribel Fernández and Ian Mackie. Interaction nets and term-rewriting systems. Theoretical Computer Science, 190(1):3–39, January 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
190/1/2618.pdf.

Kaivola:1998:AEC

[834] Roope Kaivola. Axiomatising extended computation tree logic. Theoretical Computer Science, 190(1):41– 60, January 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/190/1/2619.pdf.

Lisper:1998:CUE

[835] Björn Lisper. Computing in unpredictable environments: semantics, reduction strategies, and program transformations. Theoretical Computer Science, 190(1):61–85, January 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
190/1/2620.pdf.

Cheng:1998:OMB

[836] Allan Cheng and Mogens Nielsen. Open maps, behavioural equivalences, and congruences. Theoretical Computer Science, 190(1):87–112, January 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        190/1/2621.pdf.

Gottlob:1998:F

[837] G. Gottlob and M. Y. Vardi. Foreword. Theoretical Computer Science, 190(2):113–??, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bidoit:1998:FST

[838] N. Bidoit and S. De Amo. A first step towards implementing dynamic algebraic dependences. Theoretical Computer Science, 190(2):115–149, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        190/2/2624.pdf.

Demetrovics:1998:APK

[839] J. Demetrovics, G. O. H. Katona, D. Miklos, O. Seleznjev, and B. Thalheim. Asymptotic properties of keys and functional dependencies in random databases. Theoretical Computer Science, 190(2):151–166, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        190/2/2625.pdf.

Libkin:1998:MAD

[840] Leonid Libkin. Models of approximation in databases. Theoretical Computer Science, 190(2):167–210, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
190/2/2626.pdf.

Lifschitz:1998:PVD

[841] Sérgio Lifschitz and Victor Vianu. A probabilistic view of Datalog parallelization. Theoretical Computer Science, 190(2):211–239, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
190/2/2627.pdf.

Marek:1998:RP

[842] Victor W. Marek and Miroslaw Truszczyński. Revision programming. Theoretical Computer Science, 190(2): 241–277, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/190/2/2628.pdf.

Suciu:1998:DIQ

[843] Dan Suciu. Domain-independent queries on databases with external functions. Theoretical Computer Science, 190(2):279–315, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
190/2/2629.pdf.

Tyszkiewicz:1998:KEP

[844] Jerzy Tyszkiewicz. The Kolmogorov expressive power of Boolean query languages. Theoretical Computer Science, 190(2):317–361, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        190/2/2630.pdf.

Vingralek:1998:UCC

[845] R. Vingralek, H. Hasse-Ye, Y. Breitbart, and H.-J. Schek. Unifying concurrency control and recovery of transactions with semantically rich operations. Theoretical Computer Science, 190(2): 363–396, January 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/190/2/2631.pdf.

Naudin:1998:UPF

[846] Patrice Naudin and Claude Quitté. Univariate polynomial factorization over finite fields. Theoretical Computer Science, 191(1–2):1–36, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2546.pdf.

Selivanov:1998:FHR

[847] Victor Selivanov. Fine hierarchy of regular ω-languages. Theoretical Computer Science, 191(1–2):37–59, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2501.pdf.

Frougny:1998:SDA

[848] Christiane Frougny and Jacques Sakarovitch. Synchronisation deterministe des automates à delai borné.

(French) [Synchronisation of deterministic automata with bounded delay]. Theoretical Computer Science, 191(1– 2):61–77, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/191/1-2/2505.pdf.

Mihoubi:1998:CCP

[849] Douadi Mihoubi. Characterization and closure properties of linear ωlanguages. Theoretical Computer Science, 191(1–2):79–95, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
191/1-2/2517.pdf.

Downey:1998:PCC

[850] Rodney G. Downey, Michael R. Fellows, and Kenneth W. Regan. Parameterized circuit complexity and the W hierarchy. Theoretical Computer Science, 191(1–2):97–115, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
191/1-2/2518.pdf.

Bruyere:1998:MOL

[851] Véronique Bruyère, Denis Derencourt, and Michel Latteux. The meet operation in the lattice of codes. Theoretical Computer Science, 191(1–2):117– 129, January 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/191/1-2/2519.pdf.

Breslauer:1998:STT

[852] Dany Breslauer. The suffix tree of a tree and minimizing sequential transducers. Theoretical Computer Science, 191(1–2):131–144, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2520.pdf.

Koscielski:1998:MAP

[853] Antoni Kościelski and Leszek Pacholski. Makanin’s algorithm is not primitive recursive. Theoretical Computer Science, 191(1–2):145–156, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2522.pdf.

Ferguson:1998:SCT

[854] Thomas S. Ferguson. Some chip transfer games. Theoretical Computer Science, 191(1–2):157–171, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2671.pdf.

Niemi:1998:SMC

[855] Valtteri Niemi and Ari Renvall. Secure multiparty computations without computers. Theoretical Computer Science, 191(1–2):173–183, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2643.pdf.

Kautz:1998:IZO

[856] Steven M. Kautz. An improved zeroone law for algorithmically random sequences. Theoretical Computer Science, 191(1–2):185–192, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
191/1-2/2652.pdf.

Eriksson:1998:MFT

[857] Gerd Eriksson, Henrik Eriksson, and Kimmo Eriksson. Moving a food trolley around a corner. Theoretical Computer Science, 191(1–2):193–203, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
191/1-2/2657.pdf.

Middendorf:1998:SCS

[858] Martin Middendorf. Shortest common superstrings and scheduling with coordinated starting times. Theoretical Computer Science, 191(1–2):205– 214, January 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/191/1-2/2669.pdf.

Goldreich:1998:CIA

[859] Oded Goldreich and Bernd Meyer. Computational indistinguishability: algorithms vs. circuits. Theoretical Computer Science, 191(1–2):215–218, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
191/1-2/2698.pdf.

Wang:1998:FDT

[860] Jing Wang. Finite derivation type for semi-direct products of monoids. Theoretical Computer Science, 191(1–2): 219–228, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/191/1-2/2700.pdf.

Angel:1998:ACG

[861] E. Angel and V. Zissimopoulos. Autocorrelation coefficient for the graph bipartitioning problem. Theoretical Computer Science, 191(1–2):229–243, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2712.pdf.

Beigel:1998:ASA

[862] Richard Beigel, Bill Gasarch, Ming Li, and Louxin Zhang. Addition in log_{2} n + O(1) steps on average: A simple analysis. Theoretical Computer Science, 191(1–2):245–248, January 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        191/1-2/2717.pdf.

Hsiang:1998:F

[863] Jieh Hsiang. Foreword. Theoretical Computer Science, 192(1):1, February 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        192/1/2678.pdf.

Mayr:1998:HOR

[864] Richard Mayr and Tobias Nipkow. Higher-order rewrite systems and their confluence. Theoretical Computer Science, 192(1):3–29, February 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
192/1/2679.pdf.

Marchiori:1998:BM

[865] Massimo Marchiori. Bubbles in modularity. Theoretical Computer Science, 192(1):31–54, February 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
192/1/2680.pdf.

Senizergues:1998:PAT

[866] Géraud Sénizergues. A polynomial algorithm testing partial confluence of basic semi-Thue systems. Theoretical Computer Science, 192(1):55– 75, February 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/192/1/2681.pdf.

Anantharaman:1998:RML

[867] Siva Anantharaman and Gilles Richard. A rewrite mechanism for logic programs with negation. Theoretical Computer Science, 192(1):77–106, February 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
192/1/2682.pdf.

Baader:1998:CCS

[868] Franz Baader and Klaus U. Schulz. Combination of constraint solvers for free and quasi-free structures. Theoretical Computer Science, 192(1):107– 161, February 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/192/1/2683.pdf.

Gorrieri:1998:F

[869] R. Gorrieri and C. Hankin. Foreword. Theoretical Computer Science, 192(2):163–??, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busi:1998:PAV

[870] Nadia Busi, Roberto Gorrieri, and Gianluigi Zavattaro. A process algebraic view of Linda coordination primitives. Theoretical Computer Science, 192(2): 167–199, February 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/192/2/2685.pdf.

Dami:1998:LCD

[871] Laurent Dami. A lambda-calculus for dynamic binding. Theoretical Computer Science, 192(2):201–231, February 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        192/2/2686.pdf.

Hankin:1998:RMT

[872] Chris Hankin, Daniel Le Métayer, and David Sands. Refining multiset transformers. Theoretical Computer Science, 192(2):233–258, February 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
192/2/2687.pdf.

Monteiro:1998:EBA

[873] Luís Monteiro and António Porto. Entailment-based actions for coordination. Theoretical Computer Science, 192(2):259–286, February 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
192/2/2688.pdf.

Mukherji:1998:PCB

[874] Manibrata Mukherji and Dennis Kafura. A process-calculus-based abstraction for coordinating multi-agent groups. Theoretical Computer Science, 192(2):287–314, February 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
192/2/2689.pdf.

Wegner:1998:IFC

[875] Peter Wegner. Interactive foundations of computing. Theoretical Computer Science, 192(2):315–351, February 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
192/2/2690.pdf.

Bonsangue:1998:GMS

[876] M. M. Bonsangue, F. van Breugel, and J. J. M. M. Rutten. Generalized metric spaces: completion, topology, and power domains via the Yoneda embedding. Theoretical Computer Science, 193(1–2):1–51, February 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        193/1-2/2578.pdf.

Edalat:1998:CMM

[877] Abbas Edalat and Reinhold Heckmann. A computational model for metric spaces. Theoretical Computer Science, 193(1–2):53–73, February 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        193/1-2/2445.pdf.

Ghelli:1998:BEM

[878] Giorgio Ghelli and Benjamin Pierce. Bounded existentials and minimal typing. Theoretical Computer Science, 193(1–2):75–96, February 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        193/1-2/2500.pdf.

Manabe:1998:ASG

[879] Yoshifumi Manabe, Roberto Baldoni, Michel Raynal, and Shigemi Aoyagi. k-arbiter: A safe and general scheme for h-out of-k mutual exclusion. Theoretical Computer Science, 193(1–2):97– 112, February 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/193/1-2/2503.pdf.

Alessi:1998:CDB

[880] Fabio Alessi and Paolo Baldan. A characterization of distance between 1bounded compact ultrametric spaces through a universal space. Theoretical Computer Science, 193(1–2):113– 127, February 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/193/1-2/2537.pdf.

Esbelin:1998:RRP

[881] Henri-Alex Esbelin and Malika More. Rudimentary relations and primitive recursion: A toolbox. Theoretical Computer Science, 193(1–2):129–148, February 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/193/1-2/2538.pdf.

Ross:1998:FAC

[882] Kenneth A. Ross, Divesh Srivastava, Peter J. Stuckey, and S. Sudarshan. Foundations of aggregation constraints. Theoretical Computer Science, 193(1–2):149–179, February 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
193/1-2/2547.pdf.

Drakengren:1998:DCR

[883] Thomas Drakengren. A decidable canonical representation of the compact elements in Scott’s reflexive domain in P ω. Theoretical Computer Science, 193(1–2):181–195, February 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        193/1-2/2574.pdf.

Rabinovich:1998:TTL

[884] A. Rabinovich. On translations of temporal logic of actions into monadic second-order logic. Theoretical Computer Science, 193(1–2):197– 214, February 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/193/1-2/2611.pdf.

Cadoli:1998:CDE

[885] Marco Cadoli and Luigi Palopoli. Circumscribing DATALOG: expressive power and complexity. Theoretical Computer Science, 193(1–2):215–244, February 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/193/1-2/2644.pdf.

Anonymous:1998:JCS

[886] Anonymous. Joint COMPUGRAPH/ SEMAGRAPH Workshop on Graph Rewriting and Computation (SEGRA-GRA ’95) (papers in summary form only received). Theoretical Computer Science, 194(1–2):??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flajolet:1998:CFA

[887] Philippe Flajolet and Brigitte Vallée. Continued fraction algorithms, functional operators, and structure constants. Theoretical Computer Science, 194(1–2):1–34, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
194/1-2/2659.pdf.

Fernau:1998:RRL

[888] Henning Fernau and Dietmar Wätjen. Remarks on regulated limited ET0L systems and regulated context-free grammars. Theoretical Computer Science, 194(1–2):35–55, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
194/1-2/2531.pdf.

Danyi:1998:CSD

[889] G. Dányi and Z. Fülöp. Compositions with superlinear deterministic top-down tree transformations. Theoretical Computer Science, 194(1–2): 57–85, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/194/1-2/2534.pdf.

Gyenizse:1998:LGS

[890] Pál Gyenizse and Sándor Vágvölgyi. Linear generalized semi-monadic rewrite systems effectively preserve recognizability. Theoretical Computer Science, 194(1–2):87–122, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        194/1-2/2535.pdf.

Wan:1998:TML

[891] Peng-Jun Wan. TWDM multichannel lightwave hypercube networks. Theoretical Computer Science, 194(1–2): 123–136, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/194/1-2/2539.pdf.

Niedermeier:1998:UCL

[892] Rolf Niedermeier and Peter Rossmanith. Unambiguous computations and locally definable acceptance types. Theoretical Computer Science, 194(1– 2):137–161, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/194/1-2/2541.pdf.

Irani:1998:RAM

[893] Sandy Irani and Steve Seiden. Randomized algorithms for metrical task systems. Theoretical Computer Science, 194(1–2):163–182, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        194/1-2/2542.pdf.

Chebotarev:1998:RBA

[894] Anatoli N. Chebotarev and Marina K. Morokhovets. Resolution-based approach to compatibility analysis of interacting automata. Theoretical Computer Science, 194(1–2):183–205, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/194/1-2/2561.pdf.

Hadjiat:1998:PGH

[895] Malika Hadjiat. Penelope’s graph: a hard minimum cost tension instance. Theoretical Computer Science, 194(1– 2):207–218, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/194/1-2/2722.pdf.

Ceola:1998:CMD

[896] Corine Ceola and Pierre B. A. Lecomte. Computability of a map and decidability of its graph in the model of Blum, Shub and Smale. Theoretical Computer Science, 194(1–2):219– 223, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/194/1-2/2713.pdf.

Fraigniaud:1998:XPM

[897] Pierre Fraigniaud. On XRAM and PRAM models, and on datamovement-intensive problems. Theoretical Computer Science, 194(1–2): 225–237, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/194/1-2/2725.pdf.

Banach:1998:DRA

[898] R. Banach. DPO rewriting and abstract semantics via opfibrations. Theoretical Computer Science, 194(1–2):

240–??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barendsen:1998:DSU

[899] E. Barendsen and S. Smetsers. A derivation system for uniqueness typing. Theoretical Computer Science, 194 (1–2):240–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bauderon:1998:PRG

[900] M. Bauderon. Parallel rewriting of graphs through the pullback approach. Theoretical Computer Science, 194(1– 2):240–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Best:1998:UNR

[901] E. Best and M. Koutny. Using net refinement to compute the fixpoint of a recursive expression. Theoretical Computer Science, 194(1–2):241– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brock:1998:PSF

[902] S. Brock and G. Ostheimer. A process semantics for functional programming. Theoretical Computer Science, 194(1– 2):241–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clark:1998:SPN

[903] D. Clark and R. Kennaway. Some properties of non-orthogonal term graph rewriting systems. Theoretical Computer Science, 194(1–2):241–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corradini:1998:CAS

[904] A. Corradini and R. Heckel. A compositional approach to structuring and refinement of typed graph grammars. Theoretical Computer Science, 194(1– 2):241–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Corradini:1998:CCP

[905] A. Corradini. Concurrent computing: from Petri nets to graph grammars. Theoretical Computer Science, 194(1– 2):241–242, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Courcelle:1998:LG

[906] B. Courcelle. Logic and graphs. Theoretical Computer Science, 194(1–2): 242–??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drappa:1998:UGG

[907] A. Drappa and R. Melchisedech. The use of graph grammar in a software engineering education tool. Theoretical Computer Science, 194(1–2):242– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drewes:1998:STG

[908] F. Drewes. Semirings and tree-to-graph-to-tree transductions. Theoretical Computer Science, 194(1–2):242– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehrig:1998:IC

[909] H. Ehrig. Introduction to COM-PUGRAPH. Theoretical Computer Science, 194(1–2):242–243, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Engels:1998:EHG

[910] G. Engels and A. Schurr. Encapsulated hierarchical graphs, graph types, and meta types. Theoretical Computer Science, 194(1–2):243–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Habel:1998:URN

[911] A. Habel and D. Plump. Unification, rewriting, and narrowing on term graphs. Theoretical Computer Science, 194(1–2):243–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heckel:1998:ECC

[912] R. Heckel and A. Wagner. Ensuring consistency of conditional graph rewriting — a constructive approach. Theoretical Computer Science, 194(1–2): 243–??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heimann:1998:PGR

[913] P. Heimann, G. Joeris, C.-A. Krapp, and B. Westfechtel. A programmed graph rewriting system for software process management. Theoretical Computer Science, 194(1–2):243–244, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Janssens:1998:PLE

[914] D. Janssens. Process languages for ESM systems. Theoretical Computer Science, 194(1–2):244–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Johnsson:1998:GRH

[915] T. Johnsson. Graph reduction, and how to avoid it. Theoretical Computer Science, 194(1–2):244–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kennaway:1998:IRC

[916] R. Kennaway. Infinitary rewriting and cyclic graphs. Theoretical Computer Science, 194(1–2):244–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Khasidashvili:1998:CSC

[917] Z. Khasidashvili and V. Van Oostrom. Context-sensitive conditional expression reductions systems. Theoretical Computer Science, 194(1–2):244– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Korff:1998:CDS

[918] M. Korff and L. Ribeiro. Concurrent derivations as single pushout graph grammar processes. Theoretical Computer Science, 194(1–2):245– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kreowski:1998:SPG

[919] H.-J. Kreowski. Specification and programming (by graph transformation).

Theoretical Computer Science, 194(1– 2):245–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuske:1998:IBR

[920] S. Kuske. Implementing beta-reduction by hypergraph rewriting. Theoretical Computer Science, 194(1–2):245– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Litovsky:1998:CGG

[921] I. Litovsky, Y. Metivier, and E. Sopena. Checking global graph properties by means of local computations: the majority problem. Theoretical Computer Science, 194(1–2):245–246, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Monserrat:1998:HRU

[922] M. Monserrat, F. Rossello, J. Torrens, and G. Valiente. Hypergraph rewriting using conformisms. Theoretical Computer Science, 194(1–2):246– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Plasmeijer:1998:CPE

[923] M. J. Plasmeijer. CLEAN: a programming environment based on term graph rewriting. Theoretical Computer Science, 194(1–2):246–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Quemener:1998:MCI

[924] Y.-M. Quemener and T. Jeron. Modelchecking of infinite Kripke structures defined by simple graph grammars.

Theoretical Computer Science, 194(1– 2):246–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schied:1998:LTH

[925] G. Schied and K. Barthelmann. Linear types for higher order processes with first class directed channels. Theoretical Computer Science, 194(1–2):246– 247, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schneider:1998:NOI

[926] H. J. Schneider. A note on outward and inward productions in the categorical graph-grammar approach and Deltagrammars. Theoretical Computer Science, 194(1–2):247–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seese:1998:LTC

[927] D. Seese. Linear time computable problems and logical descriptions. Theoretical Computer Science, 194(1–2):247– ??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shand:1998:PG

[928] D. Shand and S. Brock. Proofs as graphs. Theoretical Computer Science, 194(1–2):247–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sleep:1998:STP

[929] R. Sleep. SEMAGRAPH: the theory and practice of term graph rewriting. Theoretical Computer Science, 194(1– 2):248–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Taentzer:1998:DAS

[930] G. Taentzer and A. Schurr. DIEGO, another step towards a module concept for graph transformation systems. Theoretical Computer Science, 194(1–2): 248–??, March 10, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wadsworth:1998:GRR

[931] C. Wadsworth. Graph reduction: a retrospective. Theoretical Computer Science, 194(1–2):248–??, March 10, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Penczek:1998:F

[932] Wojciech Penczek. Foreword. Theoretical Computer Science, 195(1):1, March 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        195/1/2691.pdf.

deBruin:1998:LDK

[933] Arie de Bruin and Shan-Hwei Nienhuys-Cheng. Linear dynamic Kahn networks are deterministic. Theoretical Computer Science, 195(1):3–32, March 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        195/1/2692.pdf.

Demri:1998:CDI

[934] Stéphane Demri. A class of decidable information logics. Theoretical Computer Science, 195(1):33–60, March 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
195/1/2693.pdf.

Esik:1998:EPI

[935] Z. Ésik and A. Labella. Equational properties of iteration in algebraically complete categories. Theoretical Computer Science, 195(1):61–89, March 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
195/1/2694.pdf.

Seebold:1998:CSM

[936] Patrice Séébold. On the conjugation of standard morphisms. Theoretical Computer Science, 195(1):91– 109, March 20, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/195/1/2695.pdf.

Stirling:1998:DBE

[937] C. Stirling. Decidability of bisimulation equivalence for normed pushdown processes. Theoretical Computer Science, 195(2):113–131, March 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bradfield:1998:MMC

[938] J. C. Bradfield. The modal mu-calculus alternation hierarchy is strict. Theoretical Computer Science, 195(2):133– 153, March 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pitts:1998:PCB

[939] A. M. Pitts and J. R. X. Ross. Process calculus based upon evaluation to committed form. Theoretical Computer Science, 195(2):155–182, March 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Peled:1998:AAC

[940] D. Peled, T. Wilke, and P. Wolper. An algorithmic approach for checking closure properties of temporal logic specifications and ω-regular languages. Theoretical Computer Science, 195(2):183– 203, March 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boreale:1998:EIM

[941] M. Boreale. On the expressiveness of internal mobility in name-passing calculi. Theoretical Computer Science, 195(2):205–226, March 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cleaveland:1998:PAD

[942] R. Cleaveland, G. Luttgen, and V. Natarajan. A process algebra with distributed priorities. Theoretical Computer Science, 195(2):227–258, March 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Philippou:1998:TCO

[943] A. Philippou and D. Walker. On transformations of concurrent-object programs. Theoretical Computer Science, 195(2):259–289, March 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amadio:1998:BAP

[944] R. M. Amadio, I. Castellani, and D. Sangiorgi. On bisimulations for the asynchronous pi-calculus. Theoretical Computer Science, 195(2):291– 324, March 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouge:1998:E

[945] Luc Bougé and Pierre Fraigniaud. Editorial. Theoretical Computer Science, 196(1–2):1, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/196/1-2/2728.pdf.

Gibbons:1998:QRQ

[946] P. B. Gibbons, Y. Matias, and V. Ramachandran. The queue-read queuewrite asynchronous PRAM model. Theoretical Computer Science, 196(1– 2):3–29, April 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/196/1-2/2729.pdf.

Colombet:1998:SEL

[947] L. Colombet and L. Desbat. Speedup and efficiency of large-size applications on heterogeneous networks. Theoretical Computer Science, 196(1–2):31– 44, April 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/196/1-2/2730.pdf.

Kshemkalyani:1998:FVA

[948] Ajay D. Kshemkalyani. A framework for viewing atomic events in distributed computations. Theoretical Computer Science, 196(1–2):45– 70, April 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/196/1-2/2731.pdf.

Botorog:1998:EHL

[949] George Horaţiu Botorog and Herbert Kuchen. Efficient high-level parallel programming. Theoretical Computer Science, 196(1–2):71–107, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
196/1-2/2732.pdf.

Tiskin:1998:BSP

[950] Alexandre Tiskin. The bulksynchronous parallel random access machine. Theoretical Computer Science, 196(1–2):109–130, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
196/1-2/2733.pdf.

DiIanni:1998:EDR

[951] Miriam Di Ianni. Efficient delay routing. Theoretical Computer Science, 196 (1–2):131–151, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ianni:1998:EDR

[952] Miriam Di Ianni. Efficient delay routing. Theoretical Computer Science, 196 (1–2):131–151, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/196/1-2/2734.pdf.

MacKenzie:1998:EPO

[953] Philip D. MacKenzie and Vijaya Ramachandran. ERCW PRAMs and optical communication. Theoretical Computer Science, 196(1–2):153–180, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
196/1-2/2735.pdf.

MeyeraufderHeide:1998:RNO

[954] Friedhelm Meyer auf der Heide, Klaus Schröder, and Frank Schwarze. Routing on networks of optical crossbars. Theoretical Computer Science, 196(1– 2):181–200, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/196/1-2/2736.pdf.

Oberman:1998:RML

[955] Stuart F. Oberman and Michael J. Flynn. Reducing the mean latency of floating-point addition. Theoretical Computer Science, 196(1–2):201– 214, April 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/196/1-2/2737.pdf.

Podehl:1998:SMI

[956] Axel Podehl, Thomas Rauber, and Gudula Rünger. A shared-memory implementation of the hierarchical radiosity method. Theoretical Computer Science, 196(1–2):215–240, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        196/1-2/2738.pdf.

Hollingsworth:1998:UCC

[957] Jeffrey K. Hollingsworth and Barton P. Miller. Using cost to control instrumentation overhead. Theoretical Computer Science, 196(1–2):241–258, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        196/1-2/2739.pdf.

Cho:1998:TGP

[958] Cheng-Hong Cho and Jer-Tsang Wang. Triangular grid protocol: An efficient scheme for replica control with uniform access quorums. Theoretical Computer Science, 196(1–2):259–288, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        196/1-2/2740.pdf.

Hackstadt:1998:DDA

[959] Steven T. Hackstadt and Allen D. Malony. DAQV: Distributed Array Query and Visualization Framework. Theoretical Computer Science, 196(1–2): 289–317, April 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/196/1-2/2741.pdf.

Smith:1998:PFT

[960] J. A. Smith and S. K. Shrivastava. Performance of fault-tolerant data and compute intensive programs over a network of workstations. Theoretical Computer Science, 196(1–2):319– 345, April 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/196/1-2/2742.pdf.

Gasperoni:1998:LSP

[961] Franco Gasperoni and Uwe Schwiegelshohn. List scheduling in the presence of branches. A theoretical evaluation. Theoretical Computer Science, 196(1– 2):347–363, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/196/1-2/2743.pdf.

Knoop:1998:EPD

[962] Jens Knoop. Eliminating partially dead code in explicitly parallel programs. Theoretical Computer Science, 196(1– 2):365–393, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/196/1-2/2744.pdf.

LeGoueslierdArgence:1998:ASB

[963] Patrick Le Gouëslier d’Argence. Affine scheduling on bounded convex polyhedric domains is asymptotically optimal. Theoretical Computer Science, 196(1–2):395–415, April 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
196/1-2/2745.pdf.

Anonymous:1998:LLP

[964] Anonymous. Linear Logic ’96 (papers in summary form only received). Theoretical Computer Science, 197(1– 2):??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mateescu:1998:STS

[965] Alexandru Mateescu, Grzegorz Rozenberg, and Arto Salomaa. Shuffle on trajectories: Syntactic constraints. Theoretical Computer Science, 197(1–2): 1–56, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/197/1-2/2699.pdf.

Nakano:1998:ISA

[966] Koji Nakano and Koichi Wada. Integer summing algorithms on reconfigurable meshes. Theoretical Computer Science, 197(1–2):57–77, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        197/1-2/2543.pdf.

Zhong:1998:RES

[967] Ning Zhong. Recursively enumerable subsets of R^{q} in two computing models: Blum-Shub-Smale machine and Turing machine. Theoretical Computer Science, 197(1–2):79–94, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        197/1-2/2544.pdf.

Albers:1998:CAL

[968] Susanne Albers. A competitive analysis of the list update problem with lookahead. Theoretical Computer Science, 197(1–2):95–109, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
197/1-2/2562.pdf.

Sarkar:1998:MAP

[969] Palash Sarkar and Rana Barua. Multidimensional σ-automata, πpolynomials and generalised S-matrices. Theoretical Computer Science, 197(1– 2):111–138, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/197/1-2/2696.pdf.

Fortnow:1998:RSL

[970] Lance Fortnow, Rūsin¸š Freivalds, William I. Gasarch, Martin Kummer, Stuart A. Kurtz, Carl H. Smith, and Frank Stephan. On the relative sizes of learnable sets. Theoretical Computer Science, 197(1–2):139–156, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
197/1-2/2711.pdf.

Xue:1998:THT

[971] Guoliang Xue. An O(n) time hierarchical tree algorithm for computing force field in n-body simulations. Theoretical Computer Science, 197(1–2): 157–169, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/197/1-2/2768.pdf.

DePrisco:1998:TRV

[972] Roberto De Prisco, Angelo Monti, and Linda Pagli. Testing and reconfiguration of VLSI linear arrays. Theoretical Computer Science, 197(1–2): 171–188, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Prisco:1998:TRV

[973] Roberto De Prisco, Angelo Monti, and Linda Pagli. Testing and reconfiguration of VLSI linear arrays. Theoretical Computer Science, 197(1–2): 171–188, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/197/1-2/2774.pdf.

Larmore:1998:AOS

[974] Lawrence L. Larmore and Wojciech Rytter. Almost optimal sublinear time parallel recognition algorithms for three subclasses of context free languages. Theoretical Computer Science, 197(1–2):189–201, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        197/1-2/2789.pdf.

Le:1998:DOS

[975] Ngoc-Minh Lê. On determining optimal strategies in pursuit games in the plane. Theoretical Computer Science, 197(1–2):203–234, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
197/1-2/2765.pdf.

Tomescu:1998:WCA

[976] Ioan Tomescu. On words containing all short subwords. Theoretical Computer Science, 197(1–2):235–240, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
197/1-2/2767.pdf.

Abramsky:1998:LSS

[977] S. Abramsky and G. McCusker. Linearity, sharing and state: a fully abstract game semantics for idealized Algol with active expressions. Theoretical Computer Science, 197(1–2):241– ??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bierman:1998:TCL

[978] G. M. Bierman. Towards a classical linear lambda-calculus. Theoretical Computer Science, 197(1–2):242– ??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blute:1998:NFC

[979] R. F. Blute and P. J. Scott. A noncommutative full completeness theorem. Theoretical Computer Science, 197(1– 2):242–??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danos:1998:RIO

[980] V. Danos and L. Regnier. Reversible, irreversible and optimal lambdamachines. Theoretical Computer Science, 197(1–2):242–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanDalen:1998:ICB

[981] D. van Dalen. Intuitionism-counting its blessings. Theoretical Computer Science, 197(1–2):242–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fouquere:1998:LLT

[982] C. Fouquere and J. Vauzeilles. Linear logic for taxonomical networks and database updates. Theoretical Computer Science, 197(1–2):243–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girard:1998:DC

[983] J.-Y. Girard. On denotational completeness. Theoretical Computer Science, 197(1–2):243–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girard:1998:CBS

[984] J.-Y. Girard. Coherent Banach spaces: a continuous denotational semantics. Theoretical Computer Science, 197(1– 2):244–??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hayashi:1998:TEP

[985] S. Hayashi, M. Ishikawa, S. Kobayashi, H. Nakano, and S. Nakazaki.

Two extensions of PX system. Theoretical Computer Science, 197(1–2): 244–??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honda:1998:APS

[986] K. Honda. Abstract process structures. Theoretical Computer Science, 197(1– 2):244–??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kanovitch:1998:SCS

[987] M. Kanovitch. Simulating computations in second-order non-commutative linear logic. Theoretical Computer Science, 197(1–2):245–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lamarche:1998:PNG

[988] F. Lamarche. From proof nets to games. Theoretical Computer Science, 197(1–2):245–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lincoln:1998:CLP

[989] P. D. Lincoln, J. C. Mitchell, and A. Scedrov. The complexity of local proof search in linear logic. Theoretical Computer Science, 197(1–2): 245–??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Metayer:1998:SRC

[990] F. Metayer. Some remarks on cyclic linear logic. Theoretical Computer Science, 197(1–2):245–246, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McDowell:1998:ETS

[991] R. McDowell, D. Miller, and C. Palamidessi. Encoding transition systems in sequent calculus. Theoretical Computer Science, 197(1–2):246– ??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagaymam:1998:GTC

[992] M. Nagaymam and M. Okada. A graph-theoretic characterization theorem for multiplicative fragment of noncommutative linear logic. Theoretical Computer Science, 197(1–2):246– ??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okada:1998:PSH

[993] M. Okada. Phase semantics for highorder completeness, cut-elimination and normalization proofs. Theoretical Computer Science, 197(1–2):246– ??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Danos:1998:CIC

[994] V. Danos, J.-B. Joinet, and H. Schellinx. Computational isomorphisms in classical logic. Theoretical Computer Science, 197(1–2):247–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pratt:1998:BDS

[995] V. Pratt. Broadening the denotational semantics of linear logic. Theoretical Computer Science, 197(1–2): 247–??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Retore:1998:PMS

[996] C. Retore. Perfect matching and seriesparallel graphs: multiplicatives proof nets as R and B-graphs. Theoretical Computer Science, 197(1–2):247– ??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hodas:1998:FLP

[997] J. S. Hodas and J. Polakow. Forum as a logic programming language. Theoretical Computer Science, 197(1–2): 247–248, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pedicini:1998:REL

[998] M. Pedicini. Remarks on elementary linear logic. Theoretical Computer Science, 197(1–2):248–??, May 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

TortoradeFalco:1998:GSL

[999] L. Tortora de Falco. Generalized standardization lemma for the additives. Theoretical Computer Science, 197(1– 2):248–??, May 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Otto:1998:EUW

[1000] Friedrich Otto, Paliath Narendran, and Daniel J. Dougherty. Equational unification and word unification, and 2ndorder equational unification. Theoretical Computer Science, 198(1–2): 1–47, May 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/198/1-2/2639.pdf.

Nadathur:1998:NLT

[1001] Gopalan Nadathur and Debra Sue Wilson. A notation for lambda terms. A generalization of environments. Theoretical Computer Science, 198(1–2): 49–98, May 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/198/1-2/2720.pdf.

Geffert:1998:CHP

[1002] Viliam Geffert. A communication hierarchy of parallel computations. Theoretical Computer Science, 198(1–2): 99–130, May 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/198/1-2/2569.pdf.

Hartonas:1998:FAF

[1003] Chrysafis Hartonas. A fixpoint approach to finite delay and fairness. Theoretical Computer Science, 198(1–2): 131–158, May 30, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/198/1-2/2612.pdf.

Boreale:1998:SCP

[1004] Michele Boreale and Davide Sangiorgi. Some congruence properties for πcalculus bisimilarities. Theoretical Computer Science, 198(1–2):159–176, May 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
198/1-2/2661.pdf.

Goldblatt:1998:EFA

[1005] Robert Goldblatt. Enlargements of functional algebras for the lambda calculus. Theoretical Computer Science, 198(1–2):177–200, May 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
198/1-2/2662.pdf.

Egly:1998:AOP

[1006] Uwe Egly. An answer to an open problem of Urquhart. Theoretical Computer Science, 198(1–2):201–209, May 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
198/1-2/2676.pdf.

Esparza:1998:RLS

[1007] Javier Esparza. Reachability in live and safe free-choice Petri nets is NP-complete. Theoretical Computer Science, 198(1–2):211–224, May 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
198/1-2/2771.pdf.

Corradini:1998:CCW

[1008] Flavio Corradini. On the coarsest congruence within global-clock-bounded equivalence. Theoretical Computer Science, 198(1–2):225–237, May 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
198/1-2/2773.pdf.

Cagman:1998:CWR

[1009] Naim Çaǧman and J. Roger Hindley. Combinatory weak reduction in lambda calculus. Theoretical Computer Science, 198(1–2):239–247, May 30, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        198/1-2/2786.pdf.

Nijholt:1998:E

[1010] A. Nijholt and G. Scollo. Editorial. Theoretical Computer Science, 199(1– 2):1–3, June 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/199/1-2/2800.pdf.

Buszkowski:1998:ASC

[1011] Wojciech Buszkowski. Algebraic structures in categorial grammar. Theoretical Computer Science, 199(1–2): 5–24, June 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/199/1-2/2801.pdf.

Janssen:1998:ATC

[1012] Theo M. V. Janssen. Algebraic translations, correctness and algebraic compiler construction. Theoretical Computer Science, 199(1–2):25–56, June 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        199/1-2/2802.pdf.

Visser:1998:PSD

[1013] Eelco Visser. Polymorphic syntax definition. Theoretical Computer Science, 199(1–2):57–86, June 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
199/1-2/2803.pdf.

Sikkel:1998:PSC

[1014] Klaas Sikkel. Parsing schemata and correctness of parsing algorithms. Theoretical Computer Science, 199(1–2): 87–103, June 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/199/1-2/2804.pdf.

Rus:1998:APP

[1015] Teodor Rus. Algebraic processing of programming languages. Theoretical Computer Science, 199(1–2):105– 143, June 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/199/1-2/2805.pdf.

Tendeau:1998:CAD

[1016] Frédéric Tendeau. Computing abstract decorations of parse forests using dynamic programming and algebraic power series. Theoretical Computer Science, 199(1–2):145–166, June 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
199/1-2/2806.pdf.

delaClergerie:1998:IFT

[1017] Eric Villemonte de la Clergerie and François Barthélemy. Information flow in tabular interpretations for generalized push-down automata. Theoretical Computer Science, 199(1–2):167– 198, June 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/199/1-2/2807.pdf.

Rus:1998:PPM

[1018] Teodor Rus and James S. Jones. PHRASE parsers from multi-axiom grammars. Theoretical Computer Science, 199(1–2):199–229, June 15, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
199/1-2/2808.pdf.

Ramalho:1998:ASD

[1019] José Carlos Ramalho, José João Almeida, and Pedro Henriques. Algebraic specification of documents. Theoretical Computer Science, 199(1–2): 231–247, June 15, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/199/1-2/2809.pdf.

Barthelmann:1998:NOF

[1020] Klaus Barthelmann. Nondeterministic operations on finite relational structures. Theoretical Computer Science, 200(1–2):1–44, June 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/200/1-2/2799.pdf.

Schmitt:1998:STA

[1021] Vincent Schmitt. Stable trace automata vs. full trace automata. Theoretical Computer Science, 200(1–2): 45–100, June 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/200/1-2/2832.pdf.

Schmitt:1998:LAG

[1022] Lothar M. Schmitt, Chrystopher L. Nehaniv, and Robert H. Fujii. Linear analysis of genetic algorithms. Theoretical Computer Science, 200(1–2): 101–134, June 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/200/1-2/2839.pdf.

Zielonka:1998:IGF

[1023] Wieslaw Zielonka. Infinite games on finitely coloured graphs with applications to automata on infinite trees. Theoretical Computer Science, 200(1– 2):135–183, June 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/200/1-2/2844.pdf.

Jeavons:1998:ASC

[1024] Peter Jeavons. On the algebraic structure of combinatorial problems. Theoretical Computer Science, 200(1–2): 185–204, June 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/200/1-2/2766.pdf.

Harju:1998:QOW

[1025] Tero Harju and Lucian Ilie. On quasi orders of words and the confluence property. Theoretical Computer Science, 200(1–2):205–224, June 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
200/1-2/2795.pdf.

Hennessy:1998:BCB

[1026] M. Hennessy and J. Rathke. Bisimulations for a calculus of broadcasting systems. Theoretical Computer Science, 200(1–2):225–260, June 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
200/1-2/2797.pdf.

Rivals:1998:ORA

[1027] E. Rivals and J.-P. Delahaye. Optimal representation in average using Kolmogorov complexity. Theoretical Computer Science, 200(1–2):261– 287, June 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/200/1-2/2810.pdf.

vandePol:1998:OSR

[1028] Jaco van de Pol. Operational semantics of rewriting with priorities. Theoretical Computer Science, 200(1–2): 289–312, June 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/200/1-2/2818.pdf.

Blundo:1998:GBC

[1029] C. Blundo, Luiz A. Frota Mattos, and D. R. Stinson. Generalized Beimel– Chor schemes for broadcast encryption and interactive key distribution. Theoretical Computer Science, 200(1–2): 313–334, June 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/200/1-2/2837.pdf.

Mundici:1998:RMB

[1030] Daniele Mundici and Nicola Olivetti.

Resolution and model building in the infinite-valued calculus of Lukasiewicz. Theoretical Computer Science, 200(1– 2):335–366, June 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/200/1-2/2847.pdf.

Jacquet:1998:ADA

[1031] Philippe Jacquet and Wojciech Szpankowski. Analytical depoissonization and its applications. Theoretical Computer Science, 201(1–2):1–62, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
201/1-2/2703.pdf.

Cichon:1998:ORB

[1032] E. A. Cichon and E. Tahhan Bittar. Ordinal recursive bounds for Higman’s theorem. Theoretical Computer Science, 201(1–2):63–84, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cas/tree/store/tcs/sub/1998/
                                        201/1-2/2545.pdf.

Lam:1998:ISS

[1033] Tak Wah Lam and Ka Hing Lee. An improved scheme for set equality testing and updating. Theoretical Computer Science, 201(1–2):85–97, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        201/1-2/2560.pdf.

Moore:1998:DRR

[1034] Cristopher Moore. Dynamical recognizers: real-time language recognition by analog computers. Theoretical Computer Science, 201(1–2):99– 136, July 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/201/1-2/2564.pdf.

Portier:1998:RUP

[1035] Natacha Portier. Résolutions universelles pour des problèmes NPcomplets. (French) [Universal resolution of NP-complete problems]. Theoretical Computer Science, 201(1–2): 137–150, July 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/201/1-2/2565.pdf.

Rosaz:1998:IUL

[1036] Laurent Rosaz. Inventories of unavoidable languages and the wordextension conjecture. Theoretical Computer Science, 201(1–2):151–170, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
201/1-2/2567.pdf.

Cattaneo:1998:GSS

[1037] Gianpiero Cattaneo and Luciano Margara. Generalized sub-shifts in elementary cellular automata: the “strange case” of chaotic rule 180. Theoretical Computer Science, 201(1–2):171– 187, July 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/201/1-2/2746.pdf.

Flasinski:1998:PPN

[1038] M. Flasiński. Power properties of NLC graph grammars with a polynomial membership problem. Theoretical Computer Science, 201(1–2):189– 231, July 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/201/1-2/2748.pdf.

Wada:1998:EAM

[1039] Koichi Wada, Akinari Takaki, and Kimio Kawaguchi. Efficient algorithms for a mixed k-partition problem of graphs without specifying bases. Theoretical Computer Science, 201(1–2): 233–248, July 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/201/1-2/2761.pdf.

Ferragina:1998:UST

[1040] Paolo Ferragina, Roberto Grossi, and Manuela Montangero. On updating suffix tree labels. Theoretical Computer Science, 201(1–2):249–262, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        201/1-2/2779.pdf.

Park:1998:ATD

[1041] Kunsoo Park. Analysis of twodimensional approximate pattern matching algorithms. Theoretical Computer Science, 201(1–2):263–273, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        201/1-2/2812.pdf.

Chung:1998:IAS

[1042] Kuo-Liang Chung. An improved algorithm for solving the banded cyclic string-to-string correction problem. Theoretical Computer Science, 201(1– 2):275–279, July 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/201/1-2/2815.pdf.

Diaz:1998:RGC

[1043] J. Diaz, M. Serna, and P. Spirakis. On the random generation and counting of matchings in dense graphs. Theoretical Computer Science, 201(1–2): 281–290, July 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/201/1-2/2830.pdf.

Bernardo:1998:TET

[1044] Marco Bernardo and Roberto Gorrieri.

A tutorial on EMPA: A theory of concurrent processes with nondeterminism, priorities, probabilities and time. Theoretical Computer Science, 202(1– 2):1–54, July 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/202/1-2/2663.pdf.  See

corrigendum [1425].

Brzoska:1998:PMT

[1045] Christoph Brzoska. Programming in metric temporal logic. Theoretical Computer Science, 202(1–2):55–125, July 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
202/1-2/2675.pdf.

Felder:1998:TIR

[1046] Miguel Felder, Angelo Gargantini, and Angelo Morzenti. A theory of implementation and refinement in timed Petri nets. Theoretical Computer Science, 202(1–2):127–161, July 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
202/1-2/2614.pdf.

Cortesi:1998:QAI

[1047] Agostino Cortesi, Gilberto Filé, and William Winsborough. The quotient of an abstract interpretation. Theoretical Computer Science, 202(1–2): 163–192, July 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/202/1-2/2673.pdf.

Hartonas:1998:DML

[1048] Chrysafis Hartonas. Duality for modal μ-logics. Theoretical Computer Science, 202(1–2):193–222, July 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        202/1-2/2769.pdf.

vanBreugel:1998:TMS

[1049] Franck van Breugel. Terminal metric spaces of finitely branching and image finite linear processes. Theoretical Computer Science, 202(1–2):223– 230, July 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/202/1-2/2787.pdf.

Ramos-Jimenez:1998:CPC

[1050] G. Ramos-Jiménez, J. López-Muñoz, and R. Morales-Bueno. Comparisons of Parikh’s condition to other conditions for context-free languages. Theoretical Computer Science, 202(1–2): 231–244, July 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/202/1-2/2798.pdf.

Ausiello:1998:F

[1051] Giorgio Ausiello and Alberto Marcjetto-Spacemela. Foreword. Theoretical Computer Science, 203(1):1, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
203/1/2819.pdf.

Diks:1998:MGP

[1052] Krzysztof Diks and Torben Hagerup.

More general parallel tree contraction: Register allocation and broadcasting in a tree. Theoretical Computer Science, 203(1):3–29, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/203/1/2820.pdf.

Hartmann:1998:SRV

[1053] Stephan Hartmann, Markus W.

Schaffter, and Andreas S. Schulz. Switchbox routing in VLSI design: Closing the complexity gap. Theoretical Computer Science, 203(1):31– 49, August 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/203/1/2821.pdf.

Crescenzi:1998:SUD

[1054] P. Crescenzi and P. Penna. Strictlyupward drawings of ordered search trees. Theoretical Computer Science, 203(1):51–67, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/cas/tree/
store/tcs/sub/1998/203/1/2822.pdf.

Cicerone:1998:UAS

[1055] Serafino Cicerone, Daniele Frigioni, Umberto Nanni, and Francesco Pugliese. A uniform approach to semi-dynamic problems on digraphs. Theoretical Computer Science, 203(1):69–90, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
203/1/2823.pdf.

Drangmeister:1998:MEN

[1056] Kay U. Drangmeister, Sven O. Krumke, Madhav V. Marathe, Hartmut Noltemeier, and S. S. Ravi. Modifying edges of a network to obtain short subgraphs. Theoretical Computer Science, 203(1):91–121, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
203/1/2824.pdf.

Gustedt:1998:EUF

[1057] Jens Gustedt. Efficient Union-Find for planar graphs and other sparse graph classes. Theoretical Computer Science, 203(1):123–141, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
203/1/2825.pdf.

Takaoka:1998:SPA

[1058] Tadao Takaoka. Shortest path algorithms for nearly acyclic directed graphs. Theoretical Computer Science, 203(1):143–150, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
203/1/2826.pdf.

Kranakis:1998:AMF

[1059] Evangelos Kranakis, Danny Krizanc, Andrzej Pelc, and David Peleg. Approximate maxima finding of continuous functions under restricted budget. Theoretical Computer Science, 203(1): 151–162, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/203/1/2827.pdf.

Diks:1998:SDS

[1060] Krzysztof Diks and Andrzej Pelc. System diagnosis with smallest risk of error. Theoretical Computer Science, 203 (1):163–173, August 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/203/1/2828.pdf.

Baumker:1998:TEP

[1061] Armin Baumker, Wolfgang Dittrich, and Friedhelm Meyer auf der Heide. Truly efficient parallel algorithms: 1optimal multisearch for an extension of the BSP model. Theoretical Computer Science, 203(2):175–203, August 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
203/2/2851.pdf.

Chaudhuri:1998:SPD

[1062] Shiva Chaudhuri and Christos D. Zaroliagis. Shortest paths in digraphs of small treewidth. part II: Optimal parallel algorithms. Theoretical Computer Science, 203(2):205–223, August 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        203/2/2852.pdf.

Dubhashi:1998:NOD

[1063] Devdatt Dubhashi, David A. Grable, and Alessandro Panconesi. Nearoptimal, distributed edge colouring via the nibble method. Theoretical Computer Science, 203(2):225–251, August 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        203/2/2853.pdf.

Even:1998:MCC

[1064] Shimon Even, Gene Itkis, and Sergio Rajsbaum. On mixed connectivity certificates. Theoretical Computer Science, 203(2):253–269, August 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        203/2/2854.pdf.

Bruyere:1998:MCB

[1065] Véronique Bruyère. On maximal codes with bounded synchronization delay. Theoretical Computer Science, 204(1– 2):11–28, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/204/1-2/2859.pdf.

Cassaigne:1998:EUP

[1066] Julien Cassaigne and Juhani Karhumäki. Examples of undecidable problems for 2-generator matrix semigroups. Theoretical Computer Science, 204(1–2):29– 34, September 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/204/1-2/2860.pdf.

Choffrut:1998:CFI

[1067] Christian Choffrut and Flavio D’Alessandro.

Commutativity in free inverse monoids. Theoretical Computer Science, 204(1– 2):35–54, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/204/1-2/2861.pdf.

Cori:1998:CNI

[1068] Robert Cori and Michel Marcus.

Counting non-isomorphic chord diagrams. Theoretical Computer Science, 204(1–2):55–73, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
204/1-2/2862.pdf.

deLuca:1998:CCF

[1069] Aldo de Luca. A conjecture on continued fractions. Theoretical Computer Science, 204(1–2):75–86, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
204/1-2/2863.pdf.

Duval:1998:PLP

[1070] Jean-Pierre Duval. Périodes locales et propagation de périodes dans un mot. (French) [local periods and period propagation in a word]. Theoretical Computer Science, 204(1–2):87– 98, September 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/204/1-2/2864.pdf.

Goralcik:1998:DSP

[1071] P. Goralcik and V. Koubek. On the disjunctive set problem. Theoretical Computer Science, 204(1–2):99–118, September 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/204/1-2/2865.pdf.

Hansel:1998:SNI

[1072] Georges Hansel. Systèmes de numération independants et syndéticite. (French) [Systems of independent ???? numeration]. Theoretical Computer Science, 204(1–2):119–130, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        204/1-2/2866.pdf.

Ilie:1998:WQO

[1073] Lucian Ilie and Arto Salomaa. On well quasi orders of free monoids. Theoretical Computer Science, 204(1–2):131– 152, September 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/204/1-2/2867.pdf.

Mignosi:1998:PGR

[1074] Filippo Mignosi, Antonio Restivo, and Sergio Salemi. Periodicity and the golden ratio. Theoretical Computer Science, 204(1–2):153–167, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
204/1-2/2868.pdf.

McNaughton:1998:FFP

[1075] Robert McNaughton. The finiteness of finitely presented monoids. Theoretical Computer Science, 204(1–2):169– 182, September 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/204/1-2/2869.pdf.

Paun:1998:SS

[1076] Gheorghe Păun and Grzegorz Rozenberg. Sticker systems. Theoretical Computer Science, 204(1–2):183–203, September 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/204/1-2/2870.pdf.

Sakarovitch:1998:CFA

[1077] Jacques Sakarovitch. A construction on finite automata that has remained hidden. Theoretical Computer Science, 204(1–2):205–231, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
204/1-2/2871.pdf.

Schupp:1998:SHC

[1078] Paul E. Schupp. On the structure of Hamiltonian cycles in Cayley graphs of finite quotients of the modular group. Theoretical Computer Science, 204(1–2):233–248, September 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        204/1-2/2872.pdf.

Steinby:1998:GVT

[1079] Magnus Steinby. General varieties of tree languages. Theoretical Computer Science, 205(1–2):1–43, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        205/1-2/2845.pdf.

Bollig:1998:HTO

[1080] Beate Bollig, Martin Sauerhoff, Detlef Sieling, and Ingo Wegener. Hierarchy theorems for k OBDDs and k IBDDs. Theoretical Computer Science, 205(1– 2):45–60, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/205/1-2/2570.pdf.

Ehrenfeucht:1998:RRE

[1081] Andrzej Ehrenfeucht, Gheorghe Păun, and Grzegorz Rozenberg. On representing recursively enumerable languages by internal contextual languages. Theoretical Computer Science, 205(1–2):61– 83, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/205/1-2/2571.pdf.

Yu:1998:CCS

[1082] Chang-Wu Yu, Gen-Huey Chen, and Tze-Heng Ma. On the complexity of the k-chain subgraph cover problem. Theoretical Computer Science, 205(1–2):85– 98, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/205/1-2/2572.pdf.

Golin:1998:LTP

[1083] M. Golin and S. Zaks. Labelled trees and pairs of input–output permutations in priority queues. Theoretical Computer Science, 205(1–2):99–114, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/205/1-2/2573.pdf.

Flammini:1998:MIR

[1084] Michele Flammini, Giorgio Gambosi, Umberto Nanni, and Richard B. Tan. Multidimensional interval routing schemes. Theoretical Computer Science, 205(1–2):115–133, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
205/1-2/2606.pdf.

Sato:1998:ECL

[1085] Tadakazu Sato. Ergodic characterization of linear cellular automata over Z_{m}. Theoretical Computer Science, 205(1–2):135–144, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
205/1-2/2607.pdf.

Giammarresi:1998:PT

[1086] Dora Giammarresi, Sabrina Mantaci, Filippo Mignosi, and Antonio Restivo. Periodicities on trees. Theoretical Computer Science, 205(1–2):145–181, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/205/1-2/2608.pdf.

Durand-Lose:1998:PTT

[1087] J. O. Durand-Lose. Parallel transient time of one-dimensional sand pile. Theoretical Computer Science, 205(1–2): 183–193, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/205/1-2/2609.pdf.

Martin-Vide:1998:CRE

[1088] Carlos Martin-Vide, Gheorghe Păun, and Arto Salomaa. Characterizations of recursively enumerable languages by means of insertion grammars. Theoretical Computer Science, 205(1–2):195– 205, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/205/1-2/2615.pdf.

Andre:1998:EPL

[1089] Yves Andre and Francis Bossut. On the equivalence problem for letter-to-letter top-down tree transducers. Theoretical Computer Science, 205(1–2):207– 229, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/205/1-2/2616.pdf.

Wang:1998:SCP

[1090] Biing-Feng Wang. Simulating the CRCW PRAM on reconfigurable networks. Theoretical Computer Science, 205(1–2):231–242, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
205/1-2/2632.pdf.

Karpinski:1998:AIO

[1091] Marek Karpinski and Wojciech Rytter. Alphabet-independent optimal parallel search for three-dimensional patterns. Theoretical Computer Science, 205(1–2):243–260, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
205/1-2/2788.pdf.

Andreev:1998:PCA

[1092] A. E. Andreev, A. Clementi, P. Crescenzi, E. Dahlhaus, S. De Agostino, and J. D. P. Rolim. The parallel complexity of approximating the high degree subgraph problem. Theoretical Computer Science, 205(1–2):261–282, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/205/1-2/2811.pdf.

Fraenkel:1998:AWG

[1093] Aviezri S. Fraenkel and Michal Ozery. Adjoining to Wythoff’s game its P-positions as moves. Theoretical Computer Science, 205(1–2):283–296, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/205/1-2/2875.pdf.

Beal:1998:BSD

[1094] Marie-Pierre Béal and Jean Senellart. On the bound of the synchronization delay of a local automaton. Theoretical Computer Science, 205(1–2):297– 306, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/205/1-2/2834.pdf.

Beling:1998:UFM

[1095] Peter A. Beling and Nimrod Megiddo. Using fast matrix multiplication to find basic solutions. Theoretical Computer Science, 205(1–2):307–316, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        205/1-2/2838.pdf.

Hemaspaandra:1998:BOJ

[1096] Lane A. Hemaspaandra, Zhigen Jiang, Jörg Rothe, and Osamu Watanabe. Boolean operations, joins, and the extended low hierarchy. Theoretical Computer Science, 205(1–2):317–327, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/205/1-2/2841.pdf.

Wang:1998:RSR

[1097] Huaxiong Wang. On rational series and rational languages. Theoretical Computer Science, 205(1–2):329–336, September 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/205/1-2/2933.pdf.

Chuan:1998:UFC

[1098] Wai fong Chuan. Unbordered factors of the characteristic sequences of irrational numbers. Theoretical Computer Science, 205(1–2):337–344, September 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
205/1-2/2934.pdf.

Duggan:1998:UEP

[1099] Dominic Duggan. Unification with extended patterns. Theoretical Computer Science, 206(1–2):1–50, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
206/1-2/2677.pdf.

Etalle:1998:SMG

[1100] Sandro Etalle. A semantics for modular general logic programs. Theoretical Computer Science, 206(1–2):51– 80, October 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/206/1-2/2727.pdf.

Bensaou:1998:TCL

[1101] N. Bensaou and I. Guessarian. Transforming constraint logic programs. Theoretical Computer Science, 206(1– 2):81–125, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/206/1-2/2613.pdf.

Liu:1998:PCP

[1102] Xinxin Liu and David Walker. Partial confluence of processes and systems of objects. Theoretical Computer Science, 206(1–2):127–162, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        206/1-2/2647.pdf.

Dehornoy:1998:TPC

[1103] Patrick Dehornoy and Abderrahim Marzouk. Theorem proving by chain resolution. Theoretical Computer Science, 206(1–2):163–180, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        206/1-2/2664.pdf.

Eiter:1998:EPC

[1104] Thomas Eiter, Nicola Leone, and Domenico Saccá. Expressive power and complexity of partial models for disjunctive deductive databases. Theoretical Computer Science, 206(1–2):181– 218, October 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/206/1-2/2665.pdf.

Priese:1998:UAT

[1105] Lutz Priese and Harro Wimmel. A uniform approach to true-concurrency and interleaving semantics for Petri nets.

Theoretical Computer Science, 206(1– 2):219–256, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/206/1-2/2705.pdf.

Yamasaki:1998:SAP

[1106] Susumu Yamasaki and Yoshinori Kurose. Soundness of abductive proof procedure with respect to constraint for non-ground abducibles. Theoretical Computer Science, 206(1–2):257– 281, October 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/206/1-2/2707.pdf.

Levene:1998:AFD

[1107] Mark Levene and George Loizou. Axiomatisation of functional dependencies in incomplete relations. Theoretical Computer Science, 206(1–2):283– 300, October 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/206/1-2/2836.pdf.

Colson:1998:SCV

[1108] Loïc Colson and Daniel Fredholm. System T, call-by-value and the minimum problem. Theoretical Computer Science, 206(1–2):301–315, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
206/1-2/2846.pdf.

Loader:1998:UPD

[1109] Ralph Loader. Unary PCF is decidable. Theoretical Computer Science, 206(1– 2):317–329, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/206/1-2/2879.pdf.

Hirokawa:1998:IPP

[1110] Sachio Hirokawa. Infiniteness of proof(α) is polynomial-space complete. Theoretical Computer Science, 206(1– 2):331–339, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/206/1-2/2704.pdf.

Lamport:1998:PPP

[1111] Leslie Lamport. Proving possibility properties. Theoretical Computer Science, 206(1–2):341–352, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        206/1-2/2949.pdf.

Perrier:1998:CGP

[1112] Guy Perrier. Corrigendum to Galmiche’s and Perrier’s “On proof normalization in linear logic” [Theoret. Comput. Sci 135(1) (1994) 67– 110]. Theoretical Computer Science, 206(1–2):353–354, October 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        206/1-2/2965.pdf. See [3].

McNaughton:1998:CRV

[1113] Robert McNaughton. Contributions of Ronald V. Book to the theory of stringrewriting systems. Theoretical Computer Science, 207(1):13–23, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/1/2884.pdf.

Brandenburg:1998:AWG

[1114] Franz J. Brandenburg. The ancestor width of grammars and languages. Theoretical Computer Science, 207(1): 25–41, October 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/1/2885.pdf.

Otto:1998:SUR

[1115] Friedrich Otto. Some undecidability results concerning the property of preserving regularity. Theoretical Computer Science, 207(1):43–72, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/1/2886.pdf.

Salomaa:1998:SEE

[1116] Kai Salomaa and Sheng Yu. Synchronization expressions with extended join operation. Theoretical Computer Science, 207(1):73–88, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/1/2887.pdf.

Baier:1998:BQA

[1117] Herbert Baier and Klaus W. Wagner. Bounding queries in the analytic polynomial-time hierarchy. Theoretical Computer Science, 207(1):89– 104, October 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/1/2888.pdf.

Cai:1998:RPD

[1118] Jin-Yi Cai. A relation of primal– dual lattices and the complexity of shortest lattice vector problem. Theoretical Computer Science, 207(1):105– 116, October 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/1/2889.pdf.

Macarie:1998:PPP

[1119] Ioan I. Macarie and Mitsunori Ogihara. Properties of probabilistic pushdown automata. Theoretical Computer Science, 207(1):117–130, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/1/2890.pdf.

Naik:1998:HBO

[1120] Ashish V. Naik, John D. Rogers, James S. Royer, and Alan L. Selman. A hierarchy based on output multiplicity. Theoretical Computer Science, 207(1): 131–157, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/207/1/2891.pdf.

Vollmer:1998:RPT

[1121] Heribert Vollmer. Relating polynomial time to constant depth. Theoretical Computer Science, 207(1):159– 170, October 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/1/2892.pdf.

Du:1998:ASI

[1122] Xiufeng Du, Weili Wu, and Dean F.

Kelley. Approximations for subset interconnection designs. Theoretical Computer Science, 207(1):171–180, October 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/1/2893.pdf.

Lin:1998:CMP

[1123] Guo-Hui Lin and Guoliang Xue. Kcenter and K-median problems in graded distances. Theoretical Computer Science, 207(1):181–192, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/1/2894.pdf.

Wan:1998:CFC

[1124] Peng-Jun Wan. Conflict-free channel set assignment for an optical cluster interconnection network based on rotator digraphs. Theoretical Computer Science, 207(1):193–201, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/1/2895.pdf.

Wang:1998:OPV

[1125] Jie Wang and Yaorong Ge. An optimization problem in virtual endoscopy. Theoretical Computer Science, 207(1): 203–216, October 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1998/207/1/2896.pdf.

Balcazar:1998:SLA

[1126] José L. Balcázar and Montserrat Hermo. The structure of logarithmic advice complexity classes. Theoretical Computer Science, 207(1):217– 244, October 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/207/1/2897.pdf.

Lorentz:1998:GRF

[1127] Amy K. Lorentz and Jack H. Lutz. Genericity and randomness over feasible probability measures. Theoretical Computer Science, 207(1):245– 259, October 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/207/1/2898.pdf.

Muchnik:1998:MMR

[1128] Andrei A. Muchnik, Alexei L. Semenov, and Vladimir A. Uspensky. Mathematical metaphysics of randomness. Theoretical Computer Science, 207(2):263–317, November 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        207/2/2900.pdf.

Muchnik:1998:CI

[1129] An A. Muchnik. On common information. Theoretical Computer Science, 207(2):319–328, November 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/2/2901.pdf.

Vereshchagin:1998:RBD

[1130] Nikolai K. Vereshchagin. Randomized Boolean decision trees: Several remarks. Theoretical Computer Science, 207(2):329–342, November 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/2/2902.pdf.

Vyugin:1998:ETI

[1131] V. V. V’yugin. Ergodic theorems for individual random sequences. Theoretical Computer Science, 207(2):343– 361, November 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/2/2903.pdf.

Vyugin:1998:NSI

[1132] V. V. V’yugin. Non-stochastic infinite and finite sequences. Theoretical Computer Science, 207(2):363–382, November 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/2/2904.pdf.

Gorbunov:1998:CF

[1133] K. Yu. Gorbunov. On a complexity of the formula (A ∨ B) ⇒ C. Theoretical Computer Science, 207(2):383– 386, November 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/207/2/2905.pdf.

Kolmogorov:1998:TRN

[1134] A. N. Kolmogorov. On tables of random numbers. Theoretical Computer Science, 207(2):387–395, November 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
207/2/2906.pdf.

Madlener:1998:RRT

[1135] Klaus Madlener and Birgit Reinert. Relating rewriting techniques on monoids and rings: congruences on monoids and ideals in monoid rings. Theoretical Computer Science, 208(1–2):3– 31, November 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/208/1-2/2908.pdf.

Jouannaud:1998:ROH

[1136] Jean-Pierre Jouannaud and Albert Rubio. Rewrite orderings for higher-order terms in η-long β-normal form and the recursive path ordering. Theoretical Computer Science, 208(1–2):33– 58, November 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/208/1-2/2909.pdf.

KrishnaRao:1998:MAT

[1137] M. R. K. Krishna Rao. Modular aspects of term graph rewriting. Theoretical Computer Science, 208(1–2):59–??, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rao:1998:MAT

[1138] M. R. K. Krishna Rao. Modular aspects of term graph rewriting. Theoretical Computer Science, 208(1–2):59– 86, November 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/208/1-2/2910.pdf.

Sakai:1998:SSS

[1139] Masahiko Sakai and Yoshihito Toyama.

Semantics and strong sequentiality of priority term rewriting systems. Theoretical Computer Science, 208(1–2):87– 110, November 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/208/1-2/2911.pdf.

Schmidt-Schauss:1998:DAD

[1140] Manfred Schmidt-Schauß. A decision algorithm for distributive unification. Theoretical Computer Science, 208(1–2):111–148, November 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
208/1-2/2912.pdf.

Stuber:1998:STP

[1141] Jürgen Stuber. Superposition theorem proving for abelian groups represented as integer modules. Theoretical Computer Science, 208(1–2):149– 177, November 28, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/208/1-2/2913.pdf.

Treinen:1998:FOT

[1142] Ralf Treinen. The first-order theory of linear one-step rewriting is undecidable. Theoretical Computer Science, 208(1–2):179–190, November 28, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
208/1-2/2914.pdf.

Bodlaender:1998:PAG

[1143] Hans L. Bodlaender. A partial karboretum of graphs with bounded treewidth. Theoretical Computer Science, 209(1–2):1–45, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
209/1-2/2764.pdf.

Allender:1998:NCA

[1144] Eric Allender, Jia Jiao, Meena Mahajan, and V. Vinay. Non-commutative arithmetic circuits: depth reduction and size lower bounds. Theoretical Computer Science, 209(1–2):47–86, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2763.pdf.

Beguin:1998:GID

[1145] Philippe Béguin and Antonella Cresti.

General information dispersal algorithms. Theoretical Computer Science, 209(1–2):87–105, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
209/1-2/2634.pdf.

Demange:1998:DAA

[1146] Marc Demange, Pascal Grisoni, and Vangelis Th. Paschos. Differential approximation algorithms for some combinatorial optimization problems. Theoretical Computer Science, 209(1–2): 107–122, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/209/1-2/2635.pdf.

Downey:1998:TDS

[1147] Rodney G. Downey and Michael R.

Fellows. Threshold dominating sets and an improved characterization of W [2]. Theoretical Computer Science, 209(1–2):123–140, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
209/1-2/2637.pdf.

Apolloni:1998:SSL

[1148] B. Apolloni and C. Gentile. Sample size lower bounds in PAC learning by Algorithmic Complexity Theory. Theoretical Computer Science, 209(1–2):141– 162, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/209/1-2/2638.pdf.

Aceto:1998:QSE

[1149] Luca Aceto, Wan Fokkink, and Anna Ingólfsdóttir. On a question of A. Salomaa: The equational theory of regular expressions over a singleton alphabet is not finitely based. Theoretical Computer Science, 209(1–2):163– 178, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/209/1-2/2640.pdf.

Blanchard:1998:LCR

[1150] François Blanchard and Petr Kůrka. Language complexity of rotations and Sturmian sequences. Theoretical Computer Science, 209(1–2):179–193, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1998/209/1-2/2641.pdf.

AdeleA:1998:CCF

[1151] F. Adele A. Communication complexity of fault-tolerant information diffusion. Theoretical Computer Science, 209(1–2):195–??, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gargano:1998:CCF

[1152] Luisa Gargano and Adele A. Rescigno. Communication complexity of faulttolerant information diffusion. Theoretical Computer Science, 209(1–2):195– 211, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2645.pdf.

Csuhaj-Varju:1998:TBE

[1153] Erzsébet Csuhaj-Varjú and Alica Kelemenová. Team behaviour in ecogrammar systems. Theoretical Computer Science, 209(1–2):213–224, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2646.pdf.

Zimand:1998:SCW

[1154] Marius Zimand. On the size of classes with weak membership properties. Theoretical Computer Science, 209(1–2):225–235, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
209/1-2/2650.pdf.

Amaldi:1998:AMN

[1155] Edoardo Amaldi and Viggo Kann.

On the approximability of minimizing nonzero variables or unsatisfied relations in linear systems. Theoretical Computer Science, 209(1–2):237– 260, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2651.pdf.

Vuillon:1998:CMS

[1156] Laurent Vuillon. Combinatoire des motifs d’une suite sturmienne bidimensionnelle. (French) [Combination of motifs of a Sturmian bidimensional sequence]. Theoretical Computer Science, 209(1–2):261–285, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
209/1-2/2653.pdf.

Rossner:1998:HAS

[1157] Carsten Rössner and Jean-Pierre Seifert. On the hardness of approximating shortest integer relations among rational numbers. Theoretical Computer Science, 209(1–2):287–297, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
209/1-2/2654.pdf.

Beaudry:1998:LRF

[1158] Martin Beaudry. Languages recognized by finite aperiodic groupoids. Theoretical Computer Science, 209(1–2):299– 317, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2655.pdf.

Vaszil:1998:SNR

[1159] György Vaszil. On simulating nonreturning PC grammar systems with returning systems. Theoretical Computer Science, 209(1–2):319–329, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2656.pdf.

Colbourn:1998:LTA

[1160] Charles J. Colbourn and Guoliang Xue. A linear time algorithm for computing the most reliable source on a series–parallel graph with unreliable edges. Theoretical Computer Science, 209(1–2):331–345, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1998/
209/1-2/2660.pdf.

Malesinska:1998:HAF

[1161] Ewa Malesińska and Alessandro Panconesi. On the hardness of allocating frequencies for hybrid networks. Theoretical Computer Science, 209(1–2): 347–363, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1998/209/1-2/2975.pdf.

Breslauer:1998:CLP

[1162] Dany Breslauer. On competitive online paging with lookahead. Theoretical Computer Science, 209(1–2):365– 375, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2940.pdf.

Downarowicz:1998:MFM

[1163] T. Downarowicz and Y. Lacroix. Merit factors and Morse sequences. Theoretical Computer Science, 209(1–2):377– 387, December 06, 1998. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1998/209/1-2/2943.pdf.

Sengupta:1998:LBM

[1164] Rimli Sengupta and H. Venkateswaran. A lower bound for monotone arithmetic circuits computing 0–1 permanent. Theoretical Computer Science, 209(1–2):389–398, December 06, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1998/
                                        209/1-2/2950.pdf.

Brattka:1999:CI

[1165] Vasco Brattka. Computable invariance. Theoretical Computer Science, 210(1): 3–20, January 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/210/1/2925.pdf.

Bournez:1999:ATC

[1166] Olivier Bournez. Achilles and the Tortoise climbing up the hyperarithmetical hierarchy. Theoretical Computer Science, 210(1):21–71, January 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        210/1/2926.pdf.

Edalat:1999:DTA

[1167] Abbas Edalat and Philipp Sünderhauf. A domain-theoretic approach to computability on the real line. Theoretical Computer Science, 210(1):73– 98, January 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/210/1/2927.pdf.

Ho:1999:RRR

[1168] Chun-Kuen Ho. Relatively recursive reals and real functions. Theoretical Computer Science, 210(1):99–120, January 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
210/1/2928.pdf.

HotzelEscardo:1999:IRP

[1169] Martín Hötzel Escardó and Thomas Streicher. Induction and recursion on the partial real line with applications to Real PCF. Theoretical Computer Science, 210(1):121–157, January 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
210/1/2929.pdf.

Safer:1999:PRR

[1170] Taoufik Safer. Polygonal radix representations of complex numbers. Theoretical Computer Science, 210(1):159– 171, January 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/210/1/2930.pdf.

Bronnimann:1999:SDR

[1171] Herve Bronnimann, Ioannis Z. Emiris, Victor Y. Pan, and Sylvain Pion. Sign determination in residue number systems. Theoretical Computer Science, 210(1):173–197, January 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
210/1/2931.pdf.

vanderHoeven:1999:FEH

[1172] Joris van der Hoeven. Fast evaluation of holonomic functions. Theoretical Computer Science, 210(1):199– 215, January 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/210/1/2932.pdf.

Koiran:1999:CFA

[1173] Pascal Koiran and Cristopher Moore. Closed-form analytic maps in one and two dimensions can simulate universal Turing machines. Theoretical Computer Science, 210(1):217–223, January 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        210/1/2939.pdf.

Sakakibara:1999:STL

[1174] Yasubumi Sakakibara and Claudio Ferretti. Splicing on tree-like structures. Theoretical Computer Science, 210(2): 227–243, January 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/210/2/2917.pdf.

Shimozono:1999:AIA

[1175] Shinichi Shimozono. Alphabet indexing for approximating features of symbols. Theoretical Computer Science, 210(2): 245–260, January 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/210/2/2918.pdf.

Akutsu:1999:APT

[1176] Tatsuya Akutsu and Satoru Miyano.

On the approximation of protein threading. Theoretical Computer Science, 210(2):261–275, January 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
210/2/2919.pdf.

Uemura:1999:TAG

[1177] Yasuo Uemura, Aki Hasegawa, Satoshi Kobayashi, and Takashi Yokomori. Tree adjoining grammars for RNA structure prediction. Theoretical Computer Science, 210(2):277–303, January 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
210/2/2920.pdf.

Matsuda:1999:CMS

[1178] H. Matsuda, T. Ishihara, and A. Hashimoto. Classifying molecular sequences using a linkage graph with their pairwise similarities. Theoretical Computer Science, 210(2):305– 325, January 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/210/2/2921.pdf.

Gu:1999:AAG

[1179] Qian-Ping Gu, Shietung Peng, and Hal Sudborough. A 2-approximation algorithm for genome rearrangements by reversals and transpositions. Theoretical Computer Science, 210(2):327– 339, January 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/210/2/2922.pdf.

Ikeda:1999:EAM

[1180] Takahiro Ikeda and Hiroshi Imai. Enhanced A^{∗} algorithms for multiple alignments: optimal alignments for several sequences and k-opt approximate alignments for large cases. Theoretical Computer Science, 210(2):341– 374, January 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/210/2/2923.pdf.

Koutny:1999:ODS

[1181] Maciej Koutny and Eike Best. Operational and denotational semantics for the box algebra. Theoretical Computer Science, 211(1–2):1–83, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        211/1-2/2716.pdf.

Reed:1999:TFS

[1182] G. M. Reed and A. W. Roscoe. The timed failures — Stability model for CSP. Theoretical Computer Science, 211(1–2):85–127, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        211/1-2/2750.pdf.

Stone:1999:RSI

[1183] Matthew Stone. Representing scope in intuitionistic deductions. Theoretical Computer Science, 211(1–2):129– 188, January 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/211/1-2/2780.pdf.

Sun:1999:AGF

[1184] Yong Sun. An algebraic generalization of Frege structures — binding algebras. Theoretical Computer Science, 211(1– 2):189–232, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1999/211/1-2/2706.pdf.

Troelstra:1999:CCS

[1185] A. S. Troelstra. From constructivism to computer science. Theoretical Computer Science, 211(1–2):233–252, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
211/1-2/2708.pdf.

Alur:1999:ECA

[1186] Rajeev Alur, Limor Fix, and Thomas A.

Henzinger. Event-clock automata: a determinizable class of timed automata. Theoretical Computer Science, 211(1–2):253–273, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
211/1-2/2709.pdf.

Comini:1999:CPS

[1187] Marco Comini and Maria Chiara Meo.

Compositionality properties of SLDderivations. Theoretical Computer Science, 211(1–2):275–309, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        211/1-2/2710.pdf.

Engelfriet:1999:MSC

[1188] Joost Engelfriet and Tjalling Gelsema. Multisets and structural congruence of the pi-calculus with replication. Theoretical Computer Science, 211(1–2): 311–337, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/211/1-2/2715.pdf.

Aceto:1999:CEA

[1189] Luca Aceto and Jan Friso Groote. A complete equational axiomatization for MPA with string iteration. Theoretical Computer Science, 211(1–2):339– 374, January 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/211/1-2/2718.pdf.

Bloo:1999:ESE

[1190] Roel Bloo and Herman Geuvers. Explicit substitution On the edge of strong normalization. Theoretical Computer Science, 211(1–2):375–395, January 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/211/1-2/2719.pdf.

Berline:1999:DSM

[1191] Chantal Berline and Klaus Grue. A κ-denotational semantics for map theory in ZFC+SI. Theoretical Computer Science, 211(1–2):397–398, January 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
211/1-2/2936.pdf.

Benke:1999:SCB

[1192] Marcin Benke. Some complexity bounds for subtype inequalities. Theoretical Computer Science, 212(1–2):3– 27, February 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/212/1-2/2954.pdf.

Berarducci:1999:ICT

[1193] Alessandro Berarducci and Mariangiola Dezani-Ciancaglini. Infinite λcalculus and types. Theoretical Computer Science, 212(1–2):29–75, February 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
212/1-2/2955.pdf.

Boley:1999:FLI

[1194] Harold Boley. Functional-logic integration via minimal reciprocal extensions. Theoretical Computer Science, 212(1– 2):77–99, February 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1999/212/1-2/2956.pdf.

Bono:1999:MLC

[1195] Viviana Bono and Michele Bugliesi.

Matching for the lambda calculus of objects. Theoretical Computer Science, 212(1–2):101–140, February 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
212/1-2/2957.pdf.

Dyckhoff:1999:PPI

[1196] Roy Dyckhoff and Luís Pinto. Permutability of proofs in intuitionistic sequent calculi. Theoretical Computer Science, 212(1–2):141–155, February 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
212/1-2/2958.pdf.

Emms:1999:ETC

[1197] Martin Emms and Hans Leiß. Extending the type checker of Standard ML by polymorphic recursion. Theoretical Computer Science, 212(1–2):157– 181, February 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/212/1-2/2959.pdf.

Honsell:1999:SAP

[1198] Furio Honsell and Marina Lenisa. Semantical analysis of perpetual strategies in λ-calculus. Theoretical Computer Science, 212(1–2):183–209, February 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/212/1-2/2960.pdf.

Intrigila:1999:ORG

[1199] B. Intrigila and M. Venturini Zilli.

Orders, reduction graphs and spectra. Theoretical Computer Science, 212(1–2):211–231, February 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
212/1-2/2961.pdf.

Piperno:1999:AVB

[1200] Adolfo Piperno. An algebraic view of the Böhm-out technique. Theoretical Computer Science, 212(1–2):233– 246, February 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/212/1-2/2962.pdf.

Schwichtenberg:1999:TPC

[1201] Helmut Schwichtenberg. Termination of permutative conversions in intuitionistic Gentzen calculi. Theoretical Computer Science, 212(1–2):247– 260, February 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/212/1-2/2963.pdf.

Spreen:1999:FPL

[1202] Dieter Spreen. On functions preserving levels of approximation: A refined model construction for various lambda calculi. Theoretical Computer Science, 212(1–2):261–303, February 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/

                                        212/1-2/2964.pdf. See corrigendum

[1426].

Hazewinkel:1999:P

[1203] Michiel Hazewinkel. Preface. Theoretical Computer Science, 213–214(1): 1–3, February 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1999:SIV

[1204] Anonymous. Subject index volumes 1–200. Theoretical Computer Science, 213–214(1):5–436, February 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1999:RLI

[1205] Anonymous. Reference list of indexed articles. Theoretical Computer Science, 213–214(1):437–528, February 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1999:CIV

[1206] Anonymous. Cumulative index volumes 1–200. Theoretical Computer Science, 213–214(1):529–659, February 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boufkhad:1999:LPI

[1207] Y. Boufkhad and O. Dubois. Length of prime implicants and number of solutions of random CNF formulae. Theoretical Computer Science, 215(1–2):1– 30, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/215/1-2/3001.pdf.

Didier:1999:CEA

[1208] Gilles Didier. Caractérisation des N-écritures et application à l’étude des suites de complexité ultimement n + c^{ste}. (French) [Characterization of N-writings and application to the study of sequences of n + c^{th} ultimate complexity]. Theoretical Computer Science, 215(1–2):31–49, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
215/1-2/2658.pdf.

Nourani:1999:ENH

[1209] Yaghout Nourani and Bjarne Andresen. Exploration of NP-hard enumeration problems by simulated annealing — the spectrum values of permanents. Theoretical Computer Science, 215(1–2):51–68, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
215/1-2/2666.pdf.

Ziadi:1999:OPA

[1210] Djelloul Ziadi and Jean-Marc Champarnaud. An optimal parallel algorithm to convert a regular expression into its Glushkov automaton. Theoretical Computer Science, 215(1–2):69– 87, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/215/1-2/2667.pdf.

Uehara:1999:FRN

[1211] Ryuhei Uehara, Zhi-Zhong Chen, and Xin He. Fast RNC and NC algorithms for maximal path sets. Theoretical Computer Science, 215(1–2):89– 98, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/215/1-2/2668.pdf.

Clerbout:1999:SL

[1212] Mireille Clerbout, Yves Roos, and Isabelle Ryl. Synchronization languages. Theoretical Computer Science, 215(1– 2):99–121, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/215/1-2/2670.pdf.

Hagauer:1999:RMG

[1213] Johann Hagauer, Wilfried Imrich, and Sandi Klavžar. Recognizing median graphs in subquadratic time. Theoretical Computer Science, 215(1–2):123– 136, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/215/1-2/2672.pdf.

Middendorf:1999:SIT

[1214] Martin Middendorf, Welf Löwe, and Wolf Zimmermann. Scheduling inverse trees under the communication model of the LogP-machine. Theoretical Computer Science, 215(1–2):137– 168, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/215/1-2/2697.pdf.

Mihalache:1999:DPG

[1215] Valeria Mihalache. Decidability problems in grammar systems. Theoretical Computer Science, 215(1–2):169– 189, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/215/1-2/2701.pdf.

Giammarresi:1999:DGA

[1216] Dora Giammarresi and Rosa Montalbano. Deterministic generalized automata. Theoretical Computer Science, 215(1–2):191–208, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
215/1-2/2702.pdf.

Kwon:1999:SPT

[1217] Oh-Heum Kwon and Kyung-Yong Chwa. Scheduling parallel tasks with individual deadlines. Theoretical Computer Science, 215(1–2):209– 223, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/215/1-2/2714.pdf.

Durand:1999:CDD

[1218] Bruno Durand and Anne-Cécile Fabret.

On the complexity of deadlock detection in families of planar nets. Theoretical Computer Science, 215(1–2):225– 237, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/215/1-2/2721.pdf.

Kutrib:1999:PCA

[1219] Martin Kutrib. Pushdown cellular automata. Theoretical Computer Science, 215(1–2):239–261, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        215/1-2/2723.pdf.

Buchholz:1999:EPE

[1220] Peter Buchholz. Exact performance equivalence: An equivalence relation for stochastic automata. Theoretical Computer Science, 215(1–2):263– 287, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/215/1-2/2987.pdf.

Koiran:1999:EPP

[1221] Pascal Koiran. Elimination of parameters in the polynomial hierarchy. Theoretical Computer Science, 215(1–2): 289–304, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/215/1-2/3070.pdf.

Andreae:1999:TPG

[1222] Thomas Andreae, Felix Hartenstein, and Andrea Wolter. A two-person game on graphs where each player tries to encircle his opponent’s men. Theoretical Computer Science, 215(1–2): 305–323, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/215/1-2/3017.pdf.

Grigorchuk:1999:EIL

[1223] R. I. Grigorchuk and A. Machì. An example of an indexed language of intermediate growth. Theoretical Computer Science, 215(1–2):325–327, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
215/1-2/2980.pdf.

Bruyere:1999:PCT

[1224] Véronique Bruyère and Christophe Reutenauer. A proof of Choffrut’s theorem on subsequential functions. Theoretical Computer Science, 215(1–2): 329–335, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1999/215/1-2/2982.pdf.

Andersson:1999:FTC

[1225] Arne Andersson, Peter Bro Miltersen, and Mikkel Thorup. Fusion trees can be implemented with AC^{0} instructions only. Theoretical Computer Science, 215(1–2):337–344, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
215/1-2/2990.pdf.

Fischer:1999:MCL

[1226] Herbert Fischer and Harley Flanders.

A minimal code list. Theoretical Computer Science, 215(1–2):345–348, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/215/1-2/3000.pdf.

Csuhaj-Varju:1999:CCC

[1227] Erzsébet Csuhaj-Varjú and György Vaszil. On the computational completeness of context-free parallel communicating grammar systems. Theoretical Computer Science, 215(1–2):349– 358, February 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/215/1-2/3010.pdf.

Wang:1999:FTS

[1228] Lusheng Wang and Xiaohua Jia. Fixed topology Steiner trees and spanning forests. Theoretical Computer Science, 215(1–2):359–370, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        215/1-2/3026.pdf.

Flajolet:1999:SAA

[1229] Philippe Flajolet. Singularity analysis and asymptotics of Bernoulli sums. Theoretical Computer Science, 215(1–2):371–381, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        215/1-2/3030.pdf.

Autebert:1999:SRA

[1230] Jean-Michel Autebert. Some results about centralized PC grammar systems. Theoretical Computer Science, 215(1–2):383–398, February 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cas/tree/store/tcs/sub/1999/
215/1-2/3074.pdf.

Cerone:1999:TBE

[1231] Antonio Cerone and Andrea Maggiolo-Schettini. Time-based expressivity of time Petri nets for system specification. Theoretical Computer Science, 216(1– 2):1–53, March 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/216/1-2/2843.pdf.

Ferreira:1999:BTF

[1232] William Ferreira and Matthew Hennessy. A behavioural theory of firstorder CML. Theoretical Computer Science, 216(1–2):55–107, March 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
216/1-2/2880.pdf.

Zucca:1999:SDA

[1233] Elena Zucca. From static to dynamic abstract data-types: an institution transformation. Theoretical Computer Science, 216(1–2):109– 157, March 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/216/1-2/2947.pdf.

Giacobazzi:1999:RRP

[1234] Roberto Giacobazzi and Francesco Ranzato. The reduced relative power operation on abstract domains. Theoretical Computer Science, 216(1–2): 159–211, March 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/216/1-2/3011.pdf.

Zhang:1999:FVP

[1235] Cui Zhang, Ronald A. Olsson, and Karl N. Levitt. Formal verification of a programming logic for a distributed programming language. Theoretical Computer Science, 216(1–2): 213–235, March 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/216/1-2/2770.pdf.

Degano:1999:NIS

[1236] Pierpaolo Degano and Corrado Priami. Non-interleaving semantics for mobile processes. Theoretical Computer Science, 216(1–2):237–270, March 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        216/1-2/2772.pdf.

Levi:1999:CCP

[1237] Francesca Levi. A compositional μcalculus proof system for statecharts processes. Theoretical Computer Science, 216(1–2):271–310, March 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        216/1-2/2781.pdf.

Burmeister:1999:ATU

[1238] P. Burmeister, M. Monserrat, F. Rosselló, and G. Valiente. Algebraic transformation of unary partial algebras II:

Single-pushout approach. Theoretical Computer Science, 216(1–2):311– 362, March 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/216/1-2/2817.pdf.

Schmidt-Schauss:1999:DBE

[1239] Manfred Schmidt-Schauß. Decidability of behavioural equivalence in unary PCF. Theoretical Computer Science, 216(1–2):363–373, March 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
216/1-2/2855.pdf.

Leino:1999:JSS

[1240] K. Rustan M. Leino and Rajit Manohar. Joining specification statements. Theoretical Computer Science, 216(1–2):375–394, March 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
216/1-2/2984.pdf.

Ying:1999:SPU

[1241] Mingsheng Ying. A shorter proof to uniqueness of solutions of equations. Theoretical Computer Science, 216(1– 2):395–397, March 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1999/216/1-2/3004.pdf.

Worsch:1999:PTM

[1242] Thomas Worsch. Parallel Turing machines with one-head control units and cellular automata. Theoretical Computer Science, 217(1):3–30, March 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        217/1/2967.pdf.

Cattaneo:1999:DBC

[1243] G. Cattaneo, E. Formenti, L. Margara, and G. Mauri. On the dynamical behavior of chaotic cellular automata. Theoretical Computer Science, 217(1): 31–51, March 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/217/1/2968.pdf.

Mazoyer:1999:SOD

[1244] Jacques Mazoyer and Véronique Terrier. Signals in one-dimensional cellular automata. Theoretical Computer Science, 217(1):53–80, March 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        217/1/2969.pdf.

Sipper:1999:CAE

[1245] Moshe Sipper and Marco Tomassini. Computation in artificially evolved, non-uniform cellular automata. Theoretical Computer Science, 217(1):81– 98, March 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/217/1/2970.pdf.

Bandini:1999:MCA

[1246] Stefania Bandini and Giancarlo Mauri.

Multilayered cellular automata. Theoretical Computer Science, 217(1):99– 113, March 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/217/1/2971.pdf.

Chopard:1999:LBC

[1247] Bastien Chopard and Pascal O. Luthi.

Lattice Boltzmann computations and applications to physics. Theoretical Computer Science, 217(1):115– 130, March 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/217/1/2972.pdf.

Gregorio:1999:ACA

[1248] S. Di Gregorio, R. Serra, and M. Villani. Applying cellular automata to complex environmental problems: The simulation of the bioremediation of contaminated soils. Theoretical Computer Science, 217(1):131–156, March 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
217/1/2973.pdf.

Bignone:1999:CML

[1249] Franco A. Bignone. Coupled map lattices dynamics on a variable space for the study of development: A general discussion on Caenorhabditis elegans. Theoretical Computer Science, 217(1): 157–172, March 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/217/1/2974.pdf.

Sander:1999:GLA

[1250] G. Sander. Graph layout for applications in compiler construction. Theoretical Computer Science, 217(2):175– 214, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/217/2/3078.pdf.

Ganter:1999:AEB

[1251] Bernhard Ganter. Attribute exploration with background knowledge. Theoretical Computer Science, 217(2): 215–233, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/217/2/3079.pdf.

Tamassia:1999:ATP

[1252] Roberto Tamassia. Advances in the theory and practice of graph drawing. Theoretical Computer Science, 217(2): 235–254, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/217/2/3080.pdf.

Schmid:1999:BLC

[1253] J. Schmid. Boolean layer cakes. Proceedings ORDAL ’96. Theoretical Computer Science, 217(2):255– 278, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/217/2/3082.pdf.

Gratzer:1999:CL

[1254] G. Grätzer. Congruence lattices 101.

Theoretical Computer Science, 217(2): 279–289, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/217/2/3083.pdf.

Gratzer:1999:SCA

[1255] G. Grätzer and E. T. Schmidt. Some combinatorial aspects of congruence lattice representations. Theoretical Computer Science, 217(2):291– 300, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/217/2/3084.pdf.

Schroder:1999:AFP

[1256] Bernd S. W. Schröder. Algorithms for the fixed point property. Theoretical Computer Science, 217(2):301– 358, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/217/2/3081.pdf.

Fishburn:1999:PST

[1257] Peter Fishburn. Preference structures and their numerical representations. Theoretical Computer Science, 217(2): 359–383, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/217/2/3085.pdf.

Gumm:1999:GAL

[1258] H. Peter Gumm. Generating algebraic laws from imperative programs. Theoretical Computer Science, 217(2):385– 405, April 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/217/2/3086.pdf.

Duquenne:1999:LSD

[1259] Vincent Duquenne. Latticial structures in data analysis. Theoretical Computer Science, 217(2):407–436, April 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
217/2/3087.pdf.

Cassaigne:1999:LVR

[1260] Julien Cassaigne. Limit values of the recurrence quotient of Sturmian sequences. Theoretical Computer Science, 218(1):3–12, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/1/3056.pdf.

deLuca:1999:CFW

[1261] Aldo de Luca. On the combinatorics of finite words. Theoretical Computer Science, 218(1):13–39, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/1/3057.pdf.

Melancon:1999:LWS

[1262] Guy Melançon. Lyndon words and singular factors of Sturmian words. Theoretical Computer Science, 218(1):

41–59, April 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/218/1/3058.pdf.

Carpi:1999:ASS

[1263] Arturo Carpi. On Abelian squares and substitutions. Theoretical Computer Science, 218(1):61–81, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/1/3059.pdf.

Castelli:1999:FWT

[1264] M. Gabriella Castelli, Filippo Mignosi, and Antonio Restivo. Fine and Wilf’s theorem for three periods and a generalization of Sturmian words. Theoretical Computer Science, 218(1):83– 94, April 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/218/1/3060.pdf.

Fraenkel:1999:ENS

[1265] Aviezri S. Fraenkel and Jamie Simpson. The exact number of squares in Fibonacci words. Theoretical Computer Science, 218(1):95–106, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/

218/1/3061.pdf. See corrigendum [1428].

Bruyere:1999:MBC

[1266] Véronique Bruyère and Dominique Perrin. Maximal bifix codes. Theoretical Computer Science, 218(1):107– 121, April 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/218/1/3062.pdf.

Karhumaki:1999:GFW

[1267] Juhani Karhumäki, Wojciech Plandowski, and Wojciech Rytter. Generalized factorizations of words and their algorithmic properties. Theoretical Computer Science, 218(1):123–133, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        218/1/3063.pdf.

Berstel:1999:PWT

[1268] Jean Berstel and Luc Boasson. Partial words and a theorem of Fine and Wilf. Theoretical Computer Science, 218(1): 135–141, April 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/218/1/3064.pdf.

Allouche:1999:TFP

[1269] J.-P. Allouche. Transcendence of formal power series with rational coefficients. Theoretical Computer Science, 218(1):143–160, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        218/1/3065.pdf.

Kolpakov:1999:RFB

[1270] Roman Kolpakov, Gregory Kucherov, and Yuri Tarannikov. On repetitionfree binary words of minimal density. Theoretical Computer Science, 218(1): 161–175, April 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/218/1/3066.pdf.

Ferenczi:1999:CFF

[1271] Sébastien Ferenczi and Zoltán Kása.

Complexity for finite factors of infinite sequences. Theoretical Computer Science, 218(1):177–195, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/1/3067.pdf.

Crochemore:1999:CSS

[1272] Maxime Crochemore, Leszek Gaşieniec, and Wojciech Rytter. Constantspace string-matching in sublinear average time. Theoretical Computer Science, 218(1):197–203, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/1/3068.pdf.

Iliopoulos:1999:QSC

[1273] Costas S. Iliopoulos and Laurent Mouchard. Quasiperiodicity and string covering. Theoretical Computer Science, 218(1):205–216, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/1/3069.pdf.

Mazoyer:1999:AP

[1274] Jacques Mazoyer, Renzo Pinzani, and Jean-Guy Penaud. Avant Propos. Theoretical Computer Science, 218(2):217– 218, May 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/218/2/3137.pdf.

Barcucci:1999:RGT

[1275] Elena Barcucci, Alberto Del Lungo, and Elisa Pergola. Random generation of trees and other combinatorial objects. Theoretical Computer Science, 218(2):219–232, May 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        218/2/3126.pdf.

Denise:1999:URG

[1276] Alain Denise and Paul Zimmermann. Uniform random generation of decomposable structures using floatingpoint arithmetic. Theoretical Computer Science, 218(2):233–248, May 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        218/2/3127.pdf.

Louchard:1999:APS

[1277] G. Louchard. Asymptotic properties of some underdiagonal walks generation algorithms. Theoretical Computer Science, 218(2):249–262, May 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cas/tree/store/tcs/sub/1999/
218/2/3128.pdf.

Mosbah:1999:NUR

[1278] M. Mosbah and N. Saheb. Nonuniform random spanning trees on weighted graphs. Theoretical Computer Science, 218(2):263–271, May 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/2/3129.pdf.

dAndrea:1999:MSH

[1279] Gilles d’Andréa and Christophe Fiorio. Maximal superpositions of horizontally convex polyominoes. Theoretical Computer Science, 218(2):273– 283, May 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/218/2/3130.pdf.

Goles:1999:TAR

[1280] Eric Goles and Ivan Rapaport. Tiling allowing rotations only. Theoretical Computer Science, 218(2):285–295, May 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
218/2/3131.pdf.

Simplot:1999:CRP

[1281] David Simplot. A characterization of recognizable picture languages by tilings by finite sets. Theoretical Computer Science, 218(2):297–323, May 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cas/tree/store/tcs/sub/1999/
                                        218/2/3132.pdf.

Terrier:1999:TDC

[1282] Véronique Terrier. Two-dimensional cellular automata recognizer. Theoretical Computer Science, 218(2):325– 346, May 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/218/2/3133.pdf.

Delorme:1999:DPC

[1283] Marianne Delorme, Jacques Mazoyer, and Laure Tougne. Discrete parabolas and circles on 2D cellular automata. Theoretical Computer Science, 218(2): 347–417, May 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/218/2/3134.pdf.

Apsitis:1999:IIR

[1284] Kalvis Apsītis, Setsuo Arikawa, Rūsin¸š Freivalds, Eiju Hirowatari, and Carl H. Smith. On the inductive inference of recursive real-valued functions. Theoretical Computer Science, 219(1–2): 3–17, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/219/1-2/3089.pdf.

Blanck:1999:EDR

[1285] Jens Blanck. Effective domain representations of H(X), the space of compact subsets. Theoretical Computer Science, 219(1–2):19–48, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
219/1-2/3090.pdf.

Boldi:1999:EJ

[1286] Paolo Boldi and Sebastiano Vigna.

Equality is a jump. Theoretical Computer Science, 219(1–2):49–64, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
219/1-2/3091.pdf.

Brattka:1999:CSE

[1287] Vasco Brattka and Klaus Weihrauch.

Computability on subsets of Euclidean space I: closed and compact subsets. Theoretical Computer Science, 219(1– 2):65–93, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/219/1-2/3092.pdf.

Bridges:1999:CMF

[1288] Douglas S. Bridges. Constructive mathematics: a foundation for computable analysis. Theoretical Computer Science, 219(1–2):95–109, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
219/1-2/3093.pdf.

Cenzer:1999:ISC

[1289] Douglas Cenzer and Jeffrey B. Remmel. Index sets in computable analysis. Theoretical Computer Science, 219 (1–2):111–150, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/219/1-2/3094.pdf.

Chadzelek:1999:AM

[1290] Thomas Chadzelek and Günter Hotz. Analytic machines. Theoretical Computer Science, 219(1–2):151–167, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        219/1-2/3095.pdf.

Edalat:1999:CBS

[1291] Abbas Edalat and Philipp Sünderhauf. Computable Banach spaces via domain theory. Theoretical Computer Science, 219(1–2):169–184, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        219/1-2/3096.pdf.

Hemmerling:1999:AAC

[1292] Armin Hemmerling. On approximate and algebraic computability over the real numbers. Theoretical Computer Science, 219(1–2):185–223, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        219/1-2/3097.pdf.

Hertling:1999:ERM

[1293] Peter Hertling. An effective Riemann Mapping Theorem. Theoretical Computer Science, 219(1–2):225–265, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.

com/cas/tree/store/tcs/sub/1999/
219/1-2/3098.pdf.

Kushner:1999:MCA

[1294] Boris A. Kushner. Markov’s constructive analysis; a participant’s view. Theoretical Computer Science, 219(1–2): 267–285, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/219/1-2/3099.pdf.

Muller:1999:CRV

[1295] Norbert Th. Müller. Computability on random variables. Theoretical Computer Science, 219(1–2):287– 299, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/219/1-2/3100.pdf.

Novak:1999:CUN

[1296] Erich Novak and Henryk Wozńiakowski.

On the cost of uniform and nonuniform algorithms. Theoretical Computer Science, 219(1–2):301–318, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
219/1-2/3101.pdf.

Pour-El:1999:AIC

[1297] Marian Boykan Pour-El. From axiomatics to intrinsic characterization: some open problems in computable analysis. Theoretical Computer Science, 219(1–2):319–329, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.

                                        com/cas/tree/store/tcs/sub/1999/
                                        219/1-2/3102.pdf.

Schroder:1999:OCD

[1298] Matthias Schröder. Online computations of differentiable functions. Theoretical Computer Science, 219(1–2): 331–345, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/219/1-2/3103.pdf.

Stoltenberg-Hansen:1999:CMC

[1299] Viggo Stoltenberg-Hansen and John V. Tucker. Concrete models of computation for topological algebras. Theoretical Computer Science, 219(1–2): 347–378, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/219/1-2/3104.pdf.

Tucker:1999:CWP

[1300] J. V. Tucker and J. I. Zucker. Computation by ‘While’ programs on topological partial algebras. Theoretical Computer Science, 219(1–2):379–420, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        219/1-2/3105.pdf.

Weihrauch:1999:CPM

[1301] Klaus Weihrauch. Computability on the probability measures on the Borel sets of the unit interval. Theoretical Computer Science, 219(1–2):421– 437, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/219/1-2/3106.pdf.

Weihrauch:1999:EGM

[1302] Klaus Weihrauch and Xizhong Zheng.

Effectiveness of the global modulus of continuity on metric spaces. Theoretical Computer Science, 219(1–2):439– 450, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/219/1-2/3107.pdf.

Wozniakowski:1999:WDI

[1303] Henryk Wozńiakowski. Why does information-based complexity use the real number model? Theoretical Computer Science, 219(1–2):451–465, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
219/1-2/3108.pdf.

Yasugi:1999:EPS

[1304] Mariko Yasugi, Takakazu Mori, and Yoshiki Tsujii. Effective properties of sets and functions in metric spaces with computability structure. Theoretical Computer Science, 219(1–2): 467–486, May 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/219/1-2/3109.pdf.

Zhong:1999:CSS

[1305] Ning Zhong. Computability structure of the Sobolev spaces and its applications. Theoretical Computer Science, 219(1–2):487–510, May 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        219/1-2/3110.pdf.

Aguilera:1999:UHF

[1306] Marcos Kawazoe Aguilera, Wei Chen, and Sam Toueg. Using the heartbeat failure detector for quiescent reliable communication and consensus in partitionable networks. Theoretical Computer Science, 220(1):3–30, June 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        220/1/3045.pdf.

Neiger:1999:UKO

[1307] Gil Neiger and Rida A. Bazzi. Using knowledge to optimally achieve coordination in distributed systems. Theoretical Computer Science, 220(1): 31–65, June 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/1/3046.pdf.

Lynch:1999:TCL

[1308] Nancy Lynch, Nir Shavit, Alex Shvartsman, and Dan Touitou. Timing conditions for linearizability in uniform counting networks. Theoretical Computer Science, 220(1):67–91, June 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        220/1/3047.pdf.

Kutten:1999:STA

[1309] Shay Kutten and Boaz Patt-Shamir.

Stabilizing time-adaptive protocols. Theoretical Computer Science, 220(1): 93–111, June 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/220/1/3048.pdf.

Fekete:1999:ESD

[1310] Alan Fekete, David Gupta, Victor Luchangco, Nancy Lynch, and Alex Shvartsman. Eventually-serializable data services. Theoretical Computer Science, 220(1):113–156, June 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
220/1/3049.pdf.

Lui:1999:SSN

[1311] King-Shan Lui and Shmuel Zaks.

Scheduling in synchronous networks and the greedy algorithm. Theoretical Computer Science, 220(1):157– 183, June 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/220/1/3050.pdf.

Beimel:1999:RCP

[1312] Amos Beimel and Matthew Franklin.

Reliable communication over partially authenticated networks. Theoretical Computer Science, 220(1):185–210, June 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
220/1/3051.pdf.

Chaudhuri:1999:WFI

[1313] Soma Chaudhuri, Maurice Herlihy, and Mark R. Tuttle. Wait-free implementations in message-passing systems. Theoretical Computer Science, 220(1):211– 245, June 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/1/3052.pdf.

Gehrke:1999:RCL

[1314] Johannes E. Gehrke, C. Greg Plaxton, and Rajmohan Rajaraman. Rapid convergence of a local load balancing algorithm for asynchronous rings. Theoretical Computer Science, 220(1):247– 265, June 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/1/3053.pdf.

Mavronicolas:1999:LRW

[1315] Marios Mavronicolas and Dan Roth. Linearizable read/write objects. Theoretical Computer Science, 220(1):267– 319, June 06, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/1/3054.pdf.

Ambainis:1999:OMC

[1316] Andris Ambainis, Sanjay Jain, and Arun Sharma. Ordinal mind change complexity of language identification. Theoretical Computer Science, 220(2): 323–343, June 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/2/3166.pdf.

Fadel:1999:HHS

[1317] R. Fadel, K. V. Jakobsen, J. Katajainen, and J. Teuhola. Heaps and heapsort on secondary storage. Theoretical Computer Science, 220(2):345– 362, June 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/220/2/3167.pdf.

Houle:1999:ODE

[1318] Micheal E. Houle, Ewan Tempero, and Gavin Turner. Optimal dimensionexchange token distribution on complete binary trees. Theoretical Computer Science, 220(2):363–376, June 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
220/2/3168.pdf.

Liu:1999:VRS

[1319] Chuchang Liu and Mehmet A. Orgun.

Verification of reactive systems using temporal logic with clocks. Theoretical Computer Science, 220(2):377– 408, June 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/220/2/3169.pdf.

Mason:1999:ALT

[1320] Ian A. Mason and Carolyn L. Talcott.

Actor languages: Their syntax, semantics, translation, and equivalence. Theoretical Computer Science, 220(2): 409–467, June 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/2/3170.pdf.

Roberts:1999:LMR

[1321] Alan Roberts and Antonios Symvonis. On-line matching routing on trees. Theoretical Computer Science, 220(2): 469–488, June 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/2/3171.pdf.

Zhang:1999:SCA

[1322] Yan Zhang. Specifying causality in action theories: a default logic approach. Theoretical Computer Science, 220(2): 489–513, June 17, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/220/2/3172.pdf.

Andreev:1999:WCH

[1323] Alexander E. Andreev, Andrea E. F. Clementi, and José D. P. Rolim. Worstcase hardness suffices for derandomization: a new method for hardnessrandomness trade-offs. Theoretical Computer Science, 221(1–2):3–18, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3185.

Bartal:1999:LRA

[1324] Yair Bartal and Stefano Leonardi. Online routing in all-optical networks. Theoretical Computer Science, 221(1– 2):19–39, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1999&volume=221&issue=
1-2&aid=3186.

Bassino:1999:ESL

[1325] Frédérique Bassino, Marie-Pierre Béal, and Dominique Perrin. Enumerative sequences of leaves and nodes in rational trees. Theoretical Computer Science, 221(1–2):41–60, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=221&issue=1-2&aid=3187.

Durand:1999:TQ

[1326] Bruno Durand. Tilings and quasiperiodicity. Theoretical Computer Science, 221(1–2):61–75, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL

http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/
browse.cgi?year=1999&volume=221&
issue=1-2&aid=3188.

Erdos:1999:FLS

[1327] Péter L. Erdős, Michael A. Steel, László A. Székely, and Tandy J. Warnow. A few logs suffice to build (almost) all trees: Part II. Theoretical Computer Science, 221(1–2):77–118, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=221&issue=1-2&aid=3189.

Erlebach:1999:OWR

[1328] Thomas Erlebach, Klaus Jansen, Christos Kaklamanis, Milena Mihail, and Pino Persiano. Optimal wavelength routing on directed fiber trees. Theoretical Computer Science, 221(1–2):119–137, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3190.

Krumke:1999:IST

[1329] Sven O. Krumke, Hartmut Noltemeier, Hans-C. Wirth, Madhav V. Marathe, R. Ravi, S. S. Ravi, and R. Sundaram. Improving spanning trees by upgrading nodes. Theoretical Computer Science, 221(1–2):139–155, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3191.

Manzini:1999:CEC

[1330] Giovanni Manzini and Luciano Margara. A complete and efficiently computable topological classification of D-dimensional linear cellular automata over Z_{m}. Theoretical Computer Science, 221(1–2):157–177, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3192.

Apt:1999:ECP

[1331] Krzysztof R. Apt. The essence of constraint propagation. Theoretical Computer Science, 221(1–2):179–210, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=221&issue=1-2&aid=3193.

Bouajjani:1999:SRA

[1332] Ahmed Bouajjani and Peter Habermehl. Symbolic reachability analysis of FIFO-channel systems with nonregular sets of configurations. Theoretical Computer Science, 221(1–2):211–250, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=221&issue=1-2&aid=3194.

Burkart:1999:MCF

[1333] Olaf Burkart and Bernhard Steffen.

Model checking the full modal mucalculus for infinite sequential processes. Theoretical Computer Science, 221(1–2):251–270, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=221&issue=1-2&aid=3195.

deVink:1999:BPT

[1334] E. P. de Vink and J. J. M. M.

Rutten. Bisimulation for probabilistic transition systems: a coalgebraic approach. Theoretical Computer Science, 221(1–2):271–293, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3196.

DiGianantonio:1999:ADT

[1335] Pietro Di Gianantonio. An abstract data type for real numbers. Theoretical Computer Science, 221(1–2):295–326, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3197.

Fu:1999:VMP

[1336] Yuxi Fu. Variations on mobile processes. Theoretical Computer Science, 221(1–2):327–368, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3198.

Henzinger:1999:DTC

[1337] Thomas A. Henzinger and Peter W. Kopke. Discrete-time control for rectangular hybrid automata. Theoretical Computer Science, 221(1–2):369–392, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=221&issue=1-2&aid=3199.

Honda:1999:GTA

[1338] Kohei Honda and Nobuko Yoshida. Game-theoretic analysis of call-by-value computation. Theoretical Computer Science, 221(1–2):393–456, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=221&issue=1-2&aid=3200.

Sangiorgi:1999:NDU

[1339] Davide Sangiorgi. The name discipline of uniform receptiveness. Theoretical Computer Science, 221(1–2):457–493, June 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=221&issue=1-2&aid=3201.

Abramsky:1999:SSD

[1340] S. Abramsky, S. J. Gay, and R. Nagarajan. A specification structure for deadlock-freedom of synchronous processes. Theoretical Computer Science, 222(1–2):1–53, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=222&issue=1-2&aid=3006.

Cegielski:1999:AFO

[1341] Patrick Cegielski and Denis Richard.

On arithmetical first-order theories allowing encoding and decoding of lists. Theoretical Computer Science, 222(1–2):55–75, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/

                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=222&issue=1-2&aid=2816.

File:1999:POA

[1342] Gilberto Filé and Francesco Ranzato. The powerset operator on abstract interpretations. Theoretical Computer Science, 222(1–2):77–111, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=222&issue=1-2&aid=2842.

Cao:1999:OTS

[1343] Feng Cao and Al Borchers. Optimal transmission schedules for lightwave networks embedded with de Bruijn graphs. Theoretical Computer Science, 222(1–2):113–131, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=222&issue=1-2&aid=2935.

Gurevich:1999:MSR

[1344] Yuri Gurevich and Andrei Voronkov. Monadic simultaneous rigid E-unification. Theoretical Computer Science, 222(1– 2):133–152, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1999&volume=222&issue=
                                        1-2&aid=2945.

Marchiori:1999:DAD

[1345] Elena Marchiori. Design of abstract domains using first-order logic. Theoretical Computer Science, 222(1–2):

153–179, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1999&volume=222&issue=
1-2&aid=3041.

Colson:1999:DFP

[1346] Loïc Colson. On diagonal fixed points of increasing functions. Theoretical Computer Science, 222(1–2):181–186, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=222&issue=1-2&aid=3135.

Dufourd:1999:PBN

[1347] Catherine Dufourd and Alain Finkel. A polynomial λ-bisimilar normalization for reset Petri nets. Theoretical Computer Science, 222(1–2):187–194, July 06, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=222&issue=1-2&aid=3155.

Kullmann:1999:NMS

[1348] O. Kullmann. New methods for 3-SAT decision and worst-case analysis. Theoretical Computer Science, 223(1–2): 1–72, July 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/223/1-2/2848.pdf.

Droubay:1999:PSW

[1349] Xavier Droubay and Giuseppe Pirillo.

Palindromes and Sturmian words. Theoretical Computer Science, 223(1–2): 73–85, July 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/223/1-2/2724.pdf.

Rambow:1999:IPF

[1350] Owen Rambow and Giorgio Satta. Independent parallelism in finite copying parallel rewriting systems. Theoretical Computer Science, 223(1–2): 87–120, July 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/223/1-2/2726.pdf.

Fouks:1999:TAT

[1351] Jean-Denis Fouks. Towards an algorithmic theory of adaptation. Theoretical Computer Science, 223(1–2): 121–142, July 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/223/1-2/2747.pdf.

Kim:1999:HGH

[1352] Changwook Kim and Tae Eui Jeong. HRNCE grammars — a hypergraph generating system with an eNCE way of rewriting. Theoretical Computer Science, 223(1–2):143–178, July 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        223/1-2/2794.pdf.

Shih:1999:NPT

[1353] Wei-Kuan Shih and Wen-Lian Hsu. A new planarity test. Theoretical Computer Science, 223(1–2):179–191, July 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
223/1-2/2942.pdf.

Gao:1999:NMW

[1354] Sui-Xiang Gao, Xiao-Dong Hu, and Weili Wu. Nontrivial monotone weakly symmetric Boolean functions with six variables are elusive. Theoretical Computer Science, 223(1–2):193–197, July 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/tcs/sub/1999/
223/1-2/3154.pdf.

Baaz:1999:NGC

[1355] Matthias Baaz. Note on the generalization of calculations. Theoretical Computer Science, 224(1–2):3–11, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3112.

Beklemishev:1999:PFI

[1356] Lev D. Beklemishev. Parameter free induction and provably total computable functions. Theoretical Computer Science, 224(1–2):13–33, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3113.

Courcelle:1999:MSO

[1357] Bruno Courcelle. The monadic second-order logic of graphs XI: Hierarchical decompositions of connected graphs. Theoretical Computer Science, 224(1–2):35–58, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=224&issue=1-2&aid=3114.

Ershov:1999:S

[1358] Yu. L. Ershov. On d-spaces. Theoretical Computer Science, 224(1–2): 59–72, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1999&volume=224&issue=
                                        1-2&aid=3115.

Gradel:1999:LTV

[1359] Erich Grädel and Martin Otto. On logics with two variables. Theoretical Computer Science, 224(1–2): 73–113, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1999&volume=224&issue=
                                        1-2&aid=3116.

deGroote:1999:ACC

[1360] Philippe de Groote. An algebraic correctness criterion for intuitionistic multiplicative proof-nets. Theoretical Computer Science, 224(1–2):115–134, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3117.

Heinemann:1999:TAM

[1361] Bernhard Heinemann. Temporal aspects of the modal logic of subset spaces. Theoretical Computer Science, 224(1–2):135–155, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3118.

Marion:1999:MSA

[1362] Jean-Yves Marion. From multiple sequent for additive linear logic to decision procedures for free lattices. Theoretical Computer Science, 224(1–2):157–172, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3119.

Lisitsa:1999:LOG

[1363] Alexei Lisitsa and Vladimir Sazonov.

Linear ordering on graphs, antifounded sets and polynomial time computability. Theoretical Computer Science, 224(1–2):173–213, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3120.

Diekert:1999:SWE

[1364] Volker Diekert, Yuri Matiyasevich, and Anca Muscholl. Solving word equations modulo partial commutations. Theoretical Computer Science, 224(1–2):215–235, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=224&issue=1-2&aid=3121.

Otto:1999:BIP

[1365] Martin Otto. Bisimulation-invariant PTIME and higher-dimensional μcalculus. Theoretical Computer Science, 224(1–2):237–265, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=224&issue=1-2&aid=3122.

Perrier:1999:PCF

[1366] G. Perrier. A PSPACE-complete fragment of second-order linear logic. Theoretical Computer Science, 224(1–2): 267–289, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.elsevier.com/cgi-bin/cas/
                                        tree/store/tcs/cas_sub/browse/browse.
                                        cgi?year=1999&volume=224&issue=
                                        1-2&aid=3123.

Tseytin:1999:FRD

[1367] Gregory S. Tseytin. A formalization of reasoning not derived from standard predicate logic. Theoretical Computer Science, 224(1–2):291–317, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3124.

Voronkov:1999:SRU

[1368] Andrei Voronkov. Simultaneous rigid E-unification and other decision problems related to the Herbrand theorem. Theoretical Computer Science, 224(1–2):319–352, August 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=224&issue=1-2&aid=3125.

Pelletier:1999:RFD

[1369] Maryse Pelletier and Jacques Sakarovitch.

On the representation of finite deterministic 2-tape automata. Theoretical Computer Science, 225(1–2):1– 63, August 18, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/225/1-2/2996.pdf.

Crescenzi:1999:MNC

[1370] Pierluigi Crescenzi and Luca Trevisan.

Max NP-completeness made easy. Theoretical Computer Science, 225(1–2): 65–79, August 18, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/225/1-2/3014.pdf.

Roka:1999:SBC

[1371] Zsuzsanna Róka. Simulations between cellular automata on Cayley graphs.

Theoretical Computer Science, 225(1– 2):81–111, August 18, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        tcs/sub/1999/225/1-2/2749.pdf.

Clementi:1999:INA

[1372] Andrea E. F. Clementi and Luca Trevisan. Improved non-approximability results for minimum vertex cover with density constraints. Theoretical Computer Science, 225(1–2):113–128, August 18, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        225/1-2/2762.pdf.

Chuan:1999:SMW

[1373] Wai-Fong Chuan. Sturmian morphisms and α-words. Theoretical Computer Science, 225(1–2):129–148, August 18, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/tcs/sub/1999/
                                        225/1-2/2775.pdf.

Hakala:1999:SWE

[1374] Ismo Hakala and Juha Kortelainen. On the system of word equations x_{0}u_{1}^{i}x_{1}u_{2}^{i}x_{2}u_{3}^{i}x_{3} = y_{0}v_{1}^{i}y_{1}v_{2}^{i}y_{2}v_{3}^{i}y_{3}(i = 0, 1, 2, …) in a free monoid. Theoretical Computer Science, 225(1–2):149– 161, August 18, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

                                        elsevier.com/cas/tree/store/tcs/
                                        sub/1999/225/1-2/2791.pdf.

Wong:1999:OPC

[1375] Pak-Ken Wong. Optimal path cover problem on block graphs. Theoretical Computer Science, 225(1–2):163– 169, August 18, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/225/1-2/2997.pdf.

Basart:1999:SCC

[1376] J. M. Basart and P. Guitart. A solution for the coloured cubes problem. Theoretical Computer Science, 225(1– 2):171–176, August 18, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1999/225/1-2/3139.pdf.

Gobel:1999:SRP

[1377] Manfred Göbel. The “smallest” ring of polynomial invariants of a permutation group which has no finite SAGBI bases w.r.t. any admissible order. Theoretical Computer Science, 225(1–2):177– 184, August 18, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

elsevier.com/cas/tree/store/tcs/
sub/1999/225/1-2/3144.pdf.

Lutz:1999:FRK

[1378] Jack H. Lutz and David L. Schweizer.

Feasible reductions to Kolmogorov– Loveland stochastic sequences. Theoretical Computer Science, 225(1–2): 185–194, August 18, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.com/cas/tree/store/
tcs/sub/1999/225/1-2/3202.pdf.

Adleman:1999:SAD

[1379] Leonard M. Adleman, Jonathan De-Marrais, and Ming-Deh Huang. A subexponential algorithm for discrete logarithms over hyperelliptic curves of large genus over GF(q). Theoretical Computer Science, 226(1–2):7–18, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3222.

Chua:1999:ART

[1380] Seng Kiat Chua, Ka Hin Leung, and San Ling. Attack on RSA-type cryptosystems based on singular cubic curves over Z/nZ. Theoretical Computer Science, 226(1–2):19–27, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3223.

Cusick:1999:ARC

[1381] Thomas W. Cusick. The Ajtai random class of lattices. Theoretical Computer Science, 226(1–2):29–36, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3224.

Feng:1999:TCC

[1382] Dengguo Feng. Three characterizations of correlation-immune functions over rings Z_{N}. Theoretical Computer Science, 226(1–2):37–43, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3225.

Gollmann:1999:DBB

[1383] Dieter Gollmann. Dual bases and bit-serial multiplication in F_{q^{n}}. Theoretical Computer Science, 226(1–2): 45–59, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1999&volume=226&issue=
1-2&aid=3226.

Klapper:1999:AFS

[1384] Andrew Klapper and Jinzhong Xu. Algebraic feedback shift registers. Theoretical Computer Science, 226(1–2): 61–92, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.
cgi?year=1999&volume=226&issue=
1-2&aid=3227.

Niederreiter:1999:ASC

[1385] Harald Niederreiter and Michael Vielhaber. An algorithm for shifted continued fraction expansions in parallel linear time. Theoretical Computer Science, 226(1–2):93–104, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3228.

Niemi:1999:EVN

[1386] Valtteri Niemi and Ari Renvall. Efficient voting with no selling of votes. Theoretical Computer Science, 226(1–2):105–116, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=226&issue=1-2&aid=3229.

Ruanaidh:1999:CCP

[1387] Joseph Ó Ruanaidh, Holger Petersen, Alexander Herrigel, Shelby Pereira, and Thierry Pun. Cryptographic copyright protection for digital images based on watermarking techniques. Theoretical Computer Science, 226(1–2):117–142, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=226&issue=1-2&aid=3230.

Tao:1999:FAP

[1388] Renji Tao and Shihua Chen. On finite automaton public-key cryptosystem. Theoretical Computer Science, 226(1–2):143–172, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/tcs/
                                        cas_sub/browse/browse.cgi?year=
                                        1999&volume=226&issue=1-2&aid=3231.

Varadharajan:1999:DER

[1389] Vijay Varadharajan, Khanh Quoc Nguyen, and Yi Mu. On the design of efficient RSA-based off-line electronic cash schemes. Theoretical Computer Science, 226(1–2):173–184, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3232.

Zhang:1999:STC

[1390] Chunru Zhang, Kwok-Yan Lam, and Sushil Jajodia. Scalable threshold closure. Theoretical Computer Science, 226(1–2):185–206, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3233.

Zheng:1999:RTN

[1391] Yuliang Zheng, Xian-Mo Zhang, and Hideki Imai. Restriction, terms and nonlinearity of Boolean functions. Theoretical Computer Science, 226(1–2):207–223, September 17, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.
com/cgi-bin/cas/tree/store/tcs/
cas_sub/browse/browse.cgi?year=
1999&volume=226&issue=1-2&aid=3234.

Abramsky:1999:FAI

[1392] Samson Abramsky and Guy McCusker.

Full abstraction for Idealized Algol with passive expressions. Theoretical Computer Science, 227(1–2):3– 42, September 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
151/17/17/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/17/article.pdf.

Bierman:1999:CLC

[1393] G. M. Bierman. A classical linear λ-calculus. Theoretical Computer Science, 227(1–2):43–78, September 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        151/17/18/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/18/article.pdf.

Danos:1999:RIO

[1394] Vincent Danos and Laurent Regnier. Reversible, irreversible and optimal λ-machines. Theoretical Computer Science, 227(1–2):79–97, September 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        151/17/19/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/19/article.pdf.

Guerrini:1999:GTS

[1395] Stefano Guerrini. A general theory of sharing graphs. Theoretical Computer Science, 227(1–2):99– 151, September 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        151/17/20/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/20/article.pdf.

Hu:1999:CCC

[1396] Hongde Hu and Andre Joyal. Coherence completions of categories. Theoretical Computer Science, 227(1–2): 153–184, September 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
151/17/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/151/17/21/article.pdf.

Kobayashi:1999:DCL

[1397] Naoki Kobayashi, Toshihiro Shimizu, and Akinori Yonezawa. Distributed concurrent linear logic programming. Theoretical Computer Science, 227(1–2):185–220, September 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
151/17/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/151/17/22/article.pdf.

Metayer:1999:PEA

[1398] François Métayer. Polynomial equivalence among systems LLNC, LLNC a and LLNC 0. Theoretical Computer Science, 227(1–2):221–229, September 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
151/17/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/151/17/23/article.pdf.

Turner:1999:OIL

[1399] David N. Turner and Philip Wadler.

Operational interpretations of linear logic. Theoretical Computer Science, 227(1–2):231–248, September 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        151/17/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/24/article.pdf.

Girard:1999:DC

[1400] Jean-Yves Girard. On denotational completeness. Theoretical Computer Science, 227(1–2):249–273, September 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        151/17/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/25/article.pdf.

Girard:1999:CBS

[1401] Jean-Yves Girard. Coherent Banach spaces: a continuous denotational semantics. Theoretical Computer Science, 227(1–2):275–297, September 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        151/17/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/26/article.pdf.

Lincoln:1999:OCL

[1402] Patrick D. Lincoln, John C. Mitchell, and Andre Scedrov. Optimization complexity of linear logic proof games. Theoretical Computer Science, 227(1–2): 299–331, September 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        151/17/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/151/17/27/article.pdf.

Okada:1999:PSC

[1403] Mitsuhiro Okada. Phase semantic cutelimination and normalization proofs of first- and higher-order linear logic. Theoretical Computer Science, 227(1–2): 333–396, September 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
151/17/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/151/17/28/article.pdf.

Anonymous:1999:Ea

[1404] Anonymous. Editorial(s). Theoretical Computer Science, 228(1–2): 1–3, October 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/19/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/19/article.pdf.

Gordon:1999:BTF

[1405] Andrew D. Gordon. Bisimilarity as a theory of functional programming. Theoretical Computer Science, 228(1– 2):5–47, October 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/20/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/20/article.pdf.

Freyd:1999:B

[1406] P. J. Freyd et al. Bireflectivity. Theoretical Computer Science, 228(1–2): 49–76, October 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/21/article.pdf.

Gardner:1999:CAC

[1407] Philippa Gardner. Closed action calculi. Theoretical Computer Science, 228(1–2):77–103, October 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/22/article.pdf.

Jeffrey:1999:FAS

[1408] Alan Jeffrey. A fully abstract semantics for a higher-order functional language with nondeterministic computation. Theoretical Computer Science, 228(1– 2):105–150, October 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/23/article.pdf.

Jones:1999:LPC

[1409] Neil D. Jones. LOGSPACE and PTIME characterized by programming languages. Theoretical Computer Science, 228(1–2):151–174, October 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/24/article.pdf.

Maraist:1999:CNC

[1410] J. Maraist, M. Odersky, D. N. Turner, and P. Wadler. Call-by-name, callby-value, call-by-need and the linear lambda calculus. Theoretical Computer Science, 228(1–2):175–210, October 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/25/article.pdf.

OHearn:1999:SCI

[1411] P. W. O’Hearn, A. J. Power, M. Takeyama, and R. D. Tennent. Syntactic control of interference revisited. Theoretical Computer Science, 228(1– 2):211–252, October 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/26/article.pdf.

OHearn:1999:OIY

[1412] Peter W. O’Hearn and Uday S.

Reddy. Objects, interference, and the Yoneda embedding. Theoretical Computer Science, 228(1–2):253– 282, October 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/27/article.pdf.

Anonymous:1999:Ia

[1413] Anonymous.  Index.  Theoreti-

cal Computer Science, 228(1–2):283, October 28, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
158/19/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/158/19/28/article.pdf.

Anonymous:1999:Eb

[1414] Anonymous. Editorial(s). Theoretical Computer Science, 229(1–2): 1–2, November 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
157/19/19/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/157/19/19/article.pdf.

Eiben:1999:TEA

[1415] A. E. Eiben and G. Rudolph. Theory of evolutionary algorithms: a bird’s eye view. Theoretical Computer Science, 229(1–2):3–9, November 6, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
157/19/20/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/157/19/20/article.pdf.

Lozano:1999:GAB

[1416] J. A. Lozano, P. Larrañaga, M. Graña, and F. X. Albizuri. Genetic algorithms: bridging the convergence gap. Theoretical Computer Science, 229(1–2): 11–22, November 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
157/19/21/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/41/
16/157/19/21/article.pdf.

He:1999:CRG

[1417] Jun He and Lishan Kang. On the convergence rates of genetic algorithms. Theoretical Computer Science, 229(1– 2):23–39, November 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
157/19/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/157/19/22/article.pdf.

vanNimwegen:1999:SDR

[1418] Erik van Nimwegen, James P. Crutchfield, and Melanie Mitchell. Statistical Dynamics of the Royal Road Genetic Algorithm. Theoretical Computer Science, 229(1–2):41–102, November 6, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
157/19/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/157/19/23/article.pdf.

Vose:1999:RHS

[1419] Michael D. Vose. Random heuristic search. Theoretical Computer Science, 229(1–2):103–142, November 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/41/16/157/19/24/abstract.
html;   http://www.elsevier.nl/
gej-ng/10/41/16/157/19/24/article.
pdf.

Kosters:1999:FAG

[1420] Walter A. Kosters, Joost N. Kok, and Patrik Floréen. Fourier analysis of genetic algorithms. Theoretical Computer Science, 229(1–2):143– 175, November 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        157/19/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/157/19/25/article.pdf.

Cedeno:1999:ASN

[1421] Walter Cedeño and V. Rao Vemuri. Analysis of speciation and niching in the multi-niche crowding GA. Theoretical Computer Science, 229(1–2): 177–197, November 6, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        157/19/26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/157/19/26/article.pdf.

Anonymous:1999:Ib

[1422] Anonymous.  Index.  Theoreti-

cal Computer Science, 229(1–2):199, November 6, 1999. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        157/19/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/157/19/27/article.pdf.

Boudol:2000:SCN

[1423] Gérard Boudol. On the semantics of the call-by-name CPS transform. Theoretical Computer Science, 234(1–2): 309–321, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/35/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/35/article.pdf.

Hemaspaandra:2000:ERC

[1424] L. Hemaspaandra, A. Hoene, and M. Ogihara. Erratum to “Reducibility classes of P -selective sets” [Theoret. Comput. Sci 155(2) (1996) 447– 457]. Theoretical Computer Science, 234(1–2):323, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/
170/21/36/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/41/
16/170/21/36/article.pdf.  See

[294].

Bernardo:2001:CTE

[1425] Marco Bernardo and Roberto Gorrieri. Corrigendum to “A tutorial on EMPA: a theory of concurrent processes with nondeterminism, priorities, probabilities and time” — [Theoret. Comput. Sci. 202 (1998) 1–54]. Theoretical Computer Science, 254(1–2):691–694, March 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl../
43/abstract.html;   http://www.
elsevier.nl/gej-ng/10/41/16/191/
20/43/article.pdf;   http://www.
elsevier.nl/PII/S0304397597001278.

See [1044].

Spreen:2001:CFP

[1426] Dieter Spreen. Corrigendum to “On functions preserving levels of approximation: a refined model construction for various lambda calculi” [Theoret. Comput. Sci. 212 (1999) 261–303]. Theoretical Computer Science, 266(1–2): 997–998, September 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/41/16/
                                        216/27/60/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/41/
                                        16/216/27/60/article.pdf;   http:
                                        //www.elsevier.nl/PII/S0304397598001443.

See [1202].

Kurka:2010:EET

[1427] Petr Kůrka. Erratum to: Entropy of Turing machines with moving head. Theoretical Computer Science, 411(31– 33):2999–3000, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [587].

Fraenkel:2014:CEN

[1428] Aviezri S. Fraenkel and Jamie Simpson. Corrigendum to “The exact number of squares in Fibonacci words” [Theoret. Comput. Sci. 218 (1) (1999) 95– 106]. Theoretical Computer Science, 547(??):122, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004356. See

[1265].